### v1

# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

Abstract.
A flexible model for non-stationary Gaussian random fields on hypersurfaces is introduced. The class of random fields on curves and surfaces is characterized by a power spectral density of a second order elliptic differential operator. Sampling is done by a Galerkin–Chebyshev approximation based on the surface finite element method and Chebyshev polynomials. Strong error bounds are shown with convergence rates depending on the smoothness of the approximated random field.
Key words and phrases:
Gaussian random fields, non-stationary random fields, stochastic partial differential equations, surface finite element method, Chebyshev approximation, Gaussian processes.
1991 Mathematics Subject Classification:
60G60, 60H35, 60H15,65C30, 60G15, 58J05, 41A10, 65N30,65M60
Acknowledgment: The authors thank Julie Rowlett and Iosif Polterovich for helpful discussions. This work was supported in part by the European Union (ERC, StochMan, 101088589), by the Mines Paris - PSL / INRAE “Geolearning” chair, by the Swedish Research Council (VR) through grant no. 2020-04170, by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Chalmers AI Research Centre (CHAIR)
1. Introduction
Random fields are powerful tools for modeling spatially dependent data. They have found uses in a wide range of applications, for instance in geostatistics, cosmological data analysis, climate modeling, and biomedical imaging (Marinucci and Peccati, 2011; Farag, 2014). One challenge in the modeling of spatial data is non-stationary behavior, i.e., different behaviors in different parts of the domain. Another challenge is that the domain may be a non-Euclidean space, for instance, a surface such as the sphere or on the cortical surface of the brain. In this paper, we present a surface finite element-based method to sample a flexible class of non-stationary random fields on curves and surfaces and show its strong convergence. The method, building on the foundational work for stationary fields introduced in Lang and Pereira (2023), is an extension of the stochastic partial differential equation (SPDE) approach pioneered by Whittle (1963) and popularized by Lindgren et al. (2011). The idea behind our method is to color white noise by applying a function of an elliptic differential operator $\mathcal{L}$. Formally, we study Gaussian random fields on curves and surfaces of the form
(1) $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W},$
where $\gamma$ is a function, called the power spectral density, $\mathcal{L}$ is an elliptic differential operator and $\mathcal{W}$ denotes white noise. By letting the coefficients of the differential operator vary over the domain, we can obtain local, non-stationary behaviors. If $1/\gamma$ is well-defined over $\amsmathbb{R}_{+}$, one may formally view $\mathcal{Z}$ as the solution to the stochastic partial differential equation
$(1/\gamma)(\mathcal{L})\mathcal{Z}=\mathcal{W}.$
(a) A random field where the choice of coefficients are inspired the continents.
(b) A random field on the cortical surfaces only locally activated.
(c) A random field on a star-shaped domain.
Figure 1. Examples of random field samples generated with our method.
For instance, consider the three examples depicted in Figure 1. To generate random field samples, there are two components of $\mathcal{L}$ that we can vary: the diffusion matrix and the potential. In Figure 1(a), we use a diffusion matrix to give the field preferred directions, more specifically it elongates field in the northwest-southeast direction. The potential is large over the continents and small over the oceans, effectively “turning off” the random field over land. In Figure 1(b), the potential is small in the front of the brain and large elsewhere, so that the field is only large in the front of the brain. Finally, Figure 1(c) shows the method used to generate a non-stationary random field in the one-dimensional case. To illustrate the value of the field at a point, we move it in the normal direction for a distance proportional to the value of the field. In all three cases, we see that the field behaves locally varying over the domain. With the suggested model, we can achieve preferred directions, local activation, and local deactivation.
The computational method we use to solve Equation 1, i.e., sample the random fields, is based on the surface finite element method (SFEM), a computational method pioneered by Dziuk (1988); Dziuk and Elliott (2013) and that has been used in the context of the generation of random fields in for instance Bonito et al. (2024); Jansson et al. (2022); Lang and Pereira (2023). Our main mathematical contribution is a strong convergence result. Using a functional calculus approach to the finite element discretization error, we obtain a strong rate of convergence of order $\mathcal{O}(K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}})$, where $d=2$ for surfaces and $d=1$ for curves, and $K_{\alpha}(h)$ is a dimension-dependent logarithmic factor.
The SPDE approach to random fields and their approximation have been studied previously for both surfaces and Euclidean geometries, examples include Bonito et al. (2024); Borovitskiy et al. (2020); Bolin et al. (2020); Cox and Kirchner (2020); Jansson et al. (2022); Lang and Pereira (2023); Lindgren et al. (2011, 2022). However, to the best of our knowledge, we are the first to present a strong error analysis for the non-stationary case on hypersurfaces. Moreover, we achieve this without requiring an explicit approximation of the eigenfunctions of the elliptic operator, and contrary to for instance Borovitskiy et al. (2020), computing the eigenfunctions. In fact, the computation of eigenfunctions, with theoretical guarantees, is a notoriously difficult problem, see e.g., Boffi (2010). Our method circumvents this issue by introducing a Chebyshev approximation.
Our main contribution is a powerful, efficient, and flexible tool for the modeling and sampling of non-stationary random fields on curves and surfaces with proven accuracy. In particular, using tools from complex analysis and operator theory, we derive strong error bounds for approximations of arbitrary sufficiently smooth transformations of elliptic operators, where we do not require assumptions on the approximability of individual eigenfunctions.
The paper is structured as follows: In Section 2, we introduce the relevant deterministic framework. We provide the necessary background on geometry and functional analysis in Section 2.1. This is followed by a description of the main computational tool, surface finite elements in Section 2.2. Finally, Section 2.3 provides the relevant error estimates in the deterministic setting. In Section 3, we collect all material in the stochastic setting. We introduce first the class of considered random fields in Section 3.1 and their Galerkin–Chebyshev approximation in Section 3.2. The proof of its strong convergence is split into the SFEM error in Section 3.3 and the Chebyshev approximation error Section 3.4. For numerical confirmation of our theoretical results, we refer to our earlier work Lang and Pereira (2023) and the results presented in Bonito et al. (2024). The source code used to generate the figures is available at this address: https://github.com/mike-pereira/SFEMsim.
2. Deterministic theory: geometry, functional analysis and finite elements
Before we are able to approximate random fields on hypersurfaces, we need to introduce and partially extend the existing literature on surface finite element approximations due to so far unconsidered error bounds required in our stochastic setting. We introduce the functional analytic setting in Section 2.1, discuss surface finite element methods in Section 2.2 and show error bounds in the deterministic setting in Section 2.3.
2.1. Geometric and functional analytic setting
Let $\mathcal{M}\subset\amsmathbb{R}^{d+1}$ be a $d$-dimensional ($d\leq 2$) compact oriented smooth hypersurface ($k\geq 2$) without boundary, i.e., for any $x_{0}\in\mathcal{M}$, there exists an open set $U_{x_{0}}\subset\amsmathbb{R}^{d+1}$ containing $x_{0}$ and a function $\phi_{x_{0}}\in C^{\infty}(U_{x_{0}})$ such that $\nabla\phi_{x_{0}}\neq 0$ on $\mathcal{M}\cap U_{x_{0}}$ and
$\mathcal{M}\cap U_{x_{0}}=\{x\in U_{x_{0}},\phi_{x_{0}}(x)=0\}.$
The tangent space of $\mathcal{M}$ at $x\in\mathcal{M}$ is the $d$-dimensional subspace of $\amsmathbb{R}^{d+1}$ given by $T_{x}\mathcal{M}=[\nabla\phi_{x}]^{\perp}$ (where $\nabla$ denotes the usual gradient of functions of $C^{1}(\amsmathbb{R}^{d+1})$ and $\perp$ denotes the orthogonal complement in $\amsmathbb{R}^{d+1}$ with respect to the standard Euclidean inner product.). Since $\mathcal{M}$ is oriented, there exists a smooth map $\nu:\mathcal{M}\rightarrow\amsmathbb{R}^{d+1}$ assigning to each point $x\in\mathcal{M}$ a unit vector $\nu(x)=\pm\nabla\phi_{x}/\|\nabla\phi_{x}\|$ perpendicular to the tangent space $T_{x}\mathcal{M}$. Our hypersurface $\mathcal{M}$ is a Riemannian manifold equipped with the metric $g$ that is the pullback of the Euclidean metric on $\amsmathbb{R}^{d+1}$. For instance, if $\mathcal{M}=\amsmathbb{S}^{2}$, this results in the standard round metric.
Let $\nabla_{\mathcal{M}}$ be the gradient operator acting on differentiable functions of $\mathcal{M}$, and let $\Delta_{\mathcal{M}}$ denote the Laplace–Beltrami operator on $(\mathcal{M},g)$. We denote by $\,\mathrm{d}A$ the surface measure on $\mathcal{M}$, and by $L^{2}(\mathcal{M})$ the Hilbert space of $\,\mathrm{d}A$-measurable square integrable complex-valued functions, equipped with the inner product $(\cdot,\cdot)_{L^{2}(\mathcal{M})}$ defined by
$(u,v)_{L^{2}(\mathcal{M})}=\int_{\mathcal{M}}u\overline{v}\,\mathrm{d}A,\quad u,v\in L^{2}(\mathcal{M}).$
The Sobolev spaces with smoothness index $s\in\amsmathbb{R}^{+}$ are then defined via Bessel potentials by
$H^{s}(\mathcal{M})=\left(I-\Delta_{\mathcal{M}}\right)^{-s/2}L^{2}(\mathcal{M}),$
with corresponding norm $\|\cdot\|_{H^{s}(\mathcal{M})}=\|\left(I-\Delta_{\mathcal{M}}\right)^{s/2}\cdot\|_{L^{2}(\mathcal{M})}$. For $s<0$, $H^{s}(\mathcal{M})$ is defined as the space of distributions generated by
(2) $H^{s}(\mathcal{M})=\left\{u=\left(I-\Delta_{\mathcal{M}}\right)^{k}v,\leavevmode\nobreak\ v\in H^{2k+s}(\mathcal{M})\right\},$
where $k\in\amsmathbb{N}$ is the smallest integer such that $02k+s>$ 0. In this case, the corresponding norm is given by $\|u\|_{H^{s}(\mathcal{M})}=\|v\|_{H^{2k+s}(\mathcal{M})}$. We set $H^{0}(\mathcal{M})=L^{2}(\mathcal{M})$. The reader is referred to Herrmann et al. (2018), Strichartz (1983), Triebel (1985), and references therein for more details on Sobolev spaces defined using Bessel potentials.
In this work, we consider elliptic differential operators associated to bilinear forms $\mathsf{A}_{\mathcal{M}}$ given by
(3) $\mathsf{A}_{\mathcal{M}}(u,v)=\int_{\mathcal{M}}(\mathcal{D}\nabla_{\mathcal{M}}u)\cdot({\nabla_{\mathcal{M}}\overline{v}})\,\mathrm{d}A+\int_{\mathcal{M}}(Vu)\overline{v}\,\mathrm{d}A,\quad u,v\in H^{1}(\mathcal{M}),$
where for any $x_{0}\in\mathcal{M}$ the diffusion matrix $\mathcal{D}(x_{0})=[D_{ij}(x_{0})]_{i,j=1}^{d+1}$ is a real-valued, symmetric matrix such that for any $w\in T_{x_{0}}\mathcal{M}$, $\mathcal{D}(x_{0})w\in T_{x_{0}}\mathcal{M}$ and $(\mathcal{D}(x_{0})w)\cdot\overline{w}>0$ if $w\neq 0$. In particular, since $T_{x_{0}}\mathcal{M}=[\nu(x_{0})^{\perp}]$, $\mathcal{D}(x_{0})$ is simply a matrix admitting $\nu(x_{0})$ as an eigenvector with some eigenvalue $\mu_{1}(x_{0})\in\amsmathbb{R}$, and such that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 associated with its other eigenvectors are positive. Without loss of generality, we may assume that $\mu_{1}(x_{0})=0$, meaning that $\mathcal{D}\nu=0$, that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 are uniformly lower-bounded and upper-bounded on $\mathcal{M}$ by positive constants, and that $D_{ij}\in L^{\infty}(\mathcal{M})$ for any $11\leq i,j\leq d+$ 1. Finally, we assume that $V\in L^{\infty}(\mathcal{M})$ is a real-valued function that satisfies $V_{-}\leq V\leq V_{+}$ for some $0<V_{-}\leq V_{+}<+\infty$.
Throughout this paper, let $\mathsf{A}_{\mathcal{M}}$ be coercive and continuous, i.e., there exist positive constants $\delta$ and $M$ such that for all $u,v\in H^{1}(\mathcal{M})$,
(4) $\mathsf{A}_{\mathcal{M}}(u,u)\geq\delta\|u\|_{H^{1}(\mathcal{M})}^{2},$
(5) $\left|\mathsf{A}_{\mathcal{M}}(u,v)\right|\leq M\|u\|_{H^{1}(\mathcal{M})}\,\|v\|_{H^{1}(\mathcal{M})}.$
Following Yagi (2010, Equation (1.33)), $\mathsf{A}_{\mathcal{M}}$ gives rise to an associated elliptic differential operator $\mathcal{L}:H^{1}(\mathcal{M})\rightarrow H^{-1}(\mathcal{M})$ defined weakly by
$\mathsf{A}_{\mathcal{M}}(u,v)=\int_{\mathcal{M}}(\mathcal{L}u)\overline{v}\,\mathrm{d}A,\quad u,v\in H^{1}(\mathcal{M}).$
The spectral properties of this operator are detailed in the next proposition, which is proven in Appendix A.
Proposition 2.1.
Let $\delta>0$ be the coercivity constant defined in Equation 4. There exists a set of eigenpairs $\{(\lambda_{i},e_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ consisting of a sequence of increasing real-valued eigenvalues $0<\delta\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{e_{i}\}_{i\in\amsmathbb{N}}$ forms an orthonormal basis of $L^{2}(\mathcal{M})$ where each $e_{i}$ is real-valued.
Since the operator $\mathcal{L}$ differs from the Laplace–Beltrami operator only by a zeroth-order potential term and a diffusion function in the second order term, switching between the two operators corresponds to a change of metric on $\mathcal{M}$. Therefore, the eigenvalue problem for $\mathcal{L}$ is equivalent to that for the Laplace–Beltrami operator on $\mathcal{M}$ equipped with a possibly rough metric if the coefficients of $\mathcal{D}$ are not smooth. The results in Bandara et al. (2021) imply growth rates on the eigenvalues in accordance with Weyl’s law, and more specifically that there exist $c_{\mathcal{M}},C_{\mathcal{M}}>0$ such that for any $i\in\amsmathbb{N}$
(6) $c_{\mathcal{M}}i^{2/d}\leq\lambda_{i}\leq C_{\mathcal{M}}i^{2/d}.$
As a last step in this subsection, we introduce nonlinear functions of $\mathcal{L}$ which allow later in Section 3 for the definition of a variety of Gaussian random fields. For that, we call a function $\gamma:\amsmathbb{R}_{+}\rightarrow\amsmathbb{R}$ an $\alpha$-power spectral density if
  1. 1)
    $\gamma$ is extendable to a holomorphic function on $H_{\pi/2}\coloneqq\{z\in\amsmathbb{C}:|\arg z|\leq\pi/2\}$.
  2. 2)
    There exist constants $C_{\gamma}>0$ and $\alpha>0$ such that for all $z\in H_{\pi/2}$,
    (7) $\left|\gamma(z)\right|\leq C_{\gamma}|z|^{-\alpha}.$
Applying a power spectral density to $\mathcal{L}$ results in a linear operator $\gamma(\mathcal{L})$ whose action on functions $f\in L^{2}(\mathcal{M})$ is defined by
(8) $\gamma(\mathcal{L})f=\sum_{i=1}^{\infty}\gamma(\lambda_{i})(f,e_{i})_{L^{2}(\mathcal{M})}e_{i},$
where $\{(\lambda_{i},e_{i})\}_{i\in\amsmathbb{N}}$ are the eigenpairs of $\mathcal{L}$ defined in Proposition 2.1. A typical example is the function $\gamma(\lambda)=(\kappa^{2}+\lambda)^{-\alpha}$ for $\alpha>d/4$ and $\kappa>0$, which can be used to obtain Whittle–Matérn random fields (Lang and Pereira, 2023).
The goal of the remainder of this section is to study the approximation of functions of the form $u=\gamma(\mathcal{L})f$, where $f\in L^{2}(\mathcal{M})$. Formally, if $1/\gamma$ is well-defined on the spectrum of $\mathcal{L}$, then this is the solution to the partial differential equation $\gamma(\mathcal{L})^{-1}u=f$.
2.2. SFEM–Galerkin approximation
The idea behind the surface finite element method, as introduced by Dziuk (1988), is to work on a polyhedral approximation of the surface that is in some sense close to the true surface $\mathcal{M}$. More precisely, fix $h>0$ and let $\mathcal{M}_{h}$ be a piecewise polygonal surface consisting of non-degenerate simplices (for $d=2$, triangles and for $d=1$, line segments) with vertices on $\mathcal{M}$, and such that $h$ is the size of the largest simplex defined as the in-ball radius. The set of simplices making up the discretized surface is denoted by $\mathcal{T}_{h}$, thus meaning that
$\mathcal{M}_{h}=\bigcup_{T_{j}\in\mathcal{T}_{h}}T_{j},$
and we assume that for any two simplices in $\mathcal{T}_{h}$, it holds that their intersection is either empty, or a common edge or vertex.
Following (Dziuk and Elliott, 2013, Section 1.4.1), we assume that the triangulation $\mathcal{T}_{h}$ is quasi-uniform, shape-regular, and that the number of simplices sharing the same vertex can be upper-bounded by a constant independent of $h$. In turn, these two assumption imply that $N_{h}\propto h^{-d}$ where $N_{h}\in\amsmathbb{N}$ denotes the number of vertices of $\mathcal{M}_{h}$.
$\mathcal{M}$$\mathcal{M}_{h}$$\nu_{h}$$\nu$xa(x)
Figure 2. One-dimensional illustration of the lift. The lift is along the normal vector $\nu$ to the surface $\mathcal{M}$.
The discrete surface $\mathcal{M}_{h}$ is close to the true surface $\mathcal{M}$ in the sense that $\mathcal{M}_{h}$ is contained in a small neighborhood around $\mathcal{M}$ defined as follows. First, note that $\mathcal{M}$ can be seen as the boundary of some bounded open set $G\subset\amsmathbb{R}^{d+1}$ with exterior normal $\nu$. Then, following Dziuk and Elliott (2013, Section 2.3), we consider that there exists some (small) $\varpi>0$ such that $\mathcal{M}_{h}$ is contained in a so-called tubular neighborhood $U_{\varpi}$ of $\mathcal{M}$ defined by
$U_{\varpi}=\{x\in\amsmathbb{R}^{d+1}:|d_{s}(x)|<\varpi\},$
where $d_{s}:\amsmathbb{R}^{d+1}\rightarrow\amsmathbb{R}$ denotes the oriented distance function given by
$d_{s}(x)=\begin{cases}\inf_{y\in\mathcal{M}}|x-y|,\leavevmode\nobreak\ \leavevmode\nobreak\ x\in\amsmathbb{R}^{d+1}\setminus{G},\\ \inf_{y\in\mathcal{M}}-|x-y|,\leavevmode\nobreak\ \leavevmode\nobreak\ x\in G.\end{cases}$
We denote by $\,\mathrm{d}A_{h}$ the surface measure on $\mathcal{M}_{h}$, and by $L^{2}(\mathcal{M}_{h})$ the Hilbert space of $\,\mathrm{d}A_{h}$-measurable square integrable functions, equipped with the inner product $(\cdot,\cdot)_{L^{2}(\mathcal{M}_{h})}$ defined by
$(u_{h},v_{h})_{L^{2}(\mathcal{M}_{h})}=\int_{\mathcal{M}_{h}}u_{h}\overline{v}_{h}\,\mathrm{d}A_{h},\quad u_{h},v_{h}\in L^{2}(\mathcal{M}_{h}).$
Following (Bonito et al., 2020, Section 1.2.1), we denote by $\sigma:\mathcal{M}_{h}\rightarrow\amsmathbb{R}^{+}$ the area element given by $\sigma=\mathrm{d}A/\mathrm{d}A_{h}$, such that for all $v\in L^{2}(\mathcal{M})$,
(9) $\int_{\mathcal{M}}v\,\mathrm{d}A=\int_{\mathcal{M}_{h}}\sigma v^{-\ell}\,\mathrm{d}A_{h},$
where we next introduce the lift and its inverse denoted by $\ell$ and $-\ell$, respectively.
A key element of SFEM is that we can move between $\mathcal{M}$ and $\mathcal{M}_{h}$ using the so-called lift operator. To construct the lift operator, we note that $d_{s}\in C^{k}(U_{\varpi})$ for $k\geq 2$, and that for any $x\in U_{\varpi}$, there exists a unique $a(x)\in\mathcal{M}$ such that
$x=a(x)+d_{s}(x)\nu(a(x)),$
where $\nu$ denotes the normal at $a(x)$ to $\mathcal{M}$. In particular, this implies that any point $x\in U_{\varpi}$ can be uniquely described by the pair $(a(x),d_{s}(x))\in\mathcal{M}\times\amsmathbb{R}$, and this procedure defines an isomorphism $p\colon\mathcal{M}_{h}\rightarrow\mathcal{M}$ given by
$p(x)=x-d_{s}(x)\nu(a(x)),\quad x\in\mathcal{M}_{h}.$
Therefore, any function $\eta\colon\mathcal{M}_{h}\rightarrow\amsmathbb{C}$ may be lifted to $\mathcal{M}$ by $\eta^{\ell}=\eta\circ p^{-1}:\mathcal{M}\rightarrow\amsmathbb{C}$. Likewise, the inverse lift of any function $\zeta:\mathcal{M}\rightarrow\amsmathbb{C}$ is given by $\zeta^{-\ell}=\zeta\circ p:\mathcal{M}_{h}\rightarrow\amsmathbb{C}$. The procedure is illustrated in Figure 2 in the one-dimensional setting. Note that the points on the discretized surface are lifted along the normal $\nu$ to the surface $\mathcal{M}$.
The mapping $a$ is used to define the gradient of functions on $\mathcal{M}_{h}$ (Dziuk and Elliott, 2013). More specifically, for a differentiable $\eta:\mathcal{M}_{h}\to\amsmathbb{C}$, the gradient is given by
(10) $\nabla_{\mathcal{M}_{h}}\eta(x)=\nabla\check{\eta}(x)-(\nabla\check{\eta}(x)\cdot\nu_{h}(x))\nu_{h}(x)\in T_{x}\mathcal{M}_{h},$
where $\nu_{h}$ is the normal of $\mathcal{M}_{h}$ and $\check{\eta}$ is the continuous extension of $\eta$ defined by $\check{\eta}:x\in U_{\varpi}\mapsto\check{\eta}(x)=\eta^{\ell}(a(x))$. With this definition, the Laplace–Beltrami operator on $\mathcal{M}_{h}$ can be defined by $\Delta_{\mathcal{M}_{h}}=\nabla_{\mathcal{M}_{h}}\cdot\nabla_{\mathcal{M}_{h}}$, and Sobolev spaces on $\mathcal{M}_{h}$ are defined in complete analogy to those on $\mathcal{M}$.
The analogue of the bilinear form $\mathsf{A}_{\mathcal{M}}$ on $\mathcal{M}_{h}$ is given by
(11) $\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}+\int_{\mathcal{M}_{h}}(V^{-\ell}u_{h})\overline{v}_{h}\,\mathrm{d}A_{h},\quad u_{h},v_{h}\in H^{1}(\mathcal{M}_{h}),$
where $\mathcal{D}^{-\ell}=[D_{ij}^{-\ell}]_{i,j=1}^{d+1}$. As in Dziuk and Elliott (2013), we assume that there exists $h_{0}\leavevmode\nobreak\ \in\leavevmode\nobreak\ (0,1)$ small enough, such that $\mathsf{A}_{\mathcal{M}_{h}}$ is coercive (and continuous) whenever $h\leq h_{0}$. Unless stated otherwise, we now assume that this last condition on $h$ is fulfilled.
To conclude this subsection, we introduce the (linear) finite element space $S_{h}$ on $\mathcal{M}_{h}$. The finite element space $S_{h}$ is defined as the complex span of the standard real-valued nodal basis $\psi_{1},\dots,\psi_{N_{h}}\colon\mathcal{M}\rightarrow\amsmathbb{R}$, where for any $i\in\{1,\dots,N_{h}\}$, $\psi_{i}|_{T}$ is a polynomial of at most degree one taking the value $1$ at the $i$-th vertex of $\mathcal{M}_{h}$ and 00 at all the other vertices, i.e.,
$S_{h}=\operatorname{span}\left(\psi_{1},\dots,\psi_{N_{h}}\right)\subset H^{1}(\mathcal{M}_{h}).$
By construction, $S_{h}$ is a vector space of dimension $N_{h}$. Its counterpart on $\mathcal{M}$ is the lifted finite element space $S_{h}^{\ell}$ given by
$S_{h}^{\ell}=\left\{\phi_{h}^{\ell},\phi_{h}\in S_{h}\right\}\subset H^{1}(\mathcal{M}).$
On $S_{h}$, we can, as $\mathcal{L}$ on $\mathcal{M}$, associate to the bilinear form $\mathsf{A}_{\mathcal{M}_{h}}$ in Equation 11 a linear operator $\mathsf{L}_{h}:S_{h}\rightarrow S_{h}$ which maps any $u_{h}\in S_{h}$ to the unique $\mathsf{L}_{h}u_{h}\in S_{h}$ satisfying, for any $v_{h}\in S_{h}$, the equality
$\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=(\mathsf{L}_{h}u_{h},v_{h})_{L^{2}(\mathcal{M}_{h})}.$
Similarly, if the bilinear form $\mathsf{A}_{\mathcal{M}}$ introduced in Equation 3 is restricted to $S_{h}^{\ell}$, we can associate it to a linear operator $\mathcal{L}_{h}:S_{h}^{\ell}\rightarrow S_{h}^{\ell}$ that maps any $u_{h}^{\ell}\in S_{h}$ to the unique $\mathcal{L}_{h}u_{h}^{\ell}\in S_{h}^{\ell}$ that satisfies, for any $v_{h}^{\ell}\in S_{h}^{\ell}$, the equality
$\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})=(\mathcal{L}_{h}u_{h}^{\ell},v_{h}^{\ell})_{L^{2}(\mathcal{M})}.$
Since the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$ are coercive, positive definite, Hermitian and have real coefficients, these two operators are diagonalizable in the sense that they each give rise to a set of $N_{h}$ eigenpairs (Hall, 2013). On the one hand, there exists a sequence $0\leq\Lambda_{1}^{h}\leq\dots\leq\Lambda_{N_{h}}^{h}$ and an $L^{2}(\mathcal{M}_{h})$-orthonormal basis $E_{1}^{h},\ldots,E_{N_{h}}^{h}$ of $S_{h}$ such that
$\mathsf{L}_{h}E_{i}^{h}=\Lambda_{i}^{h}E_{i}^{h},\quad 1\leq i\leq N_{h},$
and similarly there exists a sequence $0\leq\lambda_{1}^{h}\leq\dots\leq\lambda_{N_{h}}^{h}$ and an $L^{2}(\mathcal{M})$-orthonormal basis $e_{1}^{h},\ldots,e_{N_{h}}^{h}$ of $S_{h}^{\ell}$ such that
$\mathcal{L}_{h}e_{i}^{h}=\lambda_{i}^{h}e_{i}^{h},\quad 1\leq i\leq N_{h}.$
In particular, using the same approach as in Proposition 2.1, we can assume that the eigenfunctions $\{E_{i}^{h}\}_{1\leq i\leq N_{h}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are all real-valued. The eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$ are linked to one another through the following lemma, due to Bonito et al. (2018, Lemma 3.1), Bonito et al. (2024, Lemma 4.1) and Strang and Fix (2008, Theorem 6.1).
In the following, $A\lesssim B$ is shorthand for that there is a constant $C>0$ such that $A\leq CB$.
Lemma 2.2 (Eigenvalue error bounds).
Let $\{\lambda_{i}\}_{i\in\amsmathbb{N}}$, $\{\lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ and $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ denote the eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$, respectively. Then,
(12) $\lambda_{i}\leq\lambda_{i}^{h}\lesssim(1+h^{2})\lambda_{i},\quad 1\leq i\leq N_{h},$
and
(13) $|\lambda_{i}^{h}-\Lambda_{i}^{h}|\lesssim h^{2}\lambda_{i}^{h}\lesssim h^{2}\lambda_{i},\quad 1\leq i\leq N_{h}.$
Finally, we remark that the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ of $\mathsf{L}_{h}$ can be linked to the eigenvalues of some classical finite element matrices. Let $\bm{C}$ and $\bm{R}$ be the so-called mass matrix and stiffness matrix, respectively, and defined from the nodal basis by
(14) $\bm{C}$ $=\left[(\psi_{k},\psi_{l})_{L^{2}(\mathcal{M}_{h})}\right]_{1\leq k,l\leq N_{h}},\quad\bm{R}$ $=\left[\mathsf{A}_{\mathcal{M}_{h}}(\psi_{k},\psi_{l})\right]_{1\leq k,l\leq N_{h}}.$
As defined, $\bm{C}$ is a symmetric positive definite matrix and $\bm{R}$ is a symmetric positive semi-definite matrix. Let then $\sqrt{\bm{C}}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ be an invertible matrix satisfying $\sqrt{\bm{C}}(\sqrt{\bm{C}})^{T}=\bm{C}$. Then, by Lang and Pereira (2023, Corollary 3.2), the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ are also the eigenvalues of the matrix $\bm{S}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ defined by
(15) $\bm{S}=\big{(}\sqrt{\bm{C}}\big{)}^{-1}\bm{R}\big{(}\sqrt{\bm{C}}\big{)}^{-T}.$
Besides, if $\bm{\psi}$ denotes the vector-valued function given by $\bm{\psi}=(\psi_{1},\dots,\psi_{N_{h}})^{T}$, then the mapping $F:\amsmathbb{R}^{N_{h}}\rightarrow S_{h}$, defined by
$F(\bm{v})=F(\bm{v})=\bm{\psi}^{T}\big{(}\sqrt{\bm{C}}\big{)}^{-T}\bm{v},\quad\bm{v}\in\amsmathbb{R}^{N_{h}},$
is an isomorphism whose inverse maps the eigenfunctions $\{E_{i}^{h}\}_{1\leq i\leq N_{h}}$ to (orthonormal) eigenvectors of $\bm{S}$. This means in particular that $\bm{S}$ can also be written as
(16) $\bm{S}=\bm{V}\operatorname{Diag}\big{(}\Lambda_{1}^{h},\dots,\Lambda_{1}^{N_{h}}\big{)}\bm{V}^{T},$
where $\bm{V}=\big{(}F^{-1}(E_{1}^{h})|\cdots|F^{-1}(E_{N_{h}}^{h})\big{)}\in\amsmathbb{R}^{N_{h}\times N_{h}}$.
2.3. Deterministic error analysis
Based on the introduced framework, we are now in place to quantify the error between functions of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$. Let $P_{h}:L^{2}(\mathcal{M})\rightarrow S_{h}^{\ell}$ be the $L^{2}$-projection onto $S_{h}^{\ell}$ and let $\mathsf{P}_{h}:L^{2}(\mathcal{M}_{h})\rightarrow S_{h}$ the $L^{2}$-projection onto $S_{h}$. We note that the operators $\mathcal{L}$, $\mathcal{L}_{h}$, and $\mathsf{L}_{h}$ define norms that are equivalent to the standard Sobolev norms, i.e.,
(19) $\begin{gathered}\|\mathcal{L}^{1/2}v\|_{L^{2}(\mathcal{M})}\sim\|v\|_{H^{1}(\mathcal{M})},\leavevmode\nobreak\ \|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M})}\sim\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},\\ \|\mathcal{L}_{h}^{1/2}v_{h}\|_{L^{2}(\mathcal{M})}\sim\|v_{h}\|_{H^{1}(\mathcal{M})},\end{gathered}$
for all $v\in H^{1}(\mathcal{M})$, and all $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$.
With that at hand we are ready to state our main result in this section.
Proposition 2.3.
Let $\gamma$ be an $\alpha$-power spectral density. Then, there is a constant $C>0$ such that for all $f\in L^{2}(\mathcal{M})$
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $\|_{L^{2}(\mathcal{M})}\leq\begin{cases}Ch^{2}\|f\|_{L^{2}(\mathcal{M})},&\alpha>1,\\ Ch^{2}|\log h|\|f\|_{L^{2}(\mathcal{M})},&\alpha=1,\\ Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})},&\alpha<1.\end{cases}$
To prove this proposition we rely on a representation of functions of operators based on Cauchy–Stieltjes integrals, which are constructed as follows. Since the sesquilinear form defined by $\mathsf{A}_{\mathcal{M}}$ is continuous and coercive, the associated operators $\mathcal{L}$ and $\mathcal{L}_{h}$ are sectorial with some (common) angle $\theta\in(0,\pi/2)$ (Yagi, 2010, Theorem 2.1). Therefore, the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ are contained in the complement of the set $G_{\theta}=\{z\in\amsmathbb{C},\theta\leq\arg(z)\leq\pi\}$, as illustrated in Figure 3(a), and the following inequalities are satisfied for any $z\in G_{\theta}$ (cf. Yagi (2010, Equation (2.2))):
$\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v\|_{L^{2}(\mathcal{M})},\quad v\in L^{2}(\mathcal{M}),$
(20) $\|(z-\mathcal{L}_{h})^{-1}v_{h}^{\ell}\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v_{h}^{\ell}\|_{L^{2}(\mathcal{M})},\quad v_{h}^{\ell}\in S_{h}^{\ell},$
where $C_{\theta}>0$ is a generic constant. Note in particular that by definition, $G_{\theta}$ is contained in the resolvent sets of $\mathcal{L}$ and $\mathcal{L}_{h}$, and that any power spectral density $\gamma$ is bounded, holomorphic and satisfies the inequality $|z|^{\alpha}|\gamma(z)|\lesssim 1$ for any $z\in G_{\theta}$. Hence, the operators $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$ can be defined as functional calculi of the operators $\mathcal{L}$ and $\mathcal{L}_{h}$ as (Yagi, 2010, Chapter 16, Section 1.2) by
(21) $\gamma(\mathcal{L})=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)(z-\mathcal{L})^{-1}\,\mathrm{d}z\quad\text{and}\quad\gamma(\mathcal{L}_{h})=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)(z-\mathcal{L}_{h})^{-1}\,\mathrm{d}z,$
where $\Gamma\subset\amsmathbb{C}$ is any integral contour surrounding the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ and contained in $G_{\theta}$. In particular, these new definitions of functions of operators are independent of the choice of $\Gamma$, and coincide with the spectral definitions previously introduced in Equation 8 (cf. e.g. (Yagi, 2010, Remark 2.7)).
In the remainder, we split the contour $\Gamma$ into
(22) $\Gamma=\Gamma_{+}\cup\Gamma_{0}\cup\Gamma_{-},$
where $\Gamma_{+}$ is parametrized by $g_{+}(t)=te^{i\theta}$ for $t\in(\infty,\delta_{0}]$, $\Gamma_{0}$ by $g_{0}(t)=\delta_{0}e^{it}$ for $t\in[\theta,-\theta]$ and $\Gamma_{-}$ by $g_{-}(t)=te^{i\theta}$ for $t\in[\delta_{0},\infty)$ with $\delta_{0}<\delta/2$ (see Figure 3(b) for an illustration). We use this contour to prove Proposition 2.3, while relying on the following bounds for the resolvent error along $\Gamma$. The proof is included in Appendix A and is an adaption of results from Fujita and Suzuki (1991) and Bonito et al. (2024).
Lemma 2.4.
For any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$, and $\beta\in[0,1]$,
$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2\beta}|z|^{-(1-\beta)}.$
We now provide a proof for Proposition 2.3.
[…]
(a) Illustration of $G_{\theta}$ being the complement of the shaded blue slice.
[…]
(b) Illustration of the contour $\Gamma=\Gamma_{+}\cup\Gamma_{-}\cup\Gamma_{0}$ used in the proof of Proposition 2.3.
Figure 3. Contours used to define the Cauchy–Stieltjes integral representation of operators.
Proof of Proposition 2.3.
Let $f\in L^{2}(\mathcal{M})$. For any $z\in\Gamma$, set $e_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f$, which yields, using the integral representations of $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$,
$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)e_{h}(z)\,\mathrm{d}z.$
The definition of $\Gamma$ and its parametrization allow to decompose the integral as
$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{-1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{+}(t))e_{h}(g_{+}(t))g^{\prime}_{+}(t)\,\mathrm{d}t-\frac{1}{2\pi i}\int_{-\theta}^{\theta}\gamma(g_{0}(t))e_{h}(g_{0}(t))g_{0}^{\prime}(t)\,\mathrm{d}t$
$\quad+\frac{1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{-}(t))e_{h}(g_{-}(t))g^{\prime}_{-}(t)\,\mathrm{d}t,$
where we recall that $g_{+}^{\prime}(t)=e^{i\theta},\leavevmode\nobreak\ g_{-}^{\prime}(t)=e^{-i\theta}$ and $g_{0}^{\prime}(t)=i\delta_{0}e^{it}$. Taking norms on both sides of this equality and using the triangle inequality then gives
$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
$\quad\leq\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{+}(t))|\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{-\theta}^{\theta}|\gamma(g_{0}(t))|\|e_{h}(g_{0}(t)\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{-}(t))|\left\|e_{h}(g_{-}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
$\quad=I_{+}+I_{0}+I_{-}.$
Let us start by bounding $I_{0}$. We apply Equation 7 and Lemma 2.4 with $\beta=1$ to obtain
$I_{0}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}|g_{0}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}\delta_{0}^{-\alpha}\,\mathrm{d}t\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}.$
To bound $I_{+}$, we distinguish between the three cases $\alpha>1$, $\alpha<1$ and $\alpha=1$. When $\alpha>1$, we apply Equation 7 and Lemma 2.4 with $\beta=1$ to get
$I_{+}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\int_{\delta_{0}}^{\infty}|t|^{-\alpha}\,\mathrm{d}t=C\frac{\delta_{0}^{-(\alpha-1)}}{\alpha-1}h^{2}\|f\|_{L^{2}(\mathcal{M})}.$
For $\alpha<1$, recall that $h<h_{0}<1$ and assume without loss of generality that $h_{0}^{-2}>\delta_{0}$. We split $I_{+}=I_{+1}+I_{+2}$ with
$I_{+1}=\int_{\delta_{0}}^{h^{-2}}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,I_{+2}=\int_{h^{-2}}^{\infty}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t.$
Using Equation 7 and Lemma 2.4 with $\beta=1$ we bound $I_{+1}$ by
$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t$
$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{0}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\frac{h^{-2(1-\alpha)}}{1-\alpha}\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
We proceed in the same manner to bound $I_{+2}$, but use Lemma 2.4 with $\beta=0$,
$I_{+2}\leq C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|g_{+}(t)|^{-\alpha}|g_{+}(t)|^{-1}\,\mathrm{d}t=C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|t|^{-\alpha-1}\,\mathrm{d}t\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
For $\alpha=1$ and the same splitting $I_{+}=I_{+1}+I_{+2}$ we obtain for $I_{+1}$ with Lemma 2.4 for $\beta=1$
$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}t^{-1}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(\log(h^{-2})-\log\delta_{0})$
$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log(h^{-2})|+|\log\delta_{0}|)\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log h|+|\log\delta_{0}|).$
Assuming again $h<h_{0}$ and $h_{0}^{-2}>\delta_{0}$, we finally bound
$I_{+1}$ $\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h}\right|\right)\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h_{0}}\right|\right)$
$\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}.$
The bound for $I_{+2}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}$ is obtained in the same way as in the case $\alpha>1$. Hence, $I_{+}\leq Ch^{2}|\log h|\leavevmode\nobreak\ \|f\|_{L^{2}(\mathcal{M})}$ for $\alpha=1$.
Finally, note that by symmetry, $I_{-}$ satisfies the same bounds as $I_{+}$, which finishes the proof adding up all terms. ∎
A result similar to Proposition 2.3 can be derived to quantify the error between functions of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$. To do so, we note that since $\mathsf{A}_{\mathcal{M}_{h}}$ is continuous and coercive, the associated operator $\mathsf{L}_{h}$ is also sectorial with some angle in $(0,\pi/2)$. Hence, without loss of generality, the angle $\theta\in(0,\pi/2)$ can be assumed to be large enough to ensure that Equation 20 also holds for $\mathsf{L}_{h}$, i.e. that for any $z\in G_{\theta}$,
(23) $\|(z-\mathsf{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M}_{h})}$ $\leq C_{\theta}|z|^{-1}\|v_{h}\|_{L^{2}(\mathcal{M}_{h})},\quad v_{h}\in S_{h},$
and that an integral representation similar to Equation 21 also holds for $\gamma(\mathsf{L}_{h})$, namely:
$\gamma(\mathsf{L}_{h})=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)(z-\mathsf{L}_{h})^{-1}\,\mathrm{d}z.$
Proposition 2.5.
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. There exists a constant $C>0$ such that, for any $\tilde{f}\in S_{h}^{\ell}$,
(24) $\left\|\left(\gamma(\mathcal{L}_{h})\tilde{f}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\right\|_{L^{2}(\mathcal{M}_{h})}\leq Ch^{2}\,\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})}.$
This proposition can be seen as an extension of Bonito et al. (2024, Lemma 4.4) relying on extensions of the estimates proven in Bonito et al. (2024, Lemma A.1). Its proof is similar and therefore postponed to Appendix B.
3. Stochastic theory: random fields on surfaces and convergence of SFEM approximation
In this section, we introduce random fields and white noise on surfaces, thus allowing us to make sense of Equation 1 in Section 1. Further, we give approximation methods based on SFEM and prove strong error bounds.
3.1. Random fields on surfaces
Let $(\Omega,\mathcal{S},\amsmathbb{P})$ be a complete probability space. We are interested in Gaussian random fields on $\mathcal{M}$ defined as $L^{2}(\mathcal{M})$-valued random variables through expansions of the form
(25) $\mathcal{Z}=\sum_{i\in\amsmathbb{N}}Z_{i}e_{i},$
where $\{e_{i}\}_{i\in\amsmathbb{N}}$ denotes a real-valued orthonormal basis of $L^{2}(\mathcal{M})$ composed of eigenfunctions of the operator $\mathcal{L}$ (cf. Proposition 2.1), and $\{Z_{i}\}_{i\in\amsmathbb{N}}$ is a sequence of real Gaussian random variables such that $\operatorname{\amsmathbb{E}}[Z_{i}]=0$ for any $i\in\amsmathbb{N}$ and $\sum_{i\in\amsmathbb{N}}\operatorname{\amsmathbb{E}}[|Z_{i}|^{2}]<\infty$. As such, $\mathcal{Z}$ can be seen as en element of the Hilbert space $L^{2}(\Omega;L^{2}(\mathcal{M}))$ of $L^{2}(\mathcal{M})$-valued random variables, to which we associate the inner product $(\cdot,\cdot)_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ (and norm $\|\cdot\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$) defined by
$(\mathcal{Z},\mathcal{Z}^{\prime})_{L^{2}(\Omega;L^{2}(\mathcal{M}))}=\operatorname{\amsmathbb{E}}\left[(\mathcal{Z},\mathcal{Z}^{\prime})_{L^{2}(\mathcal{M})}\right],\quad\mathcal{Z},\mathcal{Z}^{\prime}\in L^{2}(\Omega;L^{2}(\mathcal{M})).$
Finally, in analogy to Equation 25, we formally define the Gaussian white noise on $\mathcal{M}$ by the expansion
(26) $\mathcal{W}=\sum_{i\in\amsmathbb{N}}W_{i}e_{i},$
where $\{W_{i}\}_{i\in\amsmathbb{N}}$ is a sequence of independent real standard Gaussian random variables. We observe that even though this expansion does not converge in $L^{2}(\Omega;L^{2}(\mathcal{M}))$, it does however converge in $L^{2}(\mathcal{M};H^{s}(\mathcal{M}))$ for $s<-d/2$. Moreover, we have that for any $\phi\in L^{2}(\mathcal{M})$, the expansion $(\phi,\mathcal{W})_{L^{2}(\mathcal{M})}=\sum_{i\in\amsmathbb{N}}W_{i}(\phi,e_{i})_{L^{2}(\mathcal{M})}$ converges in $L^{2}(\Omega)$. Further, $(\mathcal{W},\phi)_{L^{2}(\mathcal{M})}$ defines a complex Gaussian variable with mean 00, and for any $\phi^{\prime}\in L^{2}(\mathcal{M})$, $\operatorname{\mathsf{Cov}}((\phi,\mathcal{W})_{L^{2}(\mathcal{M})},(\phi^{\prime},\mathcal{W})_{L^{2}(\mathcal{M})})=\operatorname{\amsmathbb{E}}[(\phi,\mathcal{W})_{L^{2}(\mathcal{M})}\overline{(\phi^{\prime},\mathcal{W})_{L^{2}(\mathcal{M})}}]=(\phi,\phi^{\prime})_{L^{2}(\mathcal{M})}$. As such, the Gaussian white noise (26) can be interpreted as a generalized Gaussian random field over $L^{2}(\mathcal{M})$.
Circling back to the class of random fields defined in Section 1, we can now make sense of Equation 1 through Equation 8 and Equation 26, thus yielding the definition
(27) $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W}=\sum_{i\in\amsmathbb{N}}\gamma(\lambda_{i})W_{i}e_{i},$
which results in $\mathcal{Z}\in L^{2}(\Omega;H^{s}(\mathcal{M}))$ for any $s\geq 0$ such that $4\alpha-d>2s$. Note in particular that all summands in Equation 27 are real-valued functions, and that therefore $\mathcal{Z}$ is real-valued. In Figure 4 we illustrate the influence of the parameter choices on the resulting field on $\amsmathbb{S}^{2}$ for generalized non-stationary Whittle–Matérn fields on $\amsmathbb{S}^{2}$
$\mathcal{Z}=(\mathcal{L})^{-\alpha}\mathcal{W},$
where $\alpha>d/2$. By selecting $D_{ij}(x)=\delta_{ij}$ and $V(x)=\kappa^{2}$ with $\kappa>0$ one recovers the classical, stationary Whittle–Matérn fields studied in various settings in for instance Cox and Kirchner (2020); Bolin et al. (2020); Bonito et al. (2024); Lindgren et al. (2011); Jansson et al. (2022); Whittle (1963). In Figures 4(a) and 4(b) we show this case with $\kappa^{2}=10$ for a rougher field with $\alpha=0.55$ and a smoother one with $\alpha=1.5$, respectively.
Two non-stationary fields are shown in Figure 4(c) and Figure 4(d), obtained by varying the coefficient functions $D_{ij}$ and $V$ and setting $\alpha=0.75$. In Figure 4(c), we keep $D_{ij}=\delta_{ij}$ but use
$V(x)=\begin{cases}10^{5}&\text{for }x_{2}^{6}+x_{1}^{3}-x_{3}^{2}\in(0.1,0.5),\\ 10&\text{else}\end{cases}$
resulting in the observed localized behavior, where the field is essentially turned off in the region with large $V$. More specifically, $V$ describes the local correlation length, where a large $V(x)$ corresponds to a small correlation length around $x$.
Finally, setting $V=10$ constant again, we show the influence of varying parameters $\mathcal{D}$ in Figure 4(d). To derive suitable coefficients, we select a smooth function $f$ and compute its gradient $\nabla_{\amsmathbb{S}^{2}}f$ as well as its skew-gradient $X_{f}$ given by $x\times\nabla f(x)$ at each point $x\in\amsmathbb{S}^{2}$. We set for fixed $\rho_{1},\rho_{2}>0$, $\mathcal{D}(x)u=\rho_{1}(\nabla f(x)\cdot u)\nabla f(x)+\rho_{2}(X_{f}(x)\cdot u)X_{f}(x)$ for any $u\in T_{x}\amsmathbb{S}^{2}$. Here, the inner product refers to the Riemannian inner product associated with the standard round metric on $\amsmathbb{S}^{2}$. Since $\nabla_{\amsmathbb{S}^{2}}f(x)$ and $X_{f}(x)$ both are in $T_{x}\amsmathbb{S}^{2}$, $\mathcal{D}(x)$ is a linear mapping from $T_{x}\amsmathbb{S}^{2}$ into itself. Further, as $\nabla_{\amsmathbb{S}^{2}}f$ is perpendicular to $X_{f}(x)$, by selecting $\rho_{1}$ and $\rho_{2}$, we obtain a field that is elongated either orthogonally to the level sets of $f$ (large $\rho_{1}$, small $\rho_{2}$) or tangentially to the level sets (small $\rho_{1}$, and large $\rho_{2}$). To generate Figure 4(d), we selected $f(x)=x_{2}$, i.e., the function returning the second coordinate in Cartesian coordinates, $\rho_{1}=1$ and $\rho_{2}=25$.
(a) Stationary Whittle–Matérn field with low decay parameter.
(b) Stationary Whittle–Matérn field with high decay parameter.
(c) Non-stationary Whittle–Matérn field by varying potential.
(d) Non-stationary Whittle–Matérn field by varying diffusion matrix.
Figure 4. Some examples to highlight the influence of the different model parameters.
3.2. Approximation of random fields with surface finite elements
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. Following the approach presented by Lang and Pereira (2023), the field $\mathcal{Z}$ is approximated by an expansion similar to that of Equation 27, but involving only quantities defined on the polyhedral surface $\mathcal{M}_{h}$. More precisely, we define the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ of the field $\mathcal{Z}$ by the relation
(28) $\mathsf{Z}_{h}=\sum_{i=1}^{N_{h}}\gamma(\Lambda_{i}^{h})W_{i}^{h}E_{i}^{h},$
where $\{W_{i}^{h}\}_{1\leq i\leq N_{h}}$ is a sequence of independent standard Gaussian random variables, whose precise definition is clarified later in this section, and $\{(\Lambda_{i}^{h},E_{i}^{h})\}_{1\leq i\leq N_{h}}$ are the eigenpairs of $\mathsf{L}_{h}$ introduced in Section 2.2. Then, as proven in Lang and Pereira (2023, Theorem 3.4), $\mathsf{Z}_{h}$ can be decomposed in the nodal basis $\{\psi_{i}\}_{1\leq i\leq N_{h}}$ of $S_{h}$ as
$\mathsf{Z}_{h}=\sum_{i=1}^{N_{h}}Z_{i}\psi_{i},$
where the weights $(Z_{1},\dots,Z_{N_{h}})$ form a centered Gaussian vector which covariance matrix $\bm{\Sigma}_{Z}$ can be expressed using the matrices $\bm{C}$ and $\bm{S}$ introduced in Equations 14 and 15 as follows:
(29) $\bm{\Sigma}_{Z}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}\,\gamma(\bm{S})^{2}\,\big{(}\sqrt{\bm{C}}\big{)}^{-1},$
and, following Equation 16, the function of matrix $\gamma(\bm{S})$ is defined as
(30) $\gamma(\bm{S})^{2}=\bm{V}\operatorname{Diag}\left(\gamma\big{(}\Lambda_{1}^{h}\big{)}^{2},\dots,\gamma\big{(}\Lambda_{1}^{N_{h}}\big{)}^{2}\right)\bm{V}^{T}.$
Note that sampling the weights $(Z_{1},\dots,Z_{N_{h}})$, and therefore the field $\mathsf{Z}_{h}$, using directly the expression of their covariance matrix requires in practice to fully diagonalize $\bm{S}$ (since Equation 29 involves a function of a matrix). Such an operation would result in a prohibitive computational cost (of order $\mathcal{O}(N_{h}^{3})$ operations). To avoid this cost, we use the Chebyshev trick proposed by Lang and Pereira (2023, Section 4), and approximate $\mathsf{Z}_{h}$ by the field $\mathsf{Z}_{h,M}$ defined by
(31) $\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}P_{\gamma,M}(\Lambda_{i}^{h})W_{i}^{h}E_{i}^{h},$
where $P_{\gamma,M}$ is a Chebyshev polynomial approximation of degree $M\in\amsmathbb{N}$ of $\gamma$ over an interval $[\lambda_{\min},\lambda_{\max}]$ containing all the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ of $\mathsf{L}_{h}$ (which we recall, coincide with the eigenvalues of $\bm{S}$). Such an interval can be obtained as follows. On the one hand, one can take $\lambda_{\min}=V_{-}$. On the other hand, following Lang and Pereira (2023), a candidate for $\lambda_{\max}$ is obtained by applying the Gershgorin circle theorem to $\bm{S}$.
Since $\mathsf{Z}_{h,M}$ is defined by just replacing the power spectral density $\gamma$ by the polynomial  $P_{\gamma,M}$, its expansion $\mathsf{Z}_{h,M}$ into the nodal basis,
$\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}Z_{i}^{(M)}\psi_{i},$
is such that the weights $(Z_{1}^{(M)},\dots,Z_{N_{h}}^{(M)})$ now form a centered Gaussian vector with covariance matrix $\bm{\Sigma}_{Z^{(M)}}$ given by
(32) $\bm{\Sigma}_{Z^{(M)}}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}\,P_{\gamma,M}^{2}(\bm{S})\,\big{(}\sqrt{\bm{C}}\big{)}^{-1}.$
Hence, the matrix function in Equation 29 is now replaced by a matrix polynomial $P_{\gamma,M}^{2}(\bm{S})$. This eliminates the eigendecomposition need associated with matrix functions and therefore speeds up the computations. Indeed, the weights $(Z_{1}^{(M)},\dots,Z_{N_{h}}^{(M)})$ can be sampled through
$\begin{pmatrix}Z_{1}^{(M)}\\ \vdots\\ Z_{N_{h}}^{(M)}\end{pmatrix}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}P_{\gamma,M}(\bm{S})\begin{pmatrix}\varepsilon_{1}\\ \vdots\\ \varepsilon_{N_{h}}\end{pmatrix}$
where $\varepsilon_{1},\dots,\varepsilon_{N_{h}}$ are independent standard Gaussian random variables, and the matrix-vector product by $P_{\gamma,M}(\bm{S})$ can be computed iteratively while just requiring products between $\bm{S}$ and vectors. In the next two subsections, we provide error estimates quantifying the error between our target random field $\mathcal{Z}$ and its successive SFEM and polynomial approximations by $\mathsf{Z}_{h}$ and $\mathsf{Z}_{h,M}$.
3.3. Error analysis of the SFEM discretization
We start with analyzing the error between the random field $\mathcal{Z}$ defined on $\mathcal{M}$ and its SFEM approximation $\mathsf{Z}_{h}$, as stated in the next theorem.
Theorem 3.1.
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. Then, there exists $h_{0}>0$, such that for any $h<h_{0}$ the strong approximation error of the random field $\mathcal{Z}$ by its discretization $\mathsf{Z}_{h}^{\ell}$ satisfies the bound
(33) $\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
where $K_{\alpha}(h)=1$ if $\alpha-d/4<2$, $K_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$, and $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ otherwise.
To prove the strong error estimate, we rely on the deterministic error bounds proven in the previous section, and on several intermediate approximations of $\mathcal{Z}$ defined on the spaces $S_{h}^{\ell}$ (i.e., on $\mathcal{M}$) and $S_{h}$ (i.e., on $\mathcal{M}_{h}$). These intermediate approximations require in turn to define approximations of the Gaussian white noise $\mathcal{W}$ on the spaces $S_{h}^{\ell}$ and $S_{h}$.
We first define on $S_{h}^{\ell}$ (i.e., on $\mathcal{M}$), the projected white noise $\widetilde{\mathcal{W}}$ as
(34) $\widetilde{\mathcal{W}}=\sum_{j=1}^{N_{h}}\xi_{j}e_{j}^{h},$
where we recall that $\{e_{j}^{h}\}_{1\leq i\leq N_{h}}$ denotes an orthonormal basis of eigenfunctions of $\mathcal{L}_{h}$, and for any $j\in\{1,\dots,N_{h}\}$, we take $\xi_{j}=\big{(}e_{j}^{h},\mathcal{W}\big{)}_{L^{2}(\mathcal{M})}$. In particular, this last relation implies (by definition of the white noise $\mathcal{W}$) that $\xi_{1},\dots,\xi_{N_{h}}$ are independent standard Gaussian random variables.
Remark 3.2.
By injecting the representation (26) of $\mathcal{W}$ in the definition of $\xi_{j}$, we get that $\widetilde{\mathcal{W}}$ can itself be formally represented by $\widetilde{\mathcal{W}}=\sum_{k=1}^{\infty}W_{k}P_{h}e_{k}=P_{h}\mathcal{W}$. This explains why we refer to it as a projected white noise.
Based on $\widetilde{\mathcal{W}}$, we can then introduce a first approximation, on the space $S_{h}^{\ell}$, of the field $\mathcal{Z}$. We denote this approximation by $\tilde{\mathcal{Z}}_{h}$ and define it in analogy to Equation 27 as
(35) $\tilde{\mathcal{Z}}_{h}$ $=\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}=\sum_{j=1}^{N_{h}}\gamma(\lambda_{j}^{h})\xi_{j}e_{j}^{h}.$
Now, on the space $S_{h}$, we define two white noise approximations $\widehat{\mathsf{W}}$ and $\mathsf{W}$ which are based on the projected white noise $\widetilde{\mathcal{W}}$:
$\widehat{\mathsf{W}}=\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\quad\text{and}\quad\mathsf{W}=\mathsf{P}_{h}(\sigma^{1/2}\widetilde{\mathcal{W}}^{-\ell}),$
where $\sigma$ is the ratio of area measures introduced in Section 2.2. On the one hand, we associate to $\widehat{\mathsf{W}}$ an approximation $\widehat{\mathsf{Z}}_{h}$ of the field $\mathcal{Z}$ on $S_{h}$, which we define in analogy to Equation 36 as
$\widehat{\mathsf{Z}}_{h}=\gamma(\mathsf{L}_{h})\widehat{\mathsf{W}}.$
On the other hand, by expanding $\widehat{\mathsf{W}}$ and $\mathsf{W}$ in the orthonormal basis $\{E_{j}^{h}\}_{1\leq i\leq N_{h}}$ of eigenfunctions of $\mathsf{L}_{h}$, we obtain alternative representations of these fields, and we can draw a link between $\mathsf{W}$ and the SFEM–Galerkin approximation $\mathsf{Z}_{h}$, as stated in the next lemma.
Lemma 3.3.
The noises $\widehat{\mathsf{W}}$ and $\mathsf{W}$ can be written as
$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h}\quad\text{and}\quad\mathsf{W}=\sum_{j=1}^{N_{h}}\beta_{j}E_{j}^{h},$
where $(\alpha_{1},\ldots,\alpha_{N_{h}})$ and $(\beta_{1},\ldots,\beta_{N_{h}})$ are multivariate normal with mean 00 and respective covariance matrices $\bm{A}=[(\sigma E_{i}^{h},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}]_{1\leq i,j\leq N_{h}}$ and $\bm{B}=\bm{I}$. In particular, it holds that the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ defined in Equation 28 satisfies
(36) $\mathsf{Z}_{h}=\gamma(\mathsf{L}_{h})\mathsf{W}$
where we take for any $j\in\{1,\dots,N_{h}\}$, $W_{j}^{h}=\beta_{j}$.
Proof.
As $\widehat{\mathsf{W}}\in S_{h}$, we can expand it in the orthonormal basis $\{E_{j}^{h}\}_{1\leq i\leq N_{h}}$ to get
$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h}$
where $\alpha_{j}=(\widehat{\mathsf{W}},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$. Then, by definition of $\sigma$, of $\widetilde{\mathcal{W}}$ and since $(E_{j}^{h})^{\ell}\in S_{h}^{\ell}$, we further get for any $1\leq j\leq N_{h}$,
$\alpha_{j}$ $=(\widetilde{\mathcal{W}},(E_{j}^{h})^{\ell})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{N_{h}}(\mathcal{W},e_{k}^{h})_{L^{2}(\mathcal{M})}(e_{k}^{h},(E_{j}^{h})^{\ell})_{L^{2}(\mathcal{M})}$
$=\bigg{(}\mathcal{W},\sum_{k=1}^{N_{h}}(e_{k}^{h},(E_{j}^{h})^{\ell})_{L^{2}(\mathcal{M})}e_{k}^{h}\bigg{)}_{L^{2}(\mathcal{M})}=(\mathcal{W},(E_{j}^{h})^{\ell})_{L^{2}(\mathcal{M})}.$
Therefore, by definition of the white noise $\mathcal{W}$, we can conclude that for any $1\leq i,j\leq N_{h}$, $\alpha_{j}$ is normally distributed with mean 00, and that
$\operatorname{\amsmathbb{E}}[\alpha_{i}\alpha_{j}]=\operatorname{\mathsf{Cov}}(\alpha_{i},\alpha_{j})=\left((E_{i}^{h})^{\ell},(E_{j}^{h})^{\ell}\right)_{L^{2}(\mathcal{M})}=\left(\sigma E_{i}^{h},E_{j}^{h}\right)_{L^{2}(\mathcal{M}_{h})}=A_{ij}.$
Hence, $(\alpha_{1},\ldots,\alpha_{N_{h}})$ is indeed multivariate normal (any linear combination of the $\alpha_{j}$ being Gaussian by definition of $\mathcal{W}$) with mean 00 and covariance matrix $\bm{A}$.
Similarly, since $\mathsf{W}\in S_{h}$, we can write again
$\mathsf{W}=\sum_{j=1}^{N_{h}}\beta_{j}E_{j}^{h},$
where $\beta_{j}=(\mathsf{W},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma^{1/2}\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},\sigma^{-1/2}E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$, and the same computations as before yield that $\beta_{j}=\big{(}\mathcal{W},(\sigma^{-1/2}E_{j}^{h})^{\ell}\big{)}_{L^{2}(\mathcal{M})}$. Hence, we can conclude this time that for any $1\leq i,j\leq N_{h}$, $\beta_{j}$ is also normally distributed with mean 00, and that
$\operatorname{\amsmathbb{E}}[\beta_{i}\beta_{j}]$ $=\left(\sigma(\sigma^{-1/2}E_{i}^{h}),(\sigma^{-1/2}E_{j}^{h})\right)_{L^{2}(\mathcal{M}_{h})}=\left(\sigma(\sigma^{-1/2}E_{i}^{h}),(\sigma^{-1/2}E_{j}^{h})\right)_{L^{2}(\mathcal{M}_{h})}=B_{ij},$
by orthonormality of $\{E_{j}^{h}\}_{1\leq i\leq N_{h}}$. In conclusion, $(\beta_{1},\ldots,\beta_{N_{h}})$ is indeed multivariate normal with mean 00 and covariance matrix $\bm{B}=\bm{I}$. In particular, this means that $\beta_{1},\dots,\beta_{N_{h}}$ are independent standard Gaussian variables.
Finally, note that we can write
$\gamma(\mathsf{L}_{h})\mathsf{P}_{h}\mathsf{W}$ $=\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})\big{(}\mathsf{P}_{h}\mathsf{W},E_{j}^{h}\big{)}_{L^{2}(\mathcal{M}_{h})}E_{j}^{h}=\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})\big{(}\mathsf{W},E_{j}^{h}\big{)}_{L^{2}(\mathcal{M}_{h})}E_{j}^{h}$
$=\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})\beta_{j}E_{j}^{h}=\mathsf{Z}_{h}$
where we take $W_{j}^{h}=\beta_{j}$ in Equation 28. ∎
We now circle back to proving Theorem 3.1. Using the intermediate approximations $\tilde{\mathcal{Z}}_{h}$ and $\widehat{\mathsf{Z}}_{h}$ and the equivalence of the $L^{2}$ norms on $\mathcal{M}$ and $\mathcal{M}_{h}$, we can upper-bound the error between the field $\mathcal{Z}$ and its SFEM–Galerkin approximation $\mathsf{Z}_{h}$ by
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}$ $\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\tilde{\mathcal{Z}}_{h}-\widehat{\mathsf{Z}}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\widehat{\mathsf{Z}}_{h}^{\ell}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}{-}\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}+\big{\|}\widehat{\mathsf{Z}}_{h}-\mathsf{Z}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
We derive error estimates for each one of the three terms obtained in the last inequality. We start with the term $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$.
Lemma 3.4.
It holds that
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
where $C_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$ and $C_{\alpha}(h)=1$, otherwise.
Proof.
Our aim is to bound the error between $\tilde{\mathcal{Z}}_{h}$ and $\mathcal{Z}$, where we remark in particular that $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W}$ and $\mathcal{Z}_{h}=\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}$ (cf. Remark 3.2). Let $\eta=\alpha+d/4>d/2$, and let $\zeta=\alpha-d/4>0$, so that $\alpha=(\eta+\zeta)/2$. Further, define the function $\tilde{\gamma}(x)=\gamma(x)x^{\eta/2}$. Note that, since $\gamma$ is an $\alpha$-power spectral density and since $\tilde{\gamma}$ decays as
$|\tilde{\gamma}(x)|\lesssim|x|^{-\left(\alpha-\eta/2\right)}=|x|^{-\zeta/2},$
$\tilde{\gamma}$ is a $(\zeta/2)$-power spectral density. Now, by definition of $\tilde{\gamma}$,
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $=\|\tilde{\gamma}(\mathcal{L})\mathcal{L}^{-\eta/2}\mathcal{W}-\tilde{\gamma}(\mathcal{L}_{h})\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
so that the triangle inequality yields
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq E_{1}+E_{2},$
where we take
$E_{1}$ $=\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
$E_{2}$ $=\|\tilde{\gamma}(\mathcal{L}_{h})\big{(}P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}-\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\big{)}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
For the term $E_{1}$, note that by Bonito et al. (2024, Lemma 4.2),
$\|\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq\frac{\eta}{\eta-d/2}.$
In particular, $\mathcal{L}^{-\eta/2}\mathcal{W}$ is in $L^{2}(\Omega;L^{2}(\mathcal{M}))$ and an application of Proposition 2.3 yields that
$E_{1}\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}},$
where
$C_{\zeta}(h)=\begin{cases}1,&\zeta\neq 2,\\ |\log h|,&\zeta=2.\\ \end{cases}$
For $E_{2}$, we first write $P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}=\sum_{i=1}^{N_{h}}\beta_{i}e_{i}^{h}$, where $\beta_{i}=(P_{h}\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=(\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}W_{k}(e_{k},e_{i}^{h})_{L^{2}(\mathcal{M})}$. Further, following the definition of $\tilde{\mathcal{W}}$ in Equation 34, we have $\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\eta/2}\xi_{i}e_{i}^{h}$ where $\xi_{i}=(e_{i}^{h},\mathcal{W})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}W_{k}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}$. Note in particular that both $\beta_{i}$ and $\xi_{i}$ are real-valued (since the eigenbases $\{e_{k}\}_{k\in\amsmathbb{N}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are real-valued). Thus,
$E_{2}^{2}=\sum_{i=1}^{N_{h}}|\tilde{\gamma}(\lambda_{i}^{h})|^{2}\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}],$
where we use the orthonormality of the eigenfunctions. Note that
$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]-2\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]+\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}].$
First, by the independence of the sequence $\{W_{k}\}_{k=1}^{\infty}$ and $(\lambda_{k},e_{k})$ eigenpairs of $\mathcal{L}$,
$\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]$ $=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}|(e_{k},e_{i}^{h})|^{2}_{L^{2}(\mathcal{M})}=\left(\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}e_{k},e_{i}^{h}\right)_{L^{2}(\mathcal{M})}$
$=(\mathcal{L}^{-\eta}e_{i}^{h},e_{i}^{h})_{L^{2}(\mathcal{M})}=\|\mathcal{L}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}.$
Similarly, and using the fact that $(\lambda_{i}^{h},e_{i}^{h})$ is an eigenpair of $\mathcal{L}_{h}$,
$\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}]=(\lambda_{i}^{h})^{-\eta}\sum_{k=1}^{\infty}|(e_{i}^{h},e_{k})|^{2}_{L^{2}(\mathcal{M})}=(\lambda_{i}^{h})^{-\eta}\|e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}=\|\mathcal{L}_{h}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}$
and
$\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]=(\lambda_{i}^{h})^{-\eta/2}\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}|(e_{k},e_{i}^{h})|^{2}=(\mathcal{L}^{-\eta/2}e_{i}^{h},\mathcal{L}^{-\eta/2}_{h}e_{i}^{h})_{L^{2}(\mathcal{M})}.$
This implies that by Proposition 2.3,
$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\|\mathcal{L}^{-\eta/2}e_{i}^{h}-\mathcal{L}^{-\eta/2}_{h}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2},$
where
$C_{\eta}(h)=\begin{cases}1,&\eta\neq 2,\\ |\log h|,&\eta=2.\end{cases}$
Hence, it holds that
$E_{2}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}.$
It remains to bound $\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$. To this end, recall Equation 12 and Equation 6 which yield that
$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\zeta}\lesssim\sum_{i=1}^{N_{h}}i^{-2\zeta/d}\lesssim\begin{cases}N_{h}^{1-2\zeta/d},&2\zeta/d<1,\\ \log(N_{h}),&2\zeta/d=1,\\ 1,&2\zeta/d>1,\end{cases}\lesssim\begin{cases}h^{-2(d/2-\zeta)},&\zeta<d/2,\\ |\log(h)|,&\zeta=d/2,\\ 1,&\zeta>d/2,\end{cases}$
where we used in the last inequality that $N_{h}\propto h^{-d}$. We obtain
$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim h^{2(\zeta-\max\{\zeta;d/2\})}K_{\zeta}(h)^{2}$
with
$K_{\zeta}(h)=\begin{cases}1,&\zeta\neq d/2\\ |\log(h)|^{1/2},&\zeta=d/2\end{cases}$
We conclude that
$E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}}$
and in particular,
$h^{-\min\{\zeta;2\}}E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\beta},$
where we take $\beta=\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}-\min\{\zeta;2\}$. We have two cases. On the one hand, if $\zeta\geq 2$ then, since $\eta>\zeta$, we also have $\eta>2$ and therefore $\beta=0$, $C_{\eta}(h)=1$ and $K_{\zeta}(h)=1$ (given that $d/2<2$). On the other hand, if $\zeta<2$, then $\max\{\zeta;d/2\}<2$ and $\max\{\zeta;d/2\}<\eta$ and therefore $\beta=\min\{\eta;2\}-\max\{\zeta;d/2\}>0$. In turn, this means that $C_{\eta}(h)K_{\zeta}(h)h^{\beta}\lesssim|\log(h)|^{3/2}|h^{\beta}\lesssim 1$. Hence, in both cases
$h^{-\min\{\zeta;2\}}E_{2}\lesssim 1,$
and we conclude that
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\lesssim h^{\min\{\zeta;2\}}\left(C_{\zeta}(h)+h^{-\min\{\zeta;2\}}E_{2}\right)\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}}.$
The result follows by inserting the definition of $\zeta$. ∎
For the error between $\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}$ and $\widehat{\mathsf{Z}}_{h}$ we get the following estimate, inspired by Bonito et al. (2024, Lemma 4.4).
Lemma 3.5.
It holds that
$\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\lesssim|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}.$
Proof.
Let $\mathcal{E}_{h}=\big{\|}\left(\tilde{\mathcal{Z}}_{h}\right)^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}=\big{\|}\left(\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$. We note that $\widetilde{\mathcal{W}}$ is an $S_{h}^{\ell}$-valued random variable. Therefore, we can apply Equation 24 to realizations of $\widetilde{\mathcal{W}}$, and take the expectation on both sides to get
$\mathcal{E}_{h}$ $\lesssim h^{2}\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
We then distinguish two cases. First, if $d=1$, then we note that $\min\{\alpha+d/4;1\}\geq\alpha+d/4>d/2$ since $\alpha>d/4$. Besides, Bonito et al. (2024, Lemma 4.2) yields that for all $r\in(d/2,2)$, there is a constant $C>0$ such that
(37) $\|\mathcal{L}_{h}^{-r/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\leq C\frac{r}{r-d/2}.$
Hence, using estimate (37), we conclude that
$\mathcal{E}_{h}$ $\lesssim h^{2}\frac{\min\{\alpha+d/4;1\}}{\min\{\alpha+d/4;1\}-d/2}\lesssim h^{2}.$
If now $d=2$, since $\alpha+d/4>d/2=1$, we have $\min\{\alpha+d/4;1\}=1$. We then note that by the proof of Bonito et al. (2024, Lemma 4.2), for any $\varepsilon>0$
$\|\mathcal{L}_{h}^{-1/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\lesssim\sum_{j=1}^{N_{h}}\lambda_{j}^{-1}\lesssim\lambda_{N_{h}}^{\varepsilon}\sum_{j=1}^{N_{h}}\lambda_{j}^{-1-\varepsilon}.$
Apply now Equation 6 to see that
$\|\mathcal{L}_{h}^{-1/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\lesssim N_{h}^{\varepsilon}\sum_{j=1}^{N_{h}}j^{-(1+\varepsilon)}\lesssim N_{h}^{\varepsilon}\frac{1+\varepsilon}{\varepsilon}\lesssim h^{-2\varepsilon}\frac{1+\varepsilon}{\varepsilon},$
meaning that for any $\varepsilon>0$,
$\mathcal{E}_{h}\lesssim\bigg{(}\frac{1+\varepsilon}{\varepsilon}\bigg{)}^{1/2}h^{2-\varepsilon}.$
In particular, taking $\varepsilon=|\log(h)|^{-1}$ yields
$\mathcal{E}_{h}\lesssim|\log(h)|^{1/2}h^{2},$
which concludes the proof. ∎
Finally, for the error between $\widehat{\mathsf{Z}}_{h}$ and $\mathsf{Z}_{h}$ we get the following estimate.
Lemma 3.6.
It holds that there is a constant $C>0$ such that
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq Ch^{2}.$
Proof.
To prove this statement, we note that, using the same notations as the ones in the proof of Lemma 3.3,
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$ $=\left\|\gamma(\mathsf{L}_{h})\left(\widehat{\mathsf{W}}-\mathsf{W}\right)\right\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$
$=\operatorname{\amsmathbb{E}}\left[\left\|\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})(\alpha_{j}-\beta_{j})E_{j}^{h}\right\|_{L^{2}(\mathcal{M}_{h})}^{2}\right]=\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2}\operatorname{\amsmathbb{E}}[|\alpha_{j}-\beta_{j}|^{2}],$
where we applied the orthogonality of the eigenfunctions in the last step. Now, following the definition of $\alpha_{j}$ and $\beta_{j}$ given in the proof of Lemma 3.3,
$\operatorname{\amsmathbb{E}}[|\alpha_{j}-\beta_{j}|^{2}]$ $=\operatorname{\amsmathbb{E}}\bigg{[}\bigg{|}\big{(}\mathcal{W},\big{(}(1-\sigma^{-1/2})E_{j}^{h}\big{)}^{\ell}\big{)}_{L^{2}(\mathcal{M})}\bigg{|}^{2}\bigg{]}=\left\|\left((1-\sigma^{-1/2})E_{j}^{h}\right)^{\ell}\right\|^{2}_{L^{2}(\mathcal{M})}$
$=\left\|(\sigma^{1/2}-1)E_{j}^{h}\right\|^{2}_{L^{2}(\mathcal{M}_{h})}\leq\left\|(\sigma^{1/2}-1)\right\|^{2}_{L^{\infty}(\mathcal{M}_{h})}.$
Besides, for all $x\in\mathcal{M}_{h}$, it holds that
$\sigma^{1/2}(x)-1=\sqrt{1+\sigma(x)-1}-1\leq\sqrt{1+|\sigma(x)-1|}-1\leq 1+\frac{1}{2}|\sigma(x)-1|-1=\frac{1}{2}|\sigma(x)-1|.$
Therefore,
$\left\|(\sigma^{1/2}-1)\right\|^{2}_{L^{\infty}(\mathcal{M}_{h})}\leq\left\|(\sigma-1)\right\|^{2}_{L^{\infty}(\mathcal{M}_{h})}\leq Ch^{2},$
where Dziuk and Elliott (2013, Lemma 4.1) was applied in the final step. We conclude that
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}\leq Ch^{4}\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2},$
and it remains to show that $\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2}$ is bounded by a constant. To this end, we use Lemma 2.2, which implies that there exists some constant $C_{\lambda}>0$ such that
$|\Lambda_{j}^{h}/\lambda_{j}^{h}-1|\leq C_{\lambda}h^{2}.$
Recall then that the mesh size $h$ satisfies $h<h_{0}$, where $h_{0}\in(0,1)$. Without loss of generality, let us further assume that $C_{\lambda}h_{0}<1$. Now, by the growth assumption on $\gamma$,
$\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2}$ $\lesssim\sum_{j=1}^{N_{h}}|\Lambda_{j}^{h}|^{-2\alpha}=\sum_{j=1}^{N_{h}}|\lambda_{j}^{h}|^{-2\alpha}\left|1+\left(\frac{\Lambda_{j}^{h}}{\lambda_{j}^{h}}-1\right)\right|^{-2\alpha}\leq(1-Ch_{0}^{2})^{-2\alpha}\sum_{j=1}^{N_{h}}|\lambda_{j}^{h}|^{-2\alpha}$
$\lesssim\sum_{j=1}^{N_{h}}|\lambda_{j}^{h}|^{-2\alpha}.$
Using Equation 12 and Equation 6, we bound
$\sum_{j=1}^{N_{h}}|\lambda_{j}^{h}|^{-2\alpha}\leq\sum_{j=1}^{N_{h}}|\lambda_{j}|^{-2\alpha}\lesssim\sum_{j=1}^{\infty}j^{-4\alpha/d}=\zeta(4\alpha/d),$
where $\zeta$ denotes the Riemann zeta function. Thus, $\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2}$ is bounded by a constant and
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}\leq Ch^{4},$
which proves the lemma. ∎
Equipped with the estimates derived in Lemmas 3.6, 3.5 and 3.4 we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
By summing the three estimates derived in Lemmas 3.6, 3.5 and 3.4, we get
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}+|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}.$
Let then $\epsilon_{h}$ be the quantity defined by $\epsilon_{h}=\big{(}K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}\big{)}^{-1}\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$, which satisfies
$\epsilon_{h}$ $\lesssim C_{\alpha}(h)K_{\alpha}(h)^{-1}+|\log(h)|^{(d-1)/2}K_{\alpha}(h)^{-1}h^{2-\min\{\alpha-d/4;2\}}.$
If $\alpha-d/4\in(0,2)$, we have $C_{\alpha}(h)=K_{\alpha}(h)=1$ and $02-\min\{\alpha-d/4;2\}>$ 0, and therefore $\epsilon_{h}\lesssim 1+$. Similarly, if $\alpha-d/4=2$, it holds $C_{\alpha}(h)=K_{\alpha}(h)=|\log h|$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim 1+|\log(h)|^{(d-3)/2}\lesssim 1$. Finally, for $\alpha-d/4>2$, $C_{\alpha}(h)=1$, $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim|\log h|^{-(d-1)/2}+1\lesssim 1$. In conclusion, we obtain for any $\alpha>d/4$, $\epsilon_{h}\lesssim 1$ which in turn yields
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}.\qed$
Having bounded the SFEM error, we are now ready to derive the additional error associated with the Chebyshev approximation which we use to compute SFEM–Galerkin approximations in practice.
3.4. Error analysis of the Chebyshev approximation
To bound the error between $\mathsf{Z}_{h,M}$ and $\mathsf{Z}_{h}$, we can directly apply Lang and Pereira (2023, Theorem 5.8). We obtain the following result.
Theorem 3.7.
Let $[\lambda_{\min},\lambda_{\max}]\subset(0,\infty)$ be the interval on which the Chebyshev polynomial approximation $P_{\gamma,M}$ is computed, and let $\xi=\lambda_{\min}/(\lambda_{\max}-\lambda_{\min})$. Then, there exists a constant $C>0$ such that the error between the discretized field $\mathsf{Z}_{h}$ and its approximation $\mathsf{Z}_{h,M}$ is upper-bounded by
(38) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C{h}^{-d/2}\epsilon_{\xi}^{-1}(1+\epsilon_{\xi})^{-M}$
with $\epsilon_{\xi}=\xi+\sqrt{\xi(2+\xi)}$. In particular, setting $\lambda_{\min}=V_{-}$ and $\lambda_{\max}=\Lambda_{N_{h}}^{h}$, the error is bounded by
(39) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C_{V}^{-1}{h}^{-(d/2+1)}\exp(-C_{V}h^{-1}M)$
for some constant $C_{V}>0$ proportional to $\sqrt{V_{-}}$.
Proof.
Let $E_{\xi}\subset\amsmathbb{C}$ be the ellipse centered at $z=(\lambda_{\min}+\lambda_{\max})/2$, with foci $z_{1}=\lambda_{\min}$ and $z_{2}=\lambda_{\max}$, and semi-major axis $a_{\xi}=\lambda_{\max}/2$. In particular, note that $E_{\xi}\subset H_{\pi/2}$ and for any $z\in E_{\xi}$, $\operatorname{Re}(z)\geq\lambda_{\min}/2>0$. Hence, since $\gamma$ is a power spectral density, by definition $\gamma$ is holomorphic and bounded inside $E_{\xi}$. We can then adapt the same proof as in Lang and Pereira (2023, Theorem 5.8) to obtain the stated proposition. ∎
For a fixed mesh size $h$, the approximation error converges to 00 as the order of the polynomial approximation $M$ goes to infinity. Choosing $M$ as a function of $h$ that grows fast enough then allows us to ensure the convergence of the approximation error as $n$ goes to infinity (Lang and Pereira, 2023, Section 5.2).
Appendix A Deterministic proofs
A.1. Proof of Proposition 2.1
Proof.
A standard result in the spectral theory of elliptic operators on compact Riemannian manifolds (see e.g. (Shubin, 2001, Section 8)) ensures that there exists a set of eigenpairs $\{(\lambda_{i},f_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ such that $0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{f_{i}\}_{i\in\amsmathbb{N}}$ is an orthonormal basis of $L^{2}(\mathcal{M})$ composed of possibly complex-valued functions.
Hence, let us prove that $\lambda_{1}\geq\delta$ and that we can build an orthonormal basis $\{e_{i}\}_{i\in\amsmathbb{N}}$ of $L^{2}(\mathcal{M})$ such that each $e_{i}$ is real-valued and an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i}$. On the one hand, by definition of $\mathcal{L}$, $\lambda_{1}$ and $f_{1}$, we obtain
$\lambda_{1}=\lambda_{1}(f_{1},f_{1})_{L^{2}(\mathcal{M})}=\mathsf{A}_{\mathcal{M}}(f_{1},f_{1})\geq\delta\|f_{1}\|_{H^{1}(\mathcal{M})}^{2}\geq\delta\|f_{1}\|_{L^{2}(\mathcal{M})}^{2}=\delta,$
where for the last two inequalities we used the coercivity of $\mathsf{A}_{\mathcal{M}}$ and the definition of the $H^{1}$-norm.
On the other hand, let $\lambda>0$ be one of the eigenvalues of $\mathcal{L}$, and $E_{\lambda}\subset L^{2}(\mathcal{M})$ the associated eigenspace. Following again the results from (Shubin, 2001, Section 8), we get that $E_{\lambda}\subset C^{\infty}(\mathcal{M})$, $\dim E_{\lambda}<\infty$ and that if $\lambda^{\prime}\neq\lambda$ is another eigenvalue of $\mathcal{L}$, then $E_{\lambda}$ and $E_{\lambda^{\prime}}$ are orthogonal. Besides, $E_{\lambda}$ is in fact generated by the set $\{f_{j}\}_{j\in J_{\lambda}}$, where $J_{\lambda}=\{i\in\amsmathbb{N}:\lambda_{i}=\lambda\}$ is finite (since $\dim E_{\lambda}<\infty$).
Take then $u\in E_{\lambda}$. Hence, for any $v\in H^{1}(\mathcal{M})$, $\mathsf{A}_{\mathcal{M}}(u,v)=\lambda(u,v)_{L^{2}(\mathcal{M})}$. But also, by definition of $\mathsf{A}_{\mathcal{M}}$,
$\mathsf{A}_{\mathcal{M}}(\overline{u},v)=\mathsf{A}_{\mathcal{M}}(\overline{v},u)=\overline{\mathsf{A}_{\mathcal{M}}(u,\overline{v})}=\overline{\lambda(u,\overline{v})_{L^{2}(\mathcal{M})}}=\lambda(\overline{u},v)_{L^{2}(\mathcal{M})},$
where used the fact that $\mathcal{D}$ is a real symmetric matrix and $V$ is real-valued for the first two equalities. Consequently, we also have $\overline{u}\in E_{\lambda}$, and so, the real-valued functions $\operatorname{Re}(u)=(u+\overline{u}))/2$ and $\operatorname{Im}(u)=(u-\overline{u}))/2i$ (corresponding to real and imaginary parts of $u$) are also in $E_{\lambda}$.
Circling back to the orthonormal basis $\{f_{j}\}_{j\in J_{\lambda}}$ of $E_{\lambda}$, we consider the set of real-valued functions $F_{\lambda}=\{\operatorname{Re}(f_{j})\}_{j\in J_{\lambda}}\cup\{\operatorname{Im}(f_{j})\}_{j\in J_{\lambda}}\subset E_{\lambda}$, and the subspace $V_{\lambda}\subset E_{\lambda}$ generated by $F_{\lambda}$. In particular $\dim V_{\lambda}\leq\dim E_{\lambda}$. By applying the Gram–Schmidt orthogonalization process to $F_{\lambda}$, we get an orthonormal basis $\{e_{k}\}_{1\leq k\leq\dim V_{\lambda}}$ of $V_{\lambda}$ which by construction is composed of real-valued functions (since $F_{\lambda}$ is composed of real-valued functions). Let us show that $\{e_{k}\}_{1\leq k\leq\dim V_{\lambda}}$ is in fact a basis of $E_{\lambda}$, or equivalently that $\dim V_{\lambda}=\dim E_{\lambda}$.
We proceed by contradiction. Assume that $\dim V_{\lambda}<\dim E_{\lambda}$. This means in particular that the orthogonal complement of $V_{\lambda}$ in $E_{\lambda}$, denoted by $V_{\lambda}^{\perp}$, is not reduced to 00. Let then $0\neq w\in V_{\lambda}^{\perp}$. By linearity, we have, for any $j\in J_{\lambda}$, $(f_{j},w)_{L^{2}(\mathcal{M})}=(\operatorname{Re}(f_{j}),w)_{L^{2}(\mathcal{M})}+i(\operatorname{Im}(f_{j}),w)_{L^{2}(\mathcal{M})}=0$, since $\operatorname{Re}(f_{j}),\operatorname{Im}(f_{j})\in F_{\lambda}\subset V_{\lambda}$. Hence, since $w\in E_{\lambda}$ and $\{f_{j}\}_{j\in J_{\lambda}}$ is an orthonormal basis of $E_{\lambda}$, it must hold that $w=0$, which contradicts our initial claim. Consequently, $\dim V_{\lambda}=\dim E_{\lambda}$, and therefore $\{e_{k}\}_{1\leq k\leq\dim V_{\lambda}}$ is an orthonormal basis of $E_{\lambda}$.
Finally, by repeating the construction above to each eigenspace $E_{\lambda}$ associated with distinct eigenvalues, and concatenating the obtained bases, we obtain an orthonormal basis $L^{2}(\mathcal{M})$ (due to the fact that these eigenspaces are orthogonal to one another and span $L^{2}(\mathcal{M})$). Each element in this basis is an eigenfunction of $\mathcal{L}$ since it is built from a given eigenspace, and is a real-valued function. This concludes our proof. ∎
A.2. Proof of Lemma 2.4
Proof.
Let us start by proving the following claims: for any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$,
(40) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2},$
(41) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq\frac{C}{|z|}.$
We omit the proof of Equation 41 and refer the reader to Fujita and Suzuki (1991, Theorem 5.1), where the proof may be adapted verbatim.
To prove Equation 40, we can largely proceed as in Fujita and Suzuki (1991, Theorem 7.1), substituting the Euclidean elliptic regularity estimates and interpolation bounds with their surface counterparts, see for instance Taylor (2011, Chapter 11), Dziuk and Elliott (2013, Lemma 4.3) and Bonito et al. (2024, Equation (4.8)). We must in essence only adapt Fujita and Suzuki (1991, Lemma 7.1) to also hold for $z\in\Gamma_{0}$, as their estimate only holds for $z\in\partial G_{\theta}$. In other words, we must show that for any $z=\delta_{0}e^{it}$, where $t\in[-\theta,\theta]$, there is a constant $\tilde{\delta}>0$ such that for all $v\in H^{1}(\mathcal{M})$,
(42) $|z|\|v\|_{L^{2}(\mathcal{M})}^{2}+\|v\|_{H^{1}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|z\|v\|_{L^{2}(\mathcal{M})}^{2}-\mathsf{A}_{\mathcal{M}}(v,v)\right|.$
To prove this inequality, note that by coercivity of $\mathsf{A}_{\mathcal{M}}$,
$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq\delta\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}.$
Let then $\tilde{\delta}=\frac{\delta_{0}}{1+\delta_{0}}$. Since the only requirement on $\delta_{0}$ was that $\delta_{0}<\delta/2$, we can then write
$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\geq 2\delta_{0}\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-\tilde{\delta}(1+\delta_{0})\|v\|_{L^{2}(\mathcal{M})}^{2}$
$=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}+\left(\delta_{0}\tilde{\delta}-(2\delta_{0}+1)\tilde{\delta}\right)\|v\|_{L^{2}(\mathcal{M})}^{2}$
$\geq 2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-(2\delta_{0}+1)\tilde{\delta}\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\tilde{\delta}\|v\|_{L^{2}(\mathcal{M})}^{2}$
$=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right)=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\right),$
since $|z|=\delta_{0}$ if $z\in\Gamma_{0}$. Therefore,
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left(\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right).$
This last inequality implies in particular that $\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq 0$ and therefore that
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right|.$
Using once again that $|z|=\delta_{0}$ and that $\mathsf{A}_{\mathcal{M}}(v,v)\geq 0$, we deduce that
$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\leq\tilde{\delta}^{-1}\left||\mathsf{A}_{\mathcal{M}}(v,v)|-|z|\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-z\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|$
by the reverse triangle inequality. To conclude, we have proven that the relation (42) holds and equipped with this estimate, we can proceed as in the proof of Fujita and Suzuki (1991, Theorem 7.1) and obtain Equation 40.
Finally, by interpolating between the bounds obtained in Equations 40 and 41, we obtain that, for any $\beta\in[0,1]$,
$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq C\frac{h^{2\beta}}{|z|^{1-\beta}}.\qed$
Appendix B Error between functions of discrete operators
The aim of this section is to prove Equation 24. To do so, we rely on a geometric consistency estimate, and on extensions of the results of Bonito et al. (2024, Lemma A.1), which we start by stating and proving.
B.1. Geometric consistency estimate
The following geometric consistency estimate quantifies the error between the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$. Its proof is a straightforward adaptation of the proof of Dziuk and Elliott (2013, Lemma 4.7) to account for the diffusion matrix $\mathcal{D}$.
Lemma B.1.
There is a constant $C>0$ such that for all $h<h_{0}$ and $u_{h},v_{h}\in S_{h}^{\ell}$,
(43) $\left|\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})\right|\leq Ch^{2}\|u_{h}^{\ell}\|_{H^{1}(\mathcal{M})}\|v_{h}^{\ell}\|_{H^{1}(\mathcal{M})}.$
Proof.
[…]
Let $\Pi=I-\nu\nu^{T}$ (resp. $\Pi_{h}=I-\nu_{h}\nu_{h}^{T}$) be the orthogonal projection onto the tangent planes of $\mathcal{M}$ (resp. $\mathcal{M}_{h}$), and let $\mathcal{H}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ be the extended Weingarten map of $\mathcal{M}$ (cf. Dziuk and Elliott (2013, Definition 2.5)). Recall in particular that $\mathcal{H}(x)\nu(x)=0$ for any $x\in\mathcal{M}$, meaning in particular that $\mathcal{H}\Pi=\Pi\mathcal{H}=\mathcal{H}$. Finally, we introduce the map $\mathcal{Q}_{h}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ defined as
$\mathcal{Q}_{h}=\frac{1}{\sigma^{\ell}}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi,$
where $d_{s}$ is the oriented distance function restricted to $\mathcal{M}_{h}$ and introduced in Section 2.2. On the one hand, note that for any $u_{h},v_{h}\in S_{h}$,
$(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot\nabla_{\mathcal{M}_{h}}\overline{v}_{h}$ $=\big{(}\mathcal{D}^{-\ell}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot\big{(}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell}\big{)}$
$=\sigma\big{(}\mathcal{Q}_{h}^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell},$
which gives, after integrating both sides over $\mathcal{M}_{h}$ and using Equation 9,
(44) $\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}=\int_{\mathcal{M}}\big{(}\mathcal{Q}_{h}\nabla_{\mathcal{M}}u_{h}^{\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A.$
Let then $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}\colon S_{h}^{\ell}\times S_{h}^{\ell}\rightarrow\amsmathbb{R}$ be the Hermitian form defined for any $u_{h}^{\ell},v_{h}^{\ell}\in S_{h}^{\ell}$ by
(45) $\begin{split}&\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\int_{\mathcal{M}}\big{(}\mathcal{Q}_{h}\nabla_{\mathcal{M}}u_{h}^{\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A+\int_{\mathcal{M}}\big{(}\sigma^{\ell}\big{)}^{-1}Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\,\mathrm{d}A.\end{split}$
Note that following Equations 44 and 9, $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}$ satisfies for any $u_{h},v_{h}\in S_{h}$ the equality $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})$. Therefore, for any $u_{h},v_{h}\in S_{h}$, we bound
(46) $\begin{split}\big{|}\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})&-\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})\big{|}=\big{|}\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})\big{|}\\ &\leq\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\right|+\left|\int_{\mathcal{M}}(1-\big{(}\sigma^{\ell}\big{)}^{-1})Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\leavevmode\nobreak\ \,\mathrm{d}A\right|.\end{split}$
We now bound these two terms. Recall that Dziuk and Elliott (2013, Lemma 4.1) shows
(47) $\|\sigma\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim 1,\quad\|\sigma^{-1}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim 1,\quad\|\sigma-1\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2},\quad\|\sigma^{-1}-1\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}.$
Hence, since $V$ takes positive values,
$\big{|}\int_{\mathcal{M}}(1-\big{(}\sigma^{\ell}\big{)}^{-1})Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\leq\int_{\mathcal{M}}|1-\big{(}\sigma^{\ell}\big{)}^{-1}|\leavevmode\nobreak\ V\leavevmode\nobreak\ |u_{h}^{\ell}|\leavevmode\nobreak\ |{v_{h}}^{\ell}|\,\mathrm{d}A$
$\leq\|1-\sigma^{-1}\|_{L^{\infty}(\mathcal{M}_{h})}\int_{\mathcal{M}}V\leavevmode\nobreak\ |u_{h}^{\ell}|\leavevmode\nobreak\ |{v_{h}}^{\ell}|\,\mathrm{d}A,$
which in turn gives (using the Cauchy–Schwartz inequality and Equation 47),
(48) $\big{|}\int_{\mathcal{M}}(1-\big{(}\sigma^{\ell}\big{)}^{-1})Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}V|u_{h}^{\ell}|^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}V\leavevmode\nobreak\ |{v_{h}}^{\ell}|^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.$
To bound the other term, we first introduce for any $B\in\amsmathbb{R}^{(d+1)\times(d+1)}$ the notation $\|B\|=\sup_{\|x\|=1}\|Bx\|$. Then we have
$\|\mathcal{Q}_{h}-\mathcal{D}\|$ $=\big{\|}(\sigma^{\ell})^{-1}\big{(}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{)}+\big{(}(\sigma^{\ell})^{-1}-1\big{)}\mathcal{D}\|$
$\leq\big{\|}(\sigma^{\ell})^{-1}\big{\|}_{L^{\infty}(\mathcal{M})}\leavevmode\nobreak\ \big{\|}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{\|}$
$\quad\quad+\big{\|}(\sigma^{\ell})^{-1}-1\big{\|}_{L^{\infty}(\mathcal{M})}\leavevmode\nobreak\ \big{\|}\mathcal{D}\big{\|}.$
By Equation 47 and since $\mathcal{D}$ has bounded eigenvalues over $\mathcal{M}$, we obtain
(49) $\|\mathcal{Q}_{h}-\mathcal{D}\|$ $\lesssim\big{\|}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{\|}+h^{2},$
where the constant in the inequality is independent of the location on $\mathcal{M}$. We split the first term on the right into
$\big{\|}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}$ $\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{\|}$
$=\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}-\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}d_{s}^{\ell}\mathcal{H}\Pi-d_{s}^{\ell}\mathcal{H}\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi\big{\|}$
$\leq\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}+\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}d_{s}^{\ell}\mathcal{H}\Pi\big{\|}+\big{\|}d_{s}^{\ell}\mathcal{H}\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi\big{\|}.$
Since $\|d_{s}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}$ by Dziuk and Elliott (2013, Lemma 4.1) and $\mathcal{H}$ is defined independently of $h$, we conclude that
(50) $\big{\|}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{\|}$ $\lesssim\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}+h^{2}.$
We notice that $\mathcal{D}=\Pi\mathcal{D}\Pi$, since by definition of $\mathcal{D}$, $\mathcal{D}\nu=0$, which implies
$\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}$ $=\big{\|}\Pi\Pi_{h}^{\ell}\Pi\mathcal{D}\Pi\Pi_{h}^{\ell}\Pi-\Pi\mathcal{D}\Pi\big{\|}=\big{\|}(\Pi\Pi_{h}^{\ell}\Pi-\Pi)\mathcal{D}\Pi\Pi_{h}^{\ell}\Pi+\Pi\mathcal{D}(\Pi\Pi_{h}^{\ell}\Pi-\Pi)\big{\|}$
$\leq\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\big{\|}\mathcal{D}\Pi\Pi_{h}^{\ell}\Pi\big{\|}+\big{\|}\Pi\mathcal{D}\big{\|}\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}.$
Using that $\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\lesssim h^{2}$ by the proof of Dziuk and Elliott (2013, Lemma 4.1), we deduce that $\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}\lesssim h^{2}$. Injecting this inequality into Equation 50, and the resulting inequality into Equation 49, we conclude that
$\|\mathcal{Q}_{h}-\mathcal{D}\|\lesssim h^{2}.$
This allows us to write
$\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})$ $\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A\big{|}\leq\int_{\mathcal{M}}\|(\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A$
$\lesssim\int_{\mathcal{M}}h^{2}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A\leq h^{2}\int_{\mathcal{M}}(\mu_{\min})^{-1}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A,$
where $\mu_{\min}:\mathcal{M}\rightarrow\amsmathbb{R}_{+}$ maps any $x\in\mathcal{M}$ to the smallest eigenvalue of $\mathcal{D}(x)$ associated with an eigenvector in $\nu^{\perp}$. This last inequality is a consequence of the fact that by construction $\nabla_{\mathcal{M}}u^{\ell}_{h},\nabla_{\mathcal{M}}v^{\ell}_{h}\in\nu^{\perp}$ and using the characterization of eigenvalues through Rayleigh quotients. Since the non-zero eigenvalues of $\mathcal{D}$ are uniformly bounded above and below by positive constants, we conclude that
$\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\lesssim h^{2}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A.$
Then, using the Cauchy–Schwartz inequality yields
(51) $\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.$
Inserting the derived bounds Equation 48 and Equation 51 into Equation 46, we derive
$\big{|}\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})\big{|}\lesssim\leavevmode\nobreak$ $h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}$
$+h^{2}\bigg{(}\int_{\mathcal{M}}V|u_{h}^{\ell}|^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}V\leavevmode\nobreak\ |{v_{h}}^{\ell}|^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.$
Note that for any $u\in H^{1}(\mathcal{M})$,
$\mathsf{A}_{\mathcal{M}}(u,u)\geq\int_{\mathcal{M}}V|u|^{2}\,\mathrm{d}A,\quad\text{and}\quad\mathsf{A}_{\mathcal{M}}(u,u)\geq\int_{\mathcal{M}}\mathcal{D}\nabla_{\mathcal{M}}u\cdot\nabla_{\mathcal{M}}\overline{u}\,\mathrm{d}A=\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u\|_{\mathcal{D}}^{2}\,\mathrm{d}A,$
so we obtain
$\left|\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})\right|\lesssim h^{2}\sqrt{\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},u_{h}^{\ell})}\sqrt{\mathsf{A}_{\mathcal{M}}(v_{h}^{\ell},v_{h}^{\ell})}.$
Finally, due to Equation 5,
$\sqrt{\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},u_{h}^{\ell})}\sqrt{\mathsf{A}_{\mathcal{M}}(v_{h}^{\ell},v_{h}^{\ell})}\lesssim\|u_{h}^{\ell}\|_{H^{1}(\mathcal{M})}\|v_{h}^{\ell}\|_{H^{1}(\mathcal{M})},$
and the result follows. ∎
B.2. Norm and error estimates of discrete operators
We now prove some estimates for the norm of shifted inverses of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$, and for the error between inverses of these two operators. These results can be seen as extensions of the ones stated in Bonito et al. (2024, Lemma A.1).
Lemma B.2.
Let $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$ be arbitrary. Then, for all $z\in\Gamma$, for any $q\in[-1,1]$, and any $r\in(0,2)$,
(52) $\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},$
(53) $\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},$
(54) $\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{r/2-1}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})},$
(55) $\left\|(\mathcal{L}_{h}^{-1}v_{h})^{-\ell}-\mathsf{L}_{h}^{-1}\mathsf{P}_{h}(\sigma v_{h}^{-\ell})\right\|_{H^{1}(\mathcal{M}_{h})}\lesssim h^{2}\|\mathcal{L}_{h}^{-1/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
Proof.
We start with the proof of Equation 52. To this end, let $z\in\Gamma$ and $q\in[-1,1]$. For any $s\in[0,1]$ and $v_{h}\in S_{h}^{\ell}$, we expand
$\mathcal{L}_{h}^{s}v_{h}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{s}(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}e_{i}^{h},$
implying that
$\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}=\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}=\sum_{i=1}^{N_{h}}\left(|\lambda_{i}^{h}||(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|\right)^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2(1-s)}.$
Hölder’s inequality yields
(56) $\begin{split}\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}&\leq\left(\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{s}\left(\sum_{i=1}^{N_{h}}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{1-s}\\ &=\|\mathcal{L}_{h}v_{h}\|_{L^{2}(\mathcal{M})}^{2s}\|v_{h}\|_{L^{2}(\mathcal{M})}^{2(1-s)}.\end{split}$
Let $w_{h}=(z-\mathcal{L}_{h})^{-1}v_{h}$, then we obtain by Equation 20 that
$\|w_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-1}\|v_{h}\|_{L^{2}(\mathcal{M})}.$
Hence, adding and subtracting $z(z-\mathcal{L}_{h})^{-1}$ yields
$\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}$ $=\|z(z-\mathcal{L}_{h})^{-1}v_{h}-v_{h}\|_{L^{2}(\mathcal{M})}\leq|z|\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}+\|v_{h}\|_{L^{2}(\mathcal{M})}$
$\lesssim 2\|v_{h}\|_{L^{2}(\mathcal{M})}.$
Combining this estimate with Equation 56, we get for any $s\in[0,1]$
(57) $\|\mathcal{L}_{h}^{s}w_{h}\|_{L^{2}(\mathcal{M})}\leq\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}^{s}\|w_{h}\|_{L^{2}(\mathcal{M})}^{(1-s)}\leq|z|^{-(1-s)}\|v_{h}\|^{s}_{L^{2}(\mathcal{M})}\|v_{h}\|_{L^{2}(\mathcal{M})}^{1-s}=|z|^{s-1}\|v_{h}\|_{L^{2}(\mathcal{M})},$
yielding for any $q\in[-1,1]$ that
(58) $\begin{split}\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}&=\|\mathcal{L}_{h}^{1-(1+q)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})}\\ &\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},\end{split}$
where we applied Equation 57 with $s=1-(1+q)/2$ in the last step. Hence, we retrieve Equation 52.
Next, to prove Equation 53, we first observe that the estimates in Equation 57 and Equation 58 carry over to the case when $\mathsf{L}_{h}$ is used instead of $\mathcal{L}_{h}$, as the proof in essence is a standard manipulation of a finite eigenexpansion. In particular, we obtain (by taking $q=1$ and $s=0$)
(59) $\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}.$
Now, Equation 59 combined with the equivalence of norms in Equation 19, yields that
$\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},$
which shows Equation 53.
To bound Equation 54, we apply Equation 57 with $s=r/2$ to obtain
$\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}=\|\mathcal{L}_{h}^{r/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{r/2-1}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
Finally, to prove the bound in Equation 55, we rely on the geometric consistency estimate of Lemma B.1. Let $u_{h}=\mathcal{L}_{h}^{-1}v_{h}$ and let $U_{h}=\mathsf{L}_{h}^{-1}\mathsf{P}_{h}(\sigma v_{h}^{-\ell})$. Note that by definition of $\mathcal{L}_{h}^{-1}$,
$(\mathcal{L}_{h}u_{h},w_{h})_{L^{2}(\mathcal{M})}=\mathsf{A}_{\mathcal{M}}(u_{h},w_{h})=(v_{h},w_{h})_{L^{2}(\mathcal{M})},$
for all $w_{h}\in S_{h}^{\ell}$. Likewise, for $\mathsf{L}_{h}^{-1}$ we obtain
$(\mathsf{L}_{h}U_{h},W_{h})_{L^{2}(\mathcal{M}_{h})}=\mathsf{A}_{\mathcal{M}_{h}}(U_{h},W_{h})=(\mathsf{P}_{h}(\sigma v_{h}^{-\ell}),W_{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma v_{h}^{-\ell},W_{h})_{L^{2}(\mathcal{M}_{h})}=(v_{h},W_{h}^{\ell})_{L^{2}(\mathcal{M})},$
for all $W_{h}\in S_{h}$, where we used the definition of $\sigma$ in the last step. Let us now select a fixed, but arbitrary, $\Xi_{h}\in S_{h}$. Then, by combining the last two equations,
$|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|$ $=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}_{h}}(U_{h},\Xi_{h})|=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-(v_{h},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}|$
$=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}}(u_{h},\Xi_{h}^{\ell})|,$
meaning that an application of Lemma B.1 results in the bound
(60) $|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|\lesssim h^{2}\|u_{h}\|_{H^{1}(\mathcal{M})}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}.$
Further, note that for any $\Xi_{h}^{\ell}\in S_{h}^{\ell}$, the equivalence of norms (19) gives
(61) $\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}\sim\|\mathcal{L}_{h}^{1/2}\Xi_{h}^{\ell}\|_{L^{2}(\mathcal{M})}^{2}=({\mathcal{L}_{h}}\Xi_{h}^{\ell},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}=\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell}),$
where the last equality comes from the definition of $\mathcal{L}_{h}$. And similarly, for any $\Xi_{h}\in S_{h}$, we have
(62) $\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}\sim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h}).$
Then, applying the triangle inequality to the (last) right-hand side of Equation 61 gives
$\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})+\big{|}\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})\big{|}\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2},$
where we used Equation 62 and Lemma B.1 to derive the second inequality. This means in particular that there exists $C>0$ independent of $h$ such that $\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}\leq C(\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2})$. Recall that $h\in(0,h_{0})$ for some $h_{0}\in(0,1)$ small enough. Assuming that especially $01-Ch_{0}^{2}>$ 0 yields $\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}\leq C(1-Ch^{2})^{-1}\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}\leq C(1-Ch_{0}^{2})^{-1}\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}$, which allows us to conclude that
(63) $\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}.$
Now, applying successively Equation 62 and Equation 60 with $\Xi_{h}=u_{h}^{-\ell}-U_{h}$, we obtain
$\|u_{h}^{-\ell}-U_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}\lesssim h^{2}\|u_{h}\|_{H^{1}(\mathcal{M})}\|u_{h}-U_{h}^{\ell}\|_{H^{1}(\mathcal{M})}\lesssim h^{2}\|u_{h}\|_{H^{1}(\mathcal{M})}\|u_{h}^{-\ell}-U_{h}\|_{H^{1}(\mathcal{M}_{h})},$
where the last inequality is derived from applying Equation 63. Therefore, we end up with
$\|u_{h}-U_{h}^{\ell}\|_{H^{1}(\mathcal{M})}\lesssim h^{2}\|u_{h}\|_{H^{1}(\mathcal{M})}\lesssim h^{2}\|\mathcal{L}_{h}^{-1/2}v_{h}\|_{L^{2}(\mathcal{M})},$
where the equivalence of norms (19) together with the definition of $u_{h}$ are used in the final step. This concludes the proof of Equation 55. ∎
B.3. Proof of Equation 24
Based on the results in the previous subsections, we can now move on to the proof of Equation 24.
Proof.
Let $\tilde{f}\in S_{h}^{\ell}$. We introduce the inverse lift operator $\mathcal{P}_{\ell}:L^{2}(\mathcal{M})\rightarrow L^{2}(\mathcal{M}_{h})$ which maps any $F\in L^{2}(\mathcal{M})$ to $\mathcal{P}_{\ell}F=F^{-\ell}$. Let then $\mathcal{E}_{h}=\big{\|}\big{(}\gamma(\mathcal{L}_{h})\tilde{f}\big{)}^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\big{\|}_{L^{2}(\mathcal{M}_{h})}=\big{\|}\mathcal{P}_{\ell}\gamma(\mathcal{L}_{h})\tilde{f}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\mathcal{P}_{\ell}\tilde{f})\big{\|}_{L^{2}(\mathcal{M}_{h})}$. Note that by the integral representations of the operators (21)
$\mathcal{E}_{h}=\big{\|}\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}(z)\,\mathrm{d}z\big{\|}_{L^{2}(\mathcal{M}_{h})},$
where we take for any $z\in\Gamma$, $\mathcal{F}(z)=\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}-(z-\mathsf{L}_{h})^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}$. Similarly, as in the proof of Proposition 2.3, we use the splitting (22) of $\Gamma$ and the triangle inequality to deduce that
(64) $\mathcal{E}_{h}$ $\leq\frac{1}{2\pi}\int_{\Omega_{+}}|\gamma(g_{+}(t))|\left\|\mathcal{F}(g_{+}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{\Omega_{0}}|\gamma(g_{0}(t))|\|\mathcal{F}(g_{0}(t)\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t$
$\qquad+\frac{1}{2\pi}\int_{\Omega_{-}}|\gamma(g_{-}(t))|\left\|\mathcal{F}(g_{-}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
[…]
$\mathcal{E}_{h}^{*}=\int_{\Omega_{*}}|\gamma(g_{*}(t))|\left\|\mathcal{F}(g_{*}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
so that Equation 64 may be rewritten as $\mathcal{E}_{h}\lesssim\mathcal{E}_{h}^{+}+\mathcal{E}_{h}^{0}+\mathcal{E}_{h}^{-}$ and in particular, $g_{*}(t)\in\Gamma$ for any $t\in\Omega_{*}$.
[…]
$\mathcal{F}(z)$ $=(z-\mathsf{L}_{h})^{-1}\mathsf{L}_{h}\left((z\mathsf{L}_{h}^{-1}-I)\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}-\mathsf{L}_{h}^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}(z\mathcal{L}_{h}^{-1}-I)\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}$
$=(z-\mathsf{L}_{h})^{-1}\mathsf{L}_{h}\left(z\mathsf{L}_{h}^{-1}(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}+\mathsf{L}_{h}^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}-\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}$
$=\mathcal{F}_{1}(z)+\mathcal{F}_{2}(z),$
where we take $\mathcal{F}_{1}(z)=(z-\mathsf{L}_{h})^{-1}\mathsf{L}_{h}\left(z\mathsf{L}_{h}^{-1}(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}=z(z-\mathsf{L}_{h})^{-1}(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}$ and $\mathcal{F}_{2}(z)=(z\leavevmode\nobreak\ -\leavevmode\nobreak\ \mathsf{L}_{h})^{-1}\mathsf{L}_{h}\left(\mathsf{L}_{h}^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}-\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}$. Hence, by the triangle inequality,
(65) $\mathcal{E}_{h}^{*}$ $\lesssim\int_{\Gamma_{t}}|\gamma(g_{*}(t))|\big{(}\left\|\mathcal{F}_{1}(g_{*}(t))\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}+\big{\|}\mathcal{F}_{2}(g_{*}(t))\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}\right)\,\mathrm{d}t.$
We first bound $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$. Using successively Equation 23 and the geometric estimates in Bonito et al. (2018, Corollary 2.2) results in
$\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$ $=|z|\left\|(z-\mathsf{L}_{h})^{-1}(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$
$\lesssim\left\|(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}\left\|(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M})}.$
Using then Equation 54, we conclude that, for any $p\in(0,2)$,
(66) $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}|z|^{-(1-p/2)}\left\|\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}^{2}.$
To bound $\big{\|}\mathcal{F}_{2}(z)\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}$, we apply Equations 53 and 55 to obtain
$\left\|\mathcal{F}_{2}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$ $\lesssim|z|^{-1/2}\left\|\left(\mathsf{L}_{h}^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}-\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{H^{1}(\mathcal{M}_{h})}$
$\lesssim|z|^{-1/2}h^{2}\left\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M})}$
and with Equation 52 for any $p\in[-1,1]$ ($p=-q$)
(67) $\left\|\mathcal{F}_{2}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}|z|^{-(1-p/2)}\left\|\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}.$
Using Equations 67 and 66 with $p=\min\{\alpha+d/4;1\}$ together with Equation 65 gives
$\mathcal{E}_{h}^{*}$ $\lesssim\int_{\Omega_{*}}|\gamma(g_{*}(t))|h^{2}|g_{*}(t)|^{-(1-p/2)}\|\mathcal{L}_{h}^{-p/2}\tilde{f}\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,$
which yields in turn (since $\gamma$ is an $\alpha$-power spectral density)
$\mathcal{E}_{h}^{*}\lesssim h^{2}\|\mathcal{L}_{h}^{-p/2}\tilde{f}\|_{L^{2}(\mathcal{M})}\int_{\Omega_{*}}|g_{*}(t)|^{-(1+\alpha-p/2)}\,\mathrm{d}t\lesssim h^{2}\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})},$
[…]

### v2

@@
 A flexible model for non-stationary Gaussian random fields on hypersurfaces is introduced.
-The class of random fields on curves and surfaces is characterized by a power spectral density of a second order elliptic differential operator.
+The class of random fields on curves and surfaces is characterized by an amplitude spectral density of a second order elliptic differential operator.
 Sampling is done by a Galerkin–Chebyshev approximation based on the surface finite element method and Chebyshev polynomials.
 Strong error bounds are shown with convergence rates depending on the smoothness of the approximated random field.
+Numerical experiments that confirm the convergence rates are presented.
 Key words and phrases:
@@
 60G60, 60H35, 60H15,65C30, 60G15, 58J05, 41A10, 65N30,65M60
-Acknowledgment: The authors thank Julie Rowlett and Iosif Polterovich for helpful discussions.
+Acknowledgment: The authors thank Mihály Kovács, Julie Rowlett and Iosif Polterovich for helpful discussions.
 This work was supported in part by the European Union (ERC, StochMan, 101088589), by the Mines Paris - PSL / INRAE “Geolearning” chair, by the Swedish Research Council (VR) through grant no.
-2020-04170, by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Chalmers AI Research Centre (CHAIR)
+2020-04170, by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Chalmers AI Research Centre (CHAIR).
+Funded by the European Union.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Neither the European Union nor the granting authority can be held responsible for them.
 1. Introduction
 Random fields are powerful tools for modeling spatially dependent data.
-They have found uses in a wide range of applications, for instance in geostatistics, cosmological data analysis, climate modeling, and biomedical imaging (Marinucci and Peccati, 2011; Farag, 2014).
+They have found uses in a wide range of applications, for instance in geostatistics, cosmological data analysis, climate modeling, and biomedical imaging [22, 12].
 One challenge in the modeling of spatial data is non-stationary behavior, i.e., different behaviors in different parts of the domain.
@@
 In this paper, we present a surface finite element-based method to sample a flexible class of non-stationary random fields on curves and surfaces and show its strong convergence.
-The method, building on the foundational work for stationary fields introduced in Lang and Pereira (2023), is an extension of the stochastic partial differential equation (SPDE) approach pioneered by Whittle (1963) and popularized by Lindgren et al.
-(2011).
+The method, building on the foundational work for stationary fields introduced in [19], is an extension of the stochastic partial differential equation (SPDE) approach pioneered by [27] and popularized by [21].
 The idea behind our method is to color white noise by applying a function of an elliptic differential operator $\mathcal{L}$.
@@
 (1) $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W},$
-where $\gamma$ is a function, called the power spectral density, $\mathcal{L}$ is an elliptic differential operator and $\mathcal{W}$ denotes white noise.
+where $\mathcal{L}$ is an elliptic differential operator, $\mathcal{W}$ denotes white noise, $\gamma$ is a function, called amplitude spectral density in analogy to the spectral analysis of time signals [15].
 By letting the coefficients of the differential operator vary over the domain, we can obtain local, non-stationary behaviors.
@@
 With the suggested model, we can achieve preferred directions, local activation, and local deactivation.
-The computational method we use to solve Equation 1, i.e., sample the random fields, is based on the surface finite element method (SFEM), a computational method pioneered by Dziuk (1988); Dziuk and Elliott (2013) and that has been used in the context of the generation of random fields in for instance Bonito et al.
-(2024); Jansson et al.
-(2022); Lang and Pereira (2023).
-Our main mathematical contribution is a strong convergence result.
-Using a functional calculus approach to the finite element discretization error, we obtain a strong rate of convergence of order $\mathcal{O}(K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}})$, where $d=2$ for surfaces and $d=1$ for curves, and $K_{\alpha}(h)$ is a dimension-dependent logarithmic factor.
-The SPDE approach to random fields and their approximation have been studied previously for both surfaces and Euclidean geometries, examples include Bonito et al.
-(2024); Borovitskiy et al.
-(2020); Bolin et al.
-(2020); Cox and Kirchner (2020); Jansson et al.
-(2022); Lang and Pereira (2023); Lindgren et al.
-(2011, 2022).
-However, to the best of our knowledge, we are the first to present a strong error analysis for the non-stationary case on hypersurfaces.
-Moreover, we achieve this without requiring an explicit approximation of the eigenfunctions of the elliptic operator, and contrary to for instance Borovitskiy et al.
-(2020), computing the eigenfunctions.
-In fact, the computation of eigenfunctions, with theoretical guarantees, is a notoriously difficult problem, see e.g., Boffi (2010).
+The computational method we use to solve Equation 1, i.e., sample the random fields, is based on the surface finite element method (SFEM), a computational method pioneered by [10, 11] and that has been used in the context of the generation of Whittle–Matérn random fields with Laplace–Beltrami operators in for instance [6, 18, 19].
+Our main mathematical contribution is a strong convergence result for equations with more general elliptic operators and amplitude spectral densities than fractional powers.
+Using a functional calculus approach to the finite element discretization error, we obtain a strong rate of convergence of order $\mathcal{O}(C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}})$, where $d=2$ for surfaces and $d=1$ for curves, and $C_{\alpha}(h)$ is a dimension-dependent logarithmic factor.
+The SPDE approach to random fields and their approximation have been studied previously for both surfaces and Euclidean geometries, examples include [6, 8, 3, 9, 18, 19, 21, 20].
+However, to the best of our knowledge, we are the first to present a strong error analysis for the non-stationary case on hypersurfaces that are only given by mesh points.
+In contrast to the method presented in [16], we do not need to explicitly know the true hypersurface and perform computations on it.
+Instead, we only require a polyhedral approximation of the hypersurface.
+Moreover, like in [9], we achieve this without requiring explicit approximation bounds on the eigenfunctions of the elliptic operator, and contrary to for instance [8], computing the eigenfunctions.
+In fact, the computation of eigenfunctions, with theoretical guarantees, is a notoriously difficult problem, see e.g., [2].
 Our method circumvents this issue by introducing a Chebyshev approximation.
@@
 We introduce first the class of considered random fields in Section 3.1 and their Galerkin–Chebyshev approximation in Section 3.2.
-The proof of its strong convergence is split into the SFEM error in Section 3.3 and the Chebyshev approximation error Section 3.4.
-For numerical confirmation of our theoretical results, we refer to our earlier work Lang and Pereira (2023) and the results presented in Bonito et al.
-(2024).
+The proof of its strong convergence is split into the SFEM error in Section 3.3 and the Chebyshev approximation error in Section 3.4.
+In Section 4, we present numerical experiments that confirm the strong error bounds.
 The source code used to generate the figures is available at this address: https://github.com/mike-pereira/SFEMsim.
@@
 We set $H^{0}(\mathcal{M})=L^{2}(\mathcal{M})$.
-The reader is referred to Herrmann et al.
-(2018), Strichartz (1983), Triebel (1985), and references therein for more details on Sobolev spaces defined using Bessel potentials.
+The reader is referred to [17], [25], [26], and references therein for more details on Sobolev spaces defined using Bessel potentials.
 In this work, we consider elliptic differential operators associated to bilinear forms $\mathsf{A}_{\mathcal{M}}$ given by
@@
 In particular, since $T_{x_{0}}\mathcal{M}=[\nu(x_{0})^{\perp}]$, $\mathcal{D}(x_{0})$ is simply a matrix admitting $\nu(x_{0})$ as an eigenvector with some eigenvalue $\mu_{1}(x_{0})\in\amsmathbb{R}$, and such that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 associated with its other eigenvectors are positive.
-Without loss of generality, we may assume that $\mu_{1}(x_{0})=0$, meaning that $\mathcal{D}\nu=0$, that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 are uniformly lower-bounded and upper-bounded on $\mathcal{M}$ by positive constants, and that $D_{ij}\in L^{\infty}(\mathcal{M})$ for any $11\leq i,j\leq d+$ 1.
+Without loss of generality, we may assume that $\mu_{1}(x_{0})=0$, meaning that $\mathcal{D}\nu=0$, that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 are uniformly lower-bounded and upper-bounded on $\mathcal{M}$ by positive constants, and for simplicity that $D_{ij}\in C^{\infty}(\mathcal{M})$ for any $11\leq i,j\leq d+$ 1.
 Finally, we assume that $V\in L^{\infty}(\mathcal{M})$ is a real-valued function that satisfies $V_{-}\leq V\leq V_{+}$ for some $0<V_{-}\leq V_{+}<+\infty$.
@@
 (5) $\left|\mathsf{A}_{\mathcal{M}}(u,v)\right|\leq M\|u\|_{H^{1}(\mathcal{M})}\,\|v\|_{H^{1}(\mathcal{M})}.$
-Following Yagi (2010, Equation (1.33)), $\mathsf{A}_{\mathcal{M}}$ gives rise to an associated elliptic differential operator $\mathcal{L}:H^{1}(\mathcal{M})\rightarrow H^{-1}(\mathcal{M})$ defined weakly by
+Following [28, Equation (1.33)], $\mathsf{A}_{\mathcal{M}}$ gives rise to an associated elliptic differential operator $\mathcal{L}:H^{1}(\mathcal{M})\rightarrow H^{-1}(\mathcal{M})$ defined weakly by
 $\mathsf{A}_{\mathcal{M}}(u,v)=\int_{\mathcal{M}}(\mathcal{L}u)\overline{v}\,\mathrm{d}A,\quad u,v\in H^{1}(\mathcal{M}).$
@@
 Therefore, the eigenvalue problem for $\mathcal{L}$ is equivalent to that for the Laplace–Beltrami operator on $\mathcal{M}$ equipped with a possibly rough metric if the coefficients of $\mathcal{D}$ are not smooth.
-The results in Bandara et al.
-(2021) imply growth rates on the eigenvalues in accordance with Weyl’s law, and more specifically that there exist $c_{\mathcal{M}},C_{\mathcal{M}}>0$ such that for any $i\in\amsmathbb{N}$
+The results in [1] imply growth rates on the eigenvalues in accordance with Weyl’s law, and more specifically that there exist $c_{\mathcal{M}},C_{\mathcal{M}}>0$ such that for any $i\in\amsmathbb{N}$
 (6) $c_{\mathcal{M}}i^{2/d}\leq\lambda_{i}\leq C_{\mathcal{M}}i^{2/d}.$
 As a last step in this subsection, we introduce nonlinear functions of $\mathcal{L}$ which allow later in Section 3 for the definition of a variety of Gaussian random fields.
-For that, we call a function $\gamma:\amsmathbb{R}_{+}\rightarrow\amsmathbb{R}$ an $\alpha$-power spectral density if
+For that, we call a function $\gamma:\amsmathbb{R}_{+}\rightarrow\amsmathbb{R}$ an $\alpha$-amplitude spectral density if
   1. 1)
@@
     (7) $\left|\gamma(z)\right|\leq C_{\gamma}|z|^{-\alpha}.$
-Applying a power spectral density to $\mathcal{L}$ results in a linear operator $\gamma(\mathcal{L})$ whose action on functions $f\in L^{2}(\mathcal{M})$ is defined by
+Applying a amplitude spectral density to $\mathcal{L}$ results in a linear operator $\gamma(\mathcal{L})$ whose action on functions $f\in L^{2}(\mathcal{M})$ is defined by
 (8) $\gamma(\mathcal{L})f=\sum_{i=1}^{\infty}\gamma(\lambda_{i})(f,e_{i})_{L^{2}(\mathcal{M})}e_{i},$
 where $\{(\lambda_{i},e_{i})\}_{i\in\amsmathbb{N}}$ are the eigenpairs of $\mathcal{L}$ defined in Proposition 2.1.
-A typical example is the function $\gamma(\lambda)=(\kappa^{2}+\lambda)^{-\alpha}$ for $\alpha>d/4$ and $\kappa>0$, which can be used to obtain Whittle–Matérn random fields (Lang and Pereira, 2023).
+A typical example is the function $\gamma(\lambda)=(\kappa^{2}+\lambda)^{-\alpha}$ for $\alpha>d/4$ and $\kappa>0$, which can be used to obtain Whittle–Matérn random fields [19].
+Remark 2.2.
+For any $\alpha$-amplitude spectral density $\gamma$ and any $f\in L^{2}(\mathcal{M})$, $\gamma(\mathcal{L})f\in L^{2}(\mathcal{M})$.
+In fact, we have for any $s\in[0,\alpha]$, $\|\mathcal{L}^{s}(\gamma(\mathcal{L})f)\|_{L^{2}(\mathcal{M})}<\infty$.
+Indeed,
+$\|\mathcal{L}^{s}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=\sum_{i=1}^{\infty}|\lambda_{i}^{s}\gamma(\lambda_{i})|^{2}|(f,e_{i})_{L^{2}(\mathcal{M})}|^{2},$
+where, since $\gamma$ is an $\alpha$-amplitude spectral density and $\lambda_{i}\in H_{\pi/2}$ for any $i\in\amsmathbb{N}$, $|\lambda_{i}^{s}\gamma(\lambda_{i})|\lesssim|\lambda_{i}|^{-(\alpha-s)}$.
+Using then the fact that for any $i\in\amsmathbb{N}$, $\lambda_{i}\geq\lambda_{1}>0$, and that $\alpha-s\geq 0$, we conclude that $|\lambda_{i}^{s}\gamma(\lambda_{i})|\lesssim|\lambda_{1}|^{-(\alpha-s)}\lesssim 1$ and therefore $\|\mathcal{L}^{s}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}\lesssim\sum_{i=1}^{\infty}|(f,e_{i})_{L^{2}(\mathcal{M})}|^{2}=\|f\|_{L^{2}(\mathcal{M})}<\infty$.
 The goal of the remainder of this section is to study the approximation of functions of the form $u=\gamma(\mathcal{L})f$, where $f\in L^{2}(\mathcal{M})$.
-Formally, if $1/\gamma$ is well-defined on the spectrum of $\mathcal{L}$, then this is the solution to the partial differential equation $\gamma(\mathcal{L})^{-1}u=f$.
+Formally, if $1/\gamma$ is well-defined on the spectrum of $\mathcal{L}$, then this is the solution to the partial differential equation $(1/\gamma)(\mathcal{L})u=f$.
 2.2. SFEM–Galerkin approximation
-The idea behind the surface finite element method, as introduced by Dziuk (1988), is to work on a polyhedral approximation of the surface that is in some sense close to the true surface $\mathcal{M}$.
+The idea behind the surface finite element method, as introduced by [10], is to work on a polyhedral approximation of the surface that is in some sense close to the true surface $\mathcal{M}$.
 More precisely, fix $h>0$ and let $\mathcal{M}_{h}$ be a piecewise polygonal surface consisting of non-degenerate simplices (for $d=2$, triangles and for $d=1$, line segments) with vertices on $\mathcal{M}$, and such that $h$ is the size of the largest simplex defined as the in-ball radius.
@@
 and we assume that for any two simplices in $\mathcal{T}_{h}$, it holds that their intersection is either empty, or a common edge or vertex.
-Following (Dziuk and Elliott, 2013, Section 1.4.1), we assume that the triangulation $\mathcal{T}_{h}$ is quasi-uniform, shape-regular, and that the number of simplices sharing the same vertex can be upper-bounded by a constant independent of $h$.
+Following [11, Section 1.4.1], we assume that the triangulation $\mathcal{T}_{h}$ is quasi-uniform, shape-regular, and that the number of simplices sharing the same vertex can be upper-bounded by a constant independent of $h$.
 In turn, these two assumption imply that $N_{h}\propto h^{-d}$ where $N_{h}\in\amsmathbb{N}$ denotes the number of vertices of $\mathcal{M}_{h}$.
@@
 First, note that $\mathcal{M}$ can be seen as the boundary of some bounded open set $G\subset\amsmathbb{R}^{d+1}$ with exterior normal $\nu$.
-Then, following Dziuk and Elliott (2013, Section 2.3), we consider that there exists some (small) $\varpi>0$ such that $\mathcal{M}_{h}$ is contained in a so-called tubular neighborhood $U_{\varpi}$ of $\mathcal{M}$ defined by
+Then, following [11, Section 2.3], we consider that there exists some (small) $\varpi>0$ such that $\mathcal{M}_{h}$ is contained in a so-called tubular neighborhood $U_{\varpi}$ of $\mathcal{M}$ defined by
 $U_{\varpi}=\{x\in\amsmathbb{R}^{d+1}:|d_{s}(x)|<\varpi\},$
@@
 $(u_{h},v_{h})_{L^{2}(\mathcal{M}_{h})}=\int_{\mathcal{M}_{h}}u_{h}\overline{v}_{h}\,\mathrm{d}A_{h},\quad u_{h},v_{h}\in L^{2}(\mathcal{M}_{h}).$
-Following (Bonito et al., 2020, Section 1.2.1), we denote by $\sigma:\mathcal{M}_{h}\rightarrow\amsmathbb{R}^{+}$ the area element given by $\sigma=\mathrm{d}A/\mathrm{d}A_{h}$, such that for all $v\in L^{2}(\mathcal{M})$,
+Following [4, Section 1.2.1], we denote by $\sigma:\mathcal{M}_{h}\rightarrow\amsmathbb{R}^{+}$ the area element given by $\sigma=\mathrm{d}A/\mathrm{d}A_{h}$, such that for all $v\in L^{2}(\mathcal{M})$,
 (9) $\int_{\mathcal{M}}v\,\mathrm{d}A=\int_{\mathcal{M}_{h}}\sigma v^{-\ell}\,\mathrm{d}A_{h},$
@@
 Note that the points on the discretized surface are lifted along the normal $\nu$ to the surface $\mathcal{M}$.
-The mapping $a$ is used to define the gradient of functions on $\mathcal{M}_{h}$ (Dziuk and Elliott, 2013).
-More specifically, for a differentiable $\eta:\mathcal{M}_{h}\to\amsmathbb{C}$, the gradient is given by
+The mapping $a$ is used to define the gradient of functions on $\mathcal{M}_{h}$ [11]. More specifically, for a differentiable $\eta:\mathcal{M}_{h}\to\amsmathbb{C}$, the gradient is given by
 (10) $\nabla_{\mathcal{M}_{h}}\eta(x)=\nabla\check{\eta}(x)-(\nabla\check{\eta}(x)\cdot\nu_{h}(x))\nu_{h}(x)\in T_{x}\mathcal{M}_{h},$
@@
 The analogue of the bilinear form $\mathsf{A}_{\mathcal{M}}$ on $\mathcal{M}_{h}$ is given by
-(11) $\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}+\int_{\mathcal{M}_{h}}(V^{-\ell}u_{h})\overline{v}_{h}\,\mathrm{d}A_{h},\quad u_{h},v_{h}\in H^{1}(\mathcal{M}_{h}),$
-where $\mathcal{D}^{-\ell}=[D_{ij}^{-\ell}]_{i,j=1}^{d+1}$.
-As in Dziuk and Elliott (2013), we assume that there exists $h_{0}\leavevmode\nobreak\ \in\leavevmode\nobreak\ (0,1)$ small enough, such that $\mathsf{A}_{\mathcal{M}_{h}}$ is coercive (and continuous) whenever $h\leq h_{0}$.
+(11) $\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}+\int_{\mathcal{M}_{h}}(V^{-\ell}u_{h})\overline{v}_{h}\,\mathrm{d}A_{h},$
+where $u_{h},v_{h}\in H^{1}(\mathcal{M}_{h})$ and $\mathcal{D}^{-\ell}=[D_{ij}^{-\ell}]_{i,j=1}^{d+1}$.
+As in [11], we assume that there exists $h_{0}\leavevmode\nobreak\ \in\leavevmode\nobreak\ (0,1)$ small enough, such that $\mathsf{A}_{\mathcal{M}_{h}}$ is coercive (and continuous) whenever $h\leq h_{0}$.
 Unless stated otherwise, we now assume that this last condition on $h$ is fulfilled.
 To conclude this subsection, we introduce the (linear) finite element space $S_{h}$ on $\mathcal{M}_{h}$.
-The finite element space $S_{h}$ is defined as the complex span of the standard real-valued nodal basis $\psi_{1},\dots,\psi_{N_{h}}\colon\mathcal{M}\rightarrow\amsmathbb{R}$, where for any $i\in\{1,\dots,N_{h}\}$, $\psi_{i}|_{T}$ is a polynomial of at most degree one taking the value $1$ at the $i$-th vertex of $\mathcal{M}_{h}$ and 00 at all the other vertices, i.e.,
+The finite element space $S_{h}$ is defined as the complex span of the standard real-valued nodal basis
+$\psi_{1},\dots,\psi_{N_{h}}\colon\mathcal{M}\rightarrow\amsmathbb{R},$
+where for any $i\in\{1,\dots,N_{h}\}$, $\psi_{i}|_{T}$ is a polynomial of at most degree one taking the value $1$ at the $i$-th vertex of $\mathcal{M}_{h}$ and 00 at all the other vertices, i.e.,
 $S_{h}=\operatorname{span}\left(\psi_{1},\dots,\psi_{N_{h}}\right)\subset H^{1}(\mathcal{M}_{h}).$
@@
 $\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})=(\mathcal{L}_{h}u_{h}^{\ell},v_{h}^{\ell})_{L^{2}(\mathcal{M})}.$
-Since the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$ are coercive, positive definite, Hermitian and have real coefficients, these two operators are diagonalizable in the sense that they each give rise to a set of $N_{h}$ eigenpairs (Hall, 2013).
+Since the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$ are coercive, positive definite, Hermitian and have real coefficients, these two operators are diagonalizable in the sense that they each give rise to a set of $N_{h}$ eigenpairs [14].
 On the one hand, there exists a sequence $0\leq\Lambda_{1}^{h}\leq\dots\leq\Lambda_{N_{h}}^{h}$ and an $L^{2}(\mathcal{M}_{h})$-orthonormal basis $E_{1}^{h},\ldots,E_{N_{h}}^{h}$ of $S_{h}$ such that
@@
 In particular, using the same approach as in Proposition 2.1, we can assume that the eigenfunctions $\{E_{i}^{h}\}_{1\leq i\leq N_{h}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are all real-valued.
-The eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$ are linked to one another through the following lemma, due to Bonito et al.
-(2018, Lemma 3.1), Bonito et al.
-(2024, Lemma 4.1) and Strang and Fix (2008, Theorem 6.1).
+The eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$ are linked to one another through the following lemma, due to [5, Lemma 3.1], [6, Lemma 4.1] and [24, Theorem 6.1].
 In the following, $A\lesssim B$ is shorthand for that there is a constant $C>0$ such that $A\leq CB$.
@@
 Let then $\sqrt{\bm{C}}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ be an invertible matrix satisfying $\sqrt{\bm{C}}(\sqrt{\bm{C}})^{T}=\bm{C}$.
-Then, by Lang and Pereira (2023, Corollary 3.2), the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ are also the eigenvalues of the matrix $\bm{S}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ defined by
+Then, by [19, Corollary 3.2], the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ are also the eigenvalues of the matrix $\bm{S}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ defined by
 (15) $\bm{S}=\big{(}\sqrt{\bm{C}}\big{)}^{-1}\bm{R}\big{(}\sqrt{\bm{C}}\big{)}^{-T}.$
@@
 2.3. Deterministic error analysis
+From now on, let us make the following assumption on the mesh size $h$.
+Let $\delta_{0}\in(0,\delta/2)$ be fixed and arbitrary and let $h_{0}\in(0,1)$ such that $h_{0}^{-2}>\delta_{0}$ and $|\log h_{0}|>1$.
+We now assume for the remainder of the paper n that mesh size satisfies $h\in(0,h_{0})$.
 Based on the introduced framework, we are now in place to quantify the error between functions of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$.
@@
 Proposition 2.4.
-Let $\gamma$ be an $\alpha$-power spectral density. Then, there is a constant $C>0$ such that for all $f\in L^{2}(\mathcal{M})$
-$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $\|_{L^{2}(\mathcal{M})}\leq\begin{cases}Ch^{2}\|f\|_{L^{2}(\mathcal{M})},&\alpha>1,\\ Ch^{2}|\log h|\|f\|_{L^{2}(\mathcal{M})},&\alpha=1,\\ Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})},&\alpha<1.\end{cases}$
+Let $\gamma$ be an $\alpha$-amplitude spectral density.
+There exists some constant such that for any $h\in(0,h_{0})$, for all $f\in L^{2}(\mathcal{M})$ and for any $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$, we have
+$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $\|_{L^{2}(\mathcal{M})}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
+where $C_{\alpha+p}(h)=|\log h|$ if $\alpha+p\leq 1$, and $C_{\alpha+p}(h)=1$ otherwise.
 To prove this proposition we rely on a representation of functions of operators based on Cauchy–Stieltjes integrals, which are constructed as follows.
-Since the sesquilinear form defined by $\mathsf{A}_{\mathcal{M}}$ is continuous and coercive, the associated operators $\mathcal{L}$ and $\mathcal{L}_{h}$ are sectorial with some (common) angle $\theta\in(0,\pi/2)$ (Yagi, 2010, Theorem 2.1).
+Since the sesquilinear form defined by $\mathsf{A}_{\mathcal{M}}$ is continuous and coercive, the associated operators $\mathcal{L}$ and $\mathcal{L}_{h}$ are sectorial with some (common) angle $\theta\in(0,\pi/2)$ [28, Theorem 2.1].
 Therefore, the spectra of $\mathcal{L}$ and $\mathcal{L}_{h}$ are contained in the complement of the set $G_{\theta}=\{z\in\amsmathbb{C},\theta\leq\arg(z)\leq\pi\}$, as illustrated in Figure 3(a), and the following inequalities are satisfied for any $z\in G_{\theta}$ (cf.
-Yagi (2010, Equation (2.2))):
-$\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v\|_{L^{2}(\mathcal{M})},\quad v\in L^{2}(\mathcal{M}),$
+[28, Equation (2.2)]):
+(20) $\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v\|_{L^{2}(\mathcal{M})},\quad v\in L^{2}(\mathcal{M}),$
 (21) $\|(z-\mathcal{L}_{h})^{-1}v_{h}^{\ell}\|_{L^{2}(\mathcal{M})}$ $\leq C_{\theta}|z|^{-1}\|v_{h}^{\ell}\|_{L^{2}(\mathcal{M})},\quad v_{h}^{\ell}\in S_{h}^{\ell},$
 where $C_{\theta}>0$ is a generic constant.
-Note in particular that by definition, $G_{\theta}$ is contained in the resolvent sets of $\mathcal{L}$ and $\mathcal{L}_{h}$, and that any power spectral density $\gamma$ is bounded, holomorphic and satisfies the inequality $|z|^{\alpha}|\gamma(z)|\lesssim 1$ for any $z\in G_{\theta}$.
-Hence, the operators $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$ can be defined as functional calculi of the operators $\mathcal{L}$ and $\mathcal{L}_{h}$ as (Yagi, 2010, Chapter 16, Section 1.2) by
+Note in particular that by definition, $G_{\theta}$ is contained in the resolvent sets of $\mathcal{L}$ and $\mathcal{L}_{h}$, and that any amplitude spectral density $\gamma$ is bounded, holomorphic and satisfies the inequality $|z|^{\alpha}|\gamma(z)|\lesssim 1$ for any $z\in G_{\theta}$.
+Hence, the operators $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$ can be defined as functional calculi of the operators $\mathcal{L}$ and $\mathcal{L}_{h}$ as [28, Chapter 16, Section 1.2] by
 (22) $\gamma(\mathcal{L})=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)(z-\mathcal{L})^{-1}\,\mathrm{d}z\quad\text{and}\quad\gamma(\mathcal{L}_{h})=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)(z-\mathcal{L}_{h})^{-1}\,\mathrm{d}z,$
@@
 e.g.
-(Yagi, 2010, Remark 2.7)).
+[28, Remark 2.7]).
 In the remainder, we split the contour $\Gamma$ into
@@
 We use this contour to prove Proposition 2.4, while relying on the following bounds for the resolvent error along $\Gamma$.
-The proof is included in Appendix A and is an adaption of results from Fujita and Suzuki (1991) and Bonito et al.
-(2024).
+The proof is included in Section B.3 and is an adaption of results from [13, 6, 7].
 Lemma 2.5.
-For any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$, and $\beta\in[0,1]$,
-$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2\beta}|z|^{-(1-\beta)}.$
+There exists some constant such that for any $h\in(0,h_{0})$, for any $z\in\Gamma$, any $f\in L^{2}(\mathcal{M})$, for any $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$, for any $\beta\in[0,1]$ such that $p\in[0,(1+\beta)/2]$,
+(24) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-P_{h}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}|z|^{-(1+p-\beta)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
 We now provide a proof for Proposition 2.4.
@@
 Proof of Proposition 2.4.
-Let $f\in L^{2}(\mathcal{M})$.
-For any $z\in\Gamma$, set $e_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f$, which yields, using the integral representations of $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$,
-$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)e_{h}(z)\,\mathrm{d}z.$
+Let $h\in(0,h_{0})$, $f\in L^{2}(\mathcal{M})$ and let $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$.
+We can therefore introduce $u=\mathcal{L}^{p}f\in L^{2}(\mathcal{M})$.
+First, note using the triangle inequality, we have
+$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$ $\leq\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
+$\quad+\|P_{h}\gamma(\mathcal{L})f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=S_{1}+S_{2}$
+where we take $S_{1}=\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$ and $S_{2}=\|P_{h}\gamma(\mathcal{L})f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$.
+We now bound these two terms.
+For the term $S_{2}$, let us introduce the function $\widehat{\gamma}$ defined as $\widehat{\gamma}(\lambda)=\gamma(\lambda)\lambda^{-p}$.
+In particular, note that since $\gamma$ is an $\alpha$-amplitude spectral density, $\widehat{\gamma}$ is an $(\alpha+p)$-amplitude spectral density, and we have, by definition of $u$,
+$S_{2}=\|(I-P_{h})\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=\|(I-P_{h})\widehat{\gamma}(\mathcal{L})u\|_{L^{2}(\mathcal{M})}$
+We then use the Bramble–Hilbert lemma (cf.
+Lemma B.4) with $t=2\min\{\alpha+p;1\}\in(0,2]$, while noting that, following Remark 2.2, $\|\mathcal{L}^{t/2}(\widehat{\gamma}(\mathcal{L})u)\|_{L^{2}(\mathcal{M})}\lesssim\|u\|_{L^{2}(\mathcal{M})}<\infty$ (since $t/2\leq\alpha+p$):
+$S_{2}\lesssim h^{t}\|\mathcal{L}^{t/2}(\widehat{\gamma}(\mathcal{L})u)\|_{L^{2}(\mathcal{M})}\lesssim h^{t}\|u\|_{L^{2}(\mathcal{M})}=h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
+Let us now bound the term $S_{1}$.
+For any $z\in\Gamma$, set $\mathcal{F}_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}f-P_{h}(z-\mathcal{L})^{-1}f$, which yields, using the integral representations of $\gamma(\mathcal{L})$ and $\gamma(\mathcal{L}_{h})$,
+$\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f$ $=\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}_{h}(z)\,\mathrm{d}z.$
 The definition of $\Gamma$ and its parametrization allow to decompose the integral as
-$\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f$ $=\frac{-1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{+}(t))e_{h}(g_{+}(t))g^{\prime}_{+}(t)\,\mathrm{d}t-\frac{1}{2\pi i}\int_{-\theta}^{\theta}\gamma(g_{0}(t))e_{h}(g_{0}(t))g_{0}^{\prime}(t)\,\mathrm{d}t$
-$\quad+\frac{1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{-}(t))e_{h}(g_{-}(t))g^{\prime}_{-}(t)\,\mathrm{d}t,$
+$\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f$ $=\frac{-1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{+}(t))\mathcal{F}_{h}(g_{+}(t))g^{\prime}_{+}(t)\,\mathrm{d}t$
+$\quad-\frac{1}{2\pi i}\int_{-\theta}^{\theta}\gamma(g_{0}(t))\mathcal{F}_{h}(g_{0}(t))g_{0}^{\prime}(t)\,\mathrm{d}t$
+$\quad+\frac{1}{2\pi i}\int_{\delta_{0}}^{\infty}\gamma(g_{-}(t))\mathcal{F}_{h}(g_{-}(t))g^{\prime}_{-}(t)\,\mathrm{d}t,$
 where we recall that $g_{+}^{\prime}(t)=e^{i\theta},\leavevmode\nobreak\ g_{-}^{\prime}(t)=e^{-i\theta}$ and $g_{0}^{\prime}(t)=i\delta_{0}e^{it}$.
 Taking norms on both sides of this equality and using the triangle inequality then gives
-$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
-$\quad\leq\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{+}(t))|\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{-\theta}^{\theta}|\gamma(g_{0}(t))|\|e_{h}(g_{0}(t)\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
-$\qquad+\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{-}(t))|\left\|e_{h}(g_{-}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
+$S_{1}=\|\gamma(\mathcal{L}_{h})P_{h}f-P_{h}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}$
+$\quad\leq\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{+}(t))|\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{-\theta}^{\theta}|\gamma(g_{0}(t))|\|\mathcal{F}_{h}(g_{0}(t))\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
+$\qquad+\frac{1}{2\pi}\int_{\delta_{0}}^{\infty}|\gamma(g_{-}(t))|\left\|\mathcal{F}_{h}(g_{-}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t$
 $\quad=I_{+}+I_{0}+I_{-}.$
-Let us start by bounding $I_{0}$. We apply Equation 7 and Lemma 2.4 with $\beta=1$ to obtain
-$I_{0}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}|g_{0}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}\delta_{0}^{-\alpha}\,\mathrm{d}t\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}.$
-To bound $I_{+}$, we distinguish between the three cases $\alpha>1$, $\alpha<1$ and $\alpha=1$. When $\alpha>1$, we apply Equation 7 and Lemma 2.4 with $\beta=1$ to get
-$I_{+}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\int_{\delta_{0}}^{\infty}|t|^{-\alpha}\,\mathrm{d}t=C\frac{\delta_{0}^{-(\alpha-1)}}{\alpha-1}h^{2}\|f\|_{L^{2}(\mathcal{M})}.$
-For $\alpha<1$, recall that $h<h_{0}<1$ and assume without loss of generality that $h_{0}^{-2}>\delta_{0}$. We split $I_{+}=I_{+1}+I_{+2}$ with
-$I_{+1}=\int_{\delta_{0}}^{h^{-2}}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,I_{+2}=\int_{h^{-2}}^{\infty}|\gamma(g_{+}(t))|\,\left\|e_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t.$
-Using Equation 7 and Lemma 2.4 with $\beta=1$ we bound $I_{+1}$ by
-$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|g_{+}(t)|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t$
-$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{0}^{h^{-2}}|t|^{-\alpha}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\frac{h^{-2(1-\alpha)}}{1-\alpha}\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
-We proceed in the same manner to bound $I_{+2}$, but use Lemma 2.4 with $\beta=0$,
-$I_{+2}\leq C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|g_{+}(t)|^{-\alpha}|g_{+}(t)|^{-1}\,\mathrm{d}t=C\|f\|_{L^{2}(\mathcal{M})}\int_{h^{-2}}^{\infty}|t|^{-\alpha-1}\,\mathrm{d}t\leq Ch^{2\alpha}\|f\|_{L^{2}(\mathcal{M})}.$
-For $\alpha=1$ and the same splitting $I_{+}=I_{+1}+I_{+2}$ we obtain for $I_{+1}$ with Lemma 2.4 for $\beta=1$
-$I_{+1}$ $\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2}}t^{-1}\,\mathrm{d}t=Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(\log(h^{-2})-\log\delta_{0})$
-$\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log(h^{-2})|+|\log\delta_{0}|)\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}(|\log h|+|\log\delta_{0}|).$
-Assuming again $h<h_{0}$ and $h_{0}^{-2}>\delta_{0}$, we finally bound
-$I_{+1}$ $\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h}\right|\right)\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}\left(1+\left|\frac{\log\delta_{0}}{\log h_{0}}\right|\right)$
-$\leq Ch^{2}|\log h|\,\|f\|_{L^{2}(\mathcal{M})}.$
-The bound for $I_{+2}\leq Ch^{2}\|f\|_{L^{2}(\mathcal{M})}$ is obtained in the same way as in the case $\alpha>1$.
-Hence, $I_{+}\leq Ch^{2}|\log h|\leavevmode\nobreak\ \|f\|_{L^{2}(\mathcal{M})}$ for $\alpha=1$.
-Finally, note that by symmetry, $I_{-}$ satisfies the same bounds as $I_{+}$, which finishes the proof adding up all terms. ∎
+Let us start by bounding $I_{0}$. We apply Equation 7 and Equation 24 (with $\beta=1$) to obtain
+$I_{0}$ $\lesssim\int_{-\theta}^{\theta}|g_{0}(t)|^{-\alpha}\leavevmode\nobreak\ h^{2}|g_{0}(t)|^{-p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{-\theta}^{\theta}\delta_{0}^{-(\alpha+p)}\,\mathrm{d}t$
+$\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
+To bound $I_{+}$, we distinguish between the three cases $\alpha+p>1$, $\alpha+p<1$ and $\alpha+p=1$. When $\alpha+p>1$, we apply Equation 7 and Equation 24 (with $\beta=1$) to get
+$I_{+}$ $\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|g_{+}(t)|^{-(\alpha+p)}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{\infty}|t|^{-(\alpha+p)}\,\mathrm{d}t$
+$=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{\delta_{0}^{-(\alpha+p-1)}}{\alpha+p-1}.$
+Hence, we can conclude that if $\alpha+p>1$, we have $I_{+}\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
+For $\alpha+p<1$, recall that $10<h<h_{0}<$ 1 for some fixed $h_{0}$ such that $h_{0}^{-2}>\delta_{0}$. We then introduce $\tilde{\alpha}=(\alpha+p)/\alpha$ and split $I_{+}=I_{+1}+I_{+2}$ with
+$I_{+1}=\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|\gamma(g_{+}(t))|\,\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,$
+$I_{+2}=\int_{h^{-2\tilde{\alpha}}}^{\infty}|\gamma(g_{+}(t))|\,\left\|\mathcal{F}_{h}(g_{+}(t))\right\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t.$
+Let then $e_{h}=(1-(\alpha+p))/|\log h|\in(0,1-(\alpha+p))$ and $\beta=\alpha+p+e_{h}\in(0,1)$.
+Note in particular that $p\in[0,(1+\beta)/2]$ since $(1+\beta)/2-p=(1-p+\alpha+e_{h})/2>0$.
+We can therefore use Equation 7 and Equation 24 to bound $I_{+1}$ by
+$I_{+1}$ $\lesssim h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|g_{+}(t)|^{-(\alpha+1+p-\beta)}\,\mathrm{d}t=h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|t|^{-(1-e_{h})}\,\mathrm{d}t$
+$\lesssim h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{0}^{h^{-2\tilde{\alpha}}}|t|^{-(1-e_{h})}\,\mathrm{d}t=h^{2\beta}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{h^{-2\tilde{\alpha}e_{h}}}{e_{h}}$
+$=h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\frac{h^{-2(\tilde{\alpha}-1)e_{h}}}{e_{h}}\lesssim h^{2(\alpha+p)}|\log h|\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
+since $h^{-2\tilde{\alpha}e_{h}}=\exp(2\tilde{\alpha}(1-(\alpha+p)))\lesssim 1$. We proceed in the same manner to bound $I_{+2}$, using this time Equation 24 (with $\beta=p$):
+$I_{+2}$ $\lesssim h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|g_{+}(t)|^{-(\alpha+1)}\,\mathrm{d}t=h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|t|^{-(\alpha+1)}\,\mathrm{d}t$
+$\lesssim h^{p}h^{2\tilde{\alpha}\alpha}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\lesssim h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
+Hence, we conclude that if $\alpha+p<1$, we have $I_{+}\lesssim|\log h|\leavevmode\nobreak\ h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
+Finally, for $\alpha+p=1$ we repeat the same approach as for the case $\alpha+p<1$. On the one hand we use Equation 24 (with $\beta=1$) to get for the term $I_{+1}$ the bound
+$I_{+1}$ $\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{\delta_{0}}^{h^{-2\tilde{\alpha}}}|t|^{-1}\,\mathrm{d}t=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}(\log(h^{-2\tilde{\alpha}})-\log\delta_{0})$
+$\lesssim h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}|\log h|,$
+where we used the fact that $h$ is upper-bounded by a fixed constant $h_{0}$ satisfying $h_{0}^{-2}>\delta_{0}$ to derive the last inequality.
+On the other hand, we once again use Equation 24 with $\beta=p$ to bound the term $I_{+2}$
+$I_{+2}$ $\lesssim h^{2p}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}\int_{h^{-2\tilde{\alpha}}}^{\infty}|t|^{-(\alpha+1)}\,\mathrm{d}t\lesssim h^{p}h^{2\tilde{\alpha}\alpha}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
+$\lesssim h^{2(\alpha+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}=h^{2}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
+Hence, we have $I_{+}\lesssim h^{2}|\log h|\leavevmode\nobreak\ \|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$ when $\alpha+p=1$.
+Putting together the three cases, we can therefore conclude that
+$I_{+}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
+where $C_{\alpha+p}(h)=|\log h|$ if $\alpha+p\leq 1$, and $C_{\alpha+p}(h)>1$ if $\alpha+p\neq 1$.
+Finally, note that by symmetry, $I_{-}$ satisfies the same bounds as $I_{+}$, meaning that we also get $I_{-}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$.
+In conclusion, putting together the estimates for $I_{0}$, $I_{+}$ and $I_{-}$ we get
+$S_{1}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
+and putting together the estimates pf $S_{1}$ and $S_{2}$ yields
+$\|\gamma(\mathcal{L}_{h})P_{h}f-\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}\lesssim C_{\alpha+p}(h)h^{2\min\{\alpha+p;1\}}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})},$
+which concludes the proof. ∎
 A result similar to Proposition 2.4 can be derived to quantify the error between functions of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$.
@@
 Proposition 2.6.
-Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. There exists a constant $C>0$ such that, for any $\tilde{f}\in S_{h}^{\ell}$,
-(24) $\left\|\left(\gamma(\mathcal{L}_{h})\tilde{f}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\right\|_{L^{2}(\mathcal{M}_{h})}\leq Ch^{2}\,\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})}.$
-This proposition can be seen as an extension of Bonito et al.
-(2024, Lemma 4.4) relying on extensions of the estimates proven in Bonito et al.
-(2024, Lemma A.1).
-Its proof is similar and therefore postponed to Appendix B.
+Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$. Then, there exists some constant such that for any $h\in(0,h_{0})$, for any $\tilde{f}\in S_{h}^{\ell}$,
+(26) $\left\|\left(\gamma(\mathcal{L}_{h})\tilde{f}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\tilde{f}^{-\ell})\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}\,\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})}.$
+This proposition can be seen as an extension of [6, Lemma 4.4] relying on extensions of the estimates proven in [6, Lemma A.1]. Its proof is similar and therefore postponed to Appendix B.
 3. Stochastic theory: random fields on surfaces and convergence of SFEM approximation
@@
 where $\alpha>d/2$.
-By selecting $D_{ij}(x)=\delta_{ij}$ and $V(x)=\kappa^{2}$ with $\kappa>0$ one recovers the classical, stationary Whittle–Matérn fields studied in various settings in for instance Cox and Kirchner (2020); Bolin et al.
-(2020); Bonito et al.
-(2024); Lindgren et al.
-(2011); Jansson et al.
-(2022); Whittle (1963).
+By selecting $D_{ij}(x)=\delta_{ij}$ and $V(x)=\kappa^{2}$ with $\kappa>0$ one recovers the classical, stationary Whittle–Matérn fields studied in various settings in for instance [9, 3, 6, 21, 18, 27].
 In Figures 4(a) and 4(b) we show this case with $\kappa^{2}=10$ for a rougher field with $\alpha=0.55$ and a smoother one with $\alpha=1.5$, respectively.
@@
 Figure 4. Some examples to highlight the influence of the different model parameters.
+Finally, we remark that other types of amplitude spectral densities can be considered.
+For instance, to generate Figure 1(c), we used an amplitude spectral density of the form $\gamma(\lambda)=\sin(\lambda)\lambda^{-0.6}$.
 3.2. Approximation of random fields with surface finite elements
-Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$.
-Following the approach presented by Lang and Pereira (2023), the field $\mathcal{Z}$ is approximated by an expansion similar to that of Equation 27, but involving only quantities defined on the polyhedral surface $\mathcal{M}_{h}$.
+Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$.
+Following the approach presented by [19], the field $\mathcal{Z}$ is approximated by an expansion similar to that of Equation 29, but involving only quantities defined on the polyhedral surface $\mathcal{M}_{h}$.
 More precisely, we define the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ of the field $\mathcal{Z}$ by the relation
@@
 where $\{W_{i}^{h}\}_{1\leq i\leq N_{h}}$ is a sequence of independent standard Gaussian random variables, whose precise definition is clarified later in this section, and $\{(\Lambda_{i}^{h},E_{i}^{h})\}_{1\leq i\leq N_{h}}$ are the eigenpairs of $\mathsf{L}_{h}$ introduced in Section 2.2.
-Then, as proven in Lang and Pereira (2023, Theorem 3.4), $\mathsf{Z}_{h}$ can be decomposed in the nodal basis $\{\psi_{i}\}_{1\leq i\leq N_{h}}$ of $S_{h}$ as
+Then, as proven in [19, Theorem 3.4], $\mathsf{Z}_{h}$ can be decomposed in the nodal basis $\{\psi_{i}\}_{1\leq i\leq N_{h}}$ of $S_{h}$ as
 $\mathsf{Z}_{h}=\sum_{i=1}^{N_{h}}Z_{i}\psi_{i},$
@@
 Such an operation would result in a prohibitive computational cost (of order $\mathcal{O}(N_{h}^{3})$ operations).
-To avoid this cost, we use the Chebyshev trick proposed by Lang and Pereira (2023, Section 4), and approximate $\mathsf{Z}_{h}$ by the field $\mathsf{Z}_{h,M}$ defined by
+To avoid this cost, we use the Chebyshev trick proposed by [19, Section 4], and approximate $\mathsf{Z}_{h}$ by the field $\mathsf{Z}_{h,M}$ defined by
 (33) $\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}P_{\gamma,M}(\Lambda_{i}^{h})W_{i}^{h}E_{i}^{h},$
@@
 On the one hand, one can take $\lambda_{\min}=V_{-}$.
-On the other hand, following Lang and Pereira (2023), a candidate for $\lambda_{\max}$ is obtained by applying the Gershgorin circle theorem to $\bm{S}$.
-Since $\mathsf{Z}_{h,M}$ is defined by just replacing the power spectral density $\gamma$ by the polynomial  $P_{\gamma,M}$, its expansion $\mathsf{Z}_{h,M}$ into the nodal basis,
-$\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}Z_{i}^{(M)}\psi_{i},$
+On the other hand, following [19], a candidate for $\lambda_{\max}$ is obtained by applying the Gershgorin circle theorem to $\bm{S}$.
+The field $\mathsf{Z}_{h,M}$, called Galerkin–Chebyshev approximation of $\mathcal{Z}$, is in essence defined by just replacing the amplitude spectral density $\gamma$ by the polynomial  $P_{\gamma,M}$ in the definition of SFEM–Galerkin approximation $\mathsf{Z}_{h}$.
+The expansion of $\mathsf{Z}_{h,M}$ into the nodal basis,
+(34) $\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}Z_{i}^{(M)}\psi_{i},$
 is such that the weights $(Z_{1}^{(M)},\dots,Z_{N_{h}}^{(M)})$ now form a centered Gaussian vector with covariance matrix $\bm{\Sigma}_{Z^{(M)}}$ given by
@@
 Indeed, the weights $(Z_{1}^{(M)},\dots,Z_{N_{h}}^{(M)})$ can be sampled through
-$\begin{pmatrix}Z_{1}^{(M)}\\ \vdots\\ Z_{N_{h}}^{(M)}\end{pmatrix}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}P_{\gamma,M}(\bm{S})\begin{pmatrix}\varepsilon_{1}\\ \vdots\\ \varepsilon_{N_{h}}\end{pmatrix}$
-where $\varepsilon_{1},\dots,\varepsilon_{N_{h}}$ are independent standard Gaussian random variables, and the matrix-vector product by $P_{\gamma,M}(\bm{S})$ can be computed iteratively while just requiring products between $\bm{S}$ and vectors.
+(36) $\begin{pmatrix}Z_{1}^{(M)}\\ \vdots\\ Z_{N_{h}}^{(M)}\end{pmatrix}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}P_{\gamma,M}(\bm{S})\bm{W}$
+where $\bm{W}=(w_{1},\dots,w_{N_{h}})^{T}$ is a vector of independent standard Gaussian random variables, and the matrix-vector product by $P_{\gamma,M}(\bm{S})$ can be computed iteratively while just requiring products between $\bm{S}$ and vectors.
 In the next two subsections, we provide error estimates quantifying the error between our target random field $\mathcal{Z}$ and its successive SFEM and polynomial approximations by $\mathsf{Z}_{h}$ and $\mathsf{Z}_{h,M}$.
@@
 Theorem 3.1.
-Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$.
-Then, there exists $h_{0}>0$, such that for any $h<h_{0}$ the strong approximation error of the random field $\mathcal{Z}$ by its discretization $\mathsf{Z}_{h}^{\ell}$ satisfies the bound
-(33) $\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
-where $K_{\alpha}(h)=1$ if $\alpha-d/4<2$, $K_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$, and $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ otherwise.
+Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$.
+Then, for any $h\in(0,h_{0})$, the strong approximation error of the random field $\mathcal{Z}$ by its discretization $\mathsf{Z}_{h}^{\ell}$ satisfies the bound
+(37) $\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
+where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
 To prove the strong error estimate, we rely on the deterministic error bounds proven in the previous section, and on several intermediate approximations of $\mathcal{Z}$ defined on the spaces $S_{h}^{\ell}$ (i.e., on $\mathcal{M}$) and $S_{h}$ (i.e., on $\mathcal{M}_{h}$).
@@
 This explains why we refer to it as a projected white noise.
+In particular, note that $\tilde{\mathcal{W}}\in L^{2}(\Omega;L^{2}(\mathcal{M}))$ and $\|\tilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{j=1}^{N_{h}}\xi_{j}e_{j}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=N_{h}<\infty$.
 Based on $\widetilde{\mathcal{W}}$, we can then introduce a first approximation, on the space $S_{h}^{\ell}$, of the field $\mathcal{Z}$.
@@
 In particular, it holds that the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ defined in Equation 30 satisfies
-(36) $\mathsf{Z}_{h}=\gamma(\mathsf{L}_{h})\mathsf{W}$
+(40) $\mathsf{Z}_{h}=\gamma(\mathsf{L}_{h})\mathsf{W},$
 where we take for any $j\in\{1,\dots,N_{h}\}$, $W_{j}^{h}=\beta_{j}$.
@@
 As $\widehat{\mathsf{W}}\in S_{h}$, we can expand it in the orthonormal basis $\{E_{j}^{h}\}_{1\leq i\leq N_{h}}$ to get
-$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h}$
+$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h},$
 where $\alpha_{j}=(\widehat{\mathsf{W}},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$.
@@
 where $\beta_{j}=(\mathsf{W},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma^{1/2}\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},\sigma^{-1/2}E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$, and the same computations as before yield that $\beta_{j}=\big{(}\mathcal{W},(\sigma^{-1/2}E_{j}^{h})^{\ell}\big{)}_{L^{2}(\mathcal{M})}$.
-Hence, we can conclude this time that for any $1\leq i,j\leq N_{h}$, $\beta_{j}$ is also normally distributed with mean 00, and that
+Hence, we conclude this time that for any $1\leq i,j\leq N_{h}$, $\beta_{j}$ is also normally distributed with mean 00, and that
 $\operatorname{\amsmathbb{E}}[\beta_{i}\beta_{j}]$ $=\left(\sigma(\sigma^{-1/2}E_{i}^{h}),(\sigma^{-1/2}E_{j}^{h})\right)_{L^{2}(\mathcal{M}_{h})}=\left(\sigma(\sigma^{-1/2}E_{i}^{h}),(\sigma^{-1/2}E_{j}^{h})\right)_{L^{2}(\mathcal{M}_{h})}=B_{ij},$
@@
 Using the intermediate approximations $\tilde{\mathcal{Z}}_{h}$ and $\widehat{\mathsf{Z}}_{h}$ and the equivalence of the $L^{2}$ norms on $\mathcal{M}$ and $\mathcal{M}_{h}$, we can upper-bound the error between the field $\mathcal{Z}$ and its SFEM–Galerkin approximation $\mathsf{Z}_{h}$ by
-$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}$ $\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
-$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\tilde{\mathcal{Z}}_{h}-\widehat{\mathsf{Z}}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\widehat{\mathsf{Z}}_{h}^{\ell}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
-$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}{-}\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}+\big{\|}\widehat{\mathsf{Z}}_{h}-\mathsf{Z}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
+[…]
+$\quad\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\tilde{\mathcal{Z}}_{h}-\widehat{\mathsf{Z}}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\widehat{\mathsf{Z}}_{h}^{\ell}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
+$\quad\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}{+}\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}{-}\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}{+}\big{\|}\widehat{\mathsf{Z}}_{h}-\mathsf{Z}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
 We derive error estimates for each one of the three terms obtained in the last inequality. We start with the term $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$.
 Lemma 3.4.
-It holds that
-$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
-where $C_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$ and $C_{\alpha}(h)=1$, otherwise.
+Let $h\in(0,h_{0})$. It holds that
+$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
+where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
 Proof.
@@
 Remark 3.2).
-Let $\eta=\alpha+d/4>d/2$, and let $\zeta=\alpha-d/4>0$, so that $\alpha=(\eta+\zeta)/2$.
-Further, define the function $\tilde{\gamma}(x)=\gamma(x)x^{\eta/2}$.
-Note that, since $\gamma$ is an $\alpha$-power spectral density and since $\tilde{\gamma}$ decays as
-$|\tilde{\gamma}(x)|\lesssim|x|^{-\left(\alpha-\eta/2\right)}=|x|^{-\zeta/2},$
-$\tilde{\gamma}$ is a $(\zeta/2)$-power spectral density. Now, by definition of $\tilde{\gamma}$,
-$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $=\|\tilde{\gamma}(\mathcal{L})\mathcal{L}^{-\eta/2}\mathcal{W}-\tilde{\gamma}(\mathcal{L}_{h})\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
-so that the triangle inequality yields
-$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq E_{1}+E_{2},$
+Recall that we consider $h\in(0,h_{0})$ (meaning in particular that $|\log h|>1$).
+Similarly to the proof of [6, Theorem 5.2], we introduce $\varepsilon_{h}=\min\{d/2;\alpha-d/2\}/(2|\log h|)$, $\eta=2(d/4+\varepsilon_{h})$, and let $\zeta=2(\alpha-d/4-\varepsilon_{h})>0$, so that $\alpha=(\eta+\zeta)/2$.
+Note in particular that $\varepsilon_{h}>0$ since $d\in\{1,2\}$ and $\alpha>1>d/2$, and $\varepsilon_{h}<(\alpha-d/2)/2<\alpha-d/2<\alpha-d/4$.
+Besides, $\eta>d/2$ and $\zeta>0$.
+We distinguish two cases: if $\alpha\leq 1$, and if $\alpha>1$.
+Case $\alpha\leq 1$
+The proof follows the same approach as [6, Theorem 5.2]. Indeed, note that the triangle inequality yields
+$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}=\|\gamma(\mathcal{L})\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq S_{1}+S_{2},$
 where we take
-$E_{1}$ $=\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
-$E_{2}$ $=\|\tilde{\gamma}(\mathcal{L}_{h})\big{(}P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}-\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\big{)}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
-For the term $E_{1}$, note that by Bonito et al. (2024, Lemma 4.2),
-$\|\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq\frac{\eta}{\eta-d/2}.$
-In particular, $\mathcal{L}^{-\eta/2}\mathcal{W}$ is in $L^{2}(\Omega;L^{2}(\mathcal{M}))$ and an application of Proposition 2.3 yields that
-$E_{1}\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}},$
-where
-$C_{\zeta}(h)=\begin{cases}1,&\zeta\neq 2,\\ |\log h|,&\zeta=2.\\ \end{cases}$
-For $E_{2}$, we first write $P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}=\sum_{i=1}^{N_{h}}\beta_{i}e_{i}^{h}$, where $\beta_{i}=(P_{h}\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=(\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}W_{k}(e_{k},e_{i}^{h})_{L^{2}(\mathcal{M})}$.
-Further, following the definition of $\tilde{\mathcal{W}}$ in Equation 34, we have $\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\eta/2}\xi_{i}e_{i}^{h}$ where $\xi_{i}=(e_{i}^{h},\mathcal{W})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}W_{k}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}$.
-Note in particular that both $\beta_{i}$ and $\xi_{i}$ are real-valued (since the eigenbases $\{e_{k}\}_{k\in\amsmathbb{N}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are real-valued).
-Thus,
-$E_{2}^{2}=\sum_{i=1}^{N_{h}}|\tilde{\gamma}(\lambda_{i}^{h})|^{2}\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}],$
-where we use the orthonormality of the eigenfunctions. Note that
-$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]-2\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]+\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}].$
-First, by the independence of the sequence $\{W_{k}\}_{k=1}^{\infty}$ and $(\lambda_{k},e_{k})$ eigenpairs of $\mathcal{L}$,
-$\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]$ $=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}|(e_{k},e_{i}^{h})|^{2}_{L^{2}(\mathcal{M})}=\left(\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}e_{k},e_{i}^{h}\right)_{L^{2}(\mathcal{M})}$
-$=(\mathcal{L}^{-\eta}e_{i}^{h},e_{i}^{h})_{L^{2}(\mathcal{M})}=\|\mathcal{L}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}.$
-Similarly, and using the fact that $(\lambda_{i}^{h},e_{i}^{h})$ is an eigenpair of $\mathcal{L}_{h}$,
-$\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}]=(\lambda_{i}^{h})^{-\eta}\sum_{k=1}^{\infty}|(e_{i}^{h},e_{k})|^{2}_{L^{2}(\mathcal{M})}=(\lambda_{i}^{h})^{-\eta}\|e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}=\|\mathcal{L}_{h}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}$
-and
-$\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]=(\lambda_{i}^{h})^{-\eta/2}\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}|(e_{k},e_{i}^{h})|^{2}=(\mathcal{L}^{-\eta/2}e_{i}^{h},\mathcal{L}^{-\eta/2}_{h}e_{i}^{h})_{L^{2}(\mathcal{M})}.$
-This implies that by Proposition 2.3,
-$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\|\mathcal{L}^{-\eta/2}e_{i}^{h}-\mathcal{L}^{-\eta/2}_{h}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2},$
-where
-$C_{\eta}(h)=\begin{cases}1,&\eta\neq 2,\\ |\log h|,&\eta=2.\end{cases}$
-Hence, it holds that
-$E_{2}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}.$
-It remains to bound $\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$. To this end, recall Equation 12 and Equation 6 which yield that
-$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\zeta}\lesssim\sum_{i=1}^{N_{h}}i^{-2\zeta/d}\lesssim\begin{cases}N_{h}^{1-2\zeta/d},&2\zeta/d<1,\\ \log(N_{h}),&2\zeta/d=1,\\ 1,&2\zeta/d>1,\end{cases}\lesssim\begin{cases}h^{-2(d/2-\zeta)},&\zeta<d/2,\\ |\log(h)|,&\zeta=d/2,\\ 1,&\zeta>d/2,\end{cases}$
-where we used in the last inequality that $N_{h}\propto h^{-d}$. We obtain
-$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim h^{2(\zeta-\max\{\zeta;d/2\})}K_{\zeta}(h)^{2}$
-with
-$K_{\zeta}(h)=\begin{cases}1,&\zeta\neq d/2\\ |\log(h)|^{1/2},&\zeta=d/2\end{cases}$
-We conclude that
-$E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}}$
-and in particular,
-$h^{-\min\{\zeta;2\}}E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\beta},$
-where we take $\beta=\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}-\min\{\zeta;2\}$.
-We have two cases.
-On the one hand, if $\zeta\geq 2$ then, since $\eta>\zeta$, we also have $\eta>2$ and therefore $\beta=0$, $C_{\eta}(h)=1$ and $K_{\zeta}(h)=1$ (given that $d/2<2$).
-On the other hand, if $\zeta<2$, then $\max\{\zeta;d/2\}<2$ and $\max\{\zeta;d/2\}<\eta$ and therefore $\beta=\min\{\eta;2\}-\max\{\zeta;d/2\}>0$.
-In turn, this means that $C_{\eta}(h)K_{\zeta}(h)h^{\beta}\lesssim|\log(h)|^{3/2}|h^{\beta}\lesssim 1$.
-Hence, in both cases
-$h^{-\min\{\zeta;2\}}E_{2}\lesssim 1,$
-and we conclude that
-$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\lesssim h^{\min\{\zeta;2\}}\left(C_{\zeta}(h)+h^{-\min\{\zeta;2\}}E_{2}\right)\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}}.$
-The result follows by inserting the definition of $\zeta$. ∎
-For the error between $\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}$ and $\widehat{\mathsf{Z}}_{h}$ we get the following estimate, inspired by Bonito et al. (2024, Lemma 4.4).
+$S_{1}$ $=\|\gamma(\mathcal{L})\mathcal{W}-\gamma(\mathcal{L})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},\quad S_{2}=\|\gamma(\mathcal{L})P_{h}\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
+To bound the term $S_{1}$, we use
+$S_{1}^{2}$ $=\|\gamma(\mathcal{L})(I-P_{h})\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{i\in\amsmathbb{N}}\gamma(\lambda_{i})W_{i}(I-P_{h})e_{i}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
+$=\sum_{i\in\amsmathbb{N}}\gamma(\lambda_{i})^{2}\|(I-P_{h})e_{i}\|_{L^{2}(\mathcal{M})}^{2},$
+and since $\gamma$ is an $\alpha$-amplitude spectral density, we get
+(41) $S_{1}^{2}\lesssim\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2\alpha}\|(I-P_{h})e_{i}\|_{L^{2}(\mathcal{M})}^{2},$
+Applying the Bramble–Hilbert lemma (69) with $t=2((\alpha-d/4)-\varepsilon_{h})\in(0,2)$ together with Equation 41 gives
+$S_{1}^{2}\lesssim\sum_{i\in\amsmathbb{N}}h^{2t}\lambda_{i}^{-2\alpha+t}=h^{4(\alpha-d/4-\varepsilon_{h})}\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2(d/4+\varepsilon_{h})}\lesssim h^{4(\alpha-d/4)}\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2(d/4+\varepsilon_{h})},$
+since $h^{-4\varepsilon_{h}}=e^{4(\alpha-d/4)}\lesssim 1$. We can then write, using Equation 6,
+$S_{1}^{2}$ $\lesssim h^{4(\alpha-d/4)}\sum_{i\in\amsmathbb{N}}i^{-(1+4\varepsilon_{h}/d)}\lesssim h^{4(\alpha-d/4)}\int_{1}^{\infty}t^{-(1+4\varepsilon_{h}/d)}\,\mathrm{d}t$
+$\lesssim\frac{1}{4\varepsilon_{h}/d}h^{4(\alpha-d/4)}\lesssim|\log h|h^{4(\alpha-d/4)},$
+where we used the definition of $\varepsilon_{h}$ for the last inequality. So, we conclude $S_{1}\lesssim|\log h|^{1/2}h^{2(\alpha-d/4)}$.
+For the term $S_{2}$, we note that
+$S_{2}=\|\gamma(\mathcal{L})P_{h}\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}(P_{h}\mathcal{W})\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
+where, following Equation 38, $\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{j=1}^{N_{h}}\xi_{j}e_{j}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=N_{h}<\infty$ (cf.
+Remark 3.2).
+Hence, we can use Proposition 2.4 to deduce that since $\alpha\leq 1$,
+$S_{2}\lesssim|\log h|h^{2\alpha}\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
+Note then that since $N_{h}\propto h^{-d}$ (uniform mesh assumption), we can write $\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}=N_{h}^{1/2}\lesssim h^{-d/2}$.
+Therefore, we can conclude that
+$S_{2}\lesssim|\log h|h^{2\alpha-d/2}.$
+Putting together the bounds obtained for $S_{1}$ and $S_{2}$, we can then conclude that if $\alpha\leq 1$,
+$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq S_{1}+S_{2}\lesssim|\log h|h^{2\alpha-d/2}=|\log h|h^{2\min\{(\alpha-d/4);1\}},$
+where we use the fact that $\alpha-d/4<\alpha<1$ to derive the last inequality.
+Case $\alpha>1$
+We now assume that $\alpha>1$.
+We define the function $\tilde{\gamma}(x)=\gamma(x)x^{\zeta/2}$.
+Note that $\gamma$ is an $\alpha$-amplitude spectral density and $\tilde{\gamma}$ decays as $|\tilde{\gamma}(x)|\lesssim|x|^{-\left(\alpha-\zeta/2\right)}=|x|^{-\eta/2}$.
+Hence, $\tilde{\gamma}$ is a $(\eta/2)$-amplitude spectral density.
+Now, by definition of $\tilde{\gamma}$,
+$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $=\|\tilde{\gamma}(\mathcal{L})\mathcal{L}^{-\zeta/2}\mathcal{W}-\tilde{\gamma}(\mathcal{L}_{h})\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
+so that the triangle inequality yields $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq E_{1}+E_{2}$ where we take
+$E_{1}$ $=\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
+$E_{2}$ $=\|\tilde{\gamma}(\mathcal{L}_{h})\big{(}P_{h}\mathcal{L}^{-\zeta/2}\mathcal{W}-\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W}\big{)}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
+For the term $E_{1}$, we first note that $\zeta=d/2+2(\alpha-d/2-\varepsilon_{h})>d/2$ since $\varepsilon_{h}<\alpha-d/2$.
+Therefore, we can use [6, Lemma 4.2] and the fact that $h$ (and therefore $\varepsilon_{h}$) is upper-bounded to deduce that
+$\|\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\leq\frac{\zeta}{\zeta-d/2}\lesssim 1.$
+In particular, $\mathcal{L}^{-\zeta/2}\mathcal{W}$ is in $L^{2}(\Omega;L^{2}(\mathcal{M}))$.
+Let then $p=\min\{\alpha-d/2;1\}-2\varepsilon_{h}$.
+In particular, we have $p\in(0,1)$ since $p\leq 1-2\varepsilon_{h}<1$ and $p>\min\{\alpha-d/2;1\}-\min\{\alpha-d/2;d/2\}\geq 0$.
+Moreover, we have $\zeta/2-p=d/4+(\alpha-d/2)-\min\{\alpha-d/2;1\}+\varepsilon_{h}\geq d/4+\varepsilon_{h}>d/4$.
+Hence, using once again [6, Lemma 4.2], we can conclude that
+$\|\mathcal{L}^{p}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$ $=\|\mathcal{L}^{-(\zeta/2-p)}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
+$\lesssim\frac{\zeta/2-p}{\zeta/2-p-d/4}\lesssim\begin{cases}|\log h|,&\text{if }\alpha-d/2\leq 1\\ 1,&\text{if }\alpha-d/2>1\end{cases}$
+where we used the definition of $\varepsilon_{h}$ to derive the last inequality. Hence, we can apply Proposition 2.4 to get
+$E_{1}^{2}$ $=\operatorname{\amsmathbb{E}}[\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\mathcal{M})}^{2}]$
+$\lesssim C_{\eta/2+p}(h)^{2}h^{4\min\{\eta/2+p;1\}}\operatorname{\amsmathbb{E}}[\|\mathcal{L}^{p}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\mathcal{M})}^{2}]$
+$\lesssim C_{\eta/2+p}(h)^{2}h^{4\min\{\eta/2+p;1\}}K^{\prime}_{\alpha}(h)^{2},$
+where we take on the one hand, $C_{\eta/2+p}(h)=|\log h|$ if $\eta/2+p\leq 1$, and $C_{\eta/2+p}(h)=1$ otherwise, and on the other hand $K^{\prime}_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\leq 1+d/2$ and $K^{\prime}_{\alpha}(h)=1$ otherwise.
+Finally, note that according to Lemma A.1, $\eta/2+p=\min\{\alpha-d/2;1\}+d/4-\varepsilon_{h}\geq\min\{\alpha-d/4;1\}-\varepsilon_{h}$ which implies that $h^{4\min\{\eta/2+p;1\}}\leq h^{4(\min\{\alpha-d/4;1\}-\varepsilon_{h})}=h^{4\min\{\alpha-d/4;1\}}\exp(2\min\{d/2;\alpha-d/2\})$.
+Besides, we note that if $\alpha>1+d/2$, then $\eta/2+p=1+d/4-\varepsilon_{h}>1$ (since by construction $\varepsilon_{h}<d/4$), and therefore $C_{\eta/2+p}(h)=1$.
+We can therefore conclude that
+$E_{1}\lesssim K_{1}(h)h^{2\min\{\alpha-d/4;1\}}$
+where $K_{1}(h)=|\log h|^{3/2}$ if $1<\alpha\leq 1+d/4+\varepsilon_{h}$, and $K_{1}(h)=|\log h|^{1/2}$ if $21+d/4+\varepsilon_{h}<\alpha\leq 1+d/$ 2 and $K_{1}(h)=1$ if $\alpha>1+d/2$.
+For $E_{2}$, we first use the definition of the projection operator $P_{h}$ and the self-adjointness of $\mathcal{L}$ and $\mathcal{L}_{h}$ (which is a consequence of the definition of the associated bilinear form $\mathsf{A}_{\mathcal{M}}$)
+$E_{2}^{2}$ $=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}P_{h}\mathcal{L}^{-\zeta/2}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
+$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}\mathcal{L}^{-\zeta/2}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}P_{h}\mathcal{W},\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
+$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}\mathcal{W},\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
+$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}.$
+Using the orthonormality of the basis $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ and the definition of the white noise $\mathcal{W}$, we conclude that
+$E_{2}^{2}$ $=\operatorname{\amsmathbb{E}}\bigg{[}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}|\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}|^{2}\bigg{]}$
+$=\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}.$
+On the one hand, using Proposition 2.4 with $f=e_{i}^{h}\in S_{h}^{\ell}$ ($1\leq i\leq N_{h}$), which has norm $\|e_{i}^{h}\|_{L^{2}(\mathcal{M})}=1$, we have
+$\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}\lesssim C_{\zeta}(h)h^{2\min\{\zeta/2;1\}},\quad\text{where}\quad C_{\zeta}(h)=\begin{cases}1,&\zeta>2,\\ |\log h|,&\zeta\leq 2.\end{cases}$
+Note that by definition of $\zeta$ and using Lemma A.1, we have $\min\{\zeta/2;1\}\geq\min\{\alpha-d/4;1\}-\varepsilon_{h}>0$ (since $\varepsilon_{h}\in(0,\alpha-d/4)$), which in turn gives $h^{2\min\{\zeta/2;1\}}\leq h^{2(\min\{\alpha-d/4;1\}-\varepsilon_{h})}\lesssim h^{2\min\{\alpha-d/4;1\}}$.
+Hence, we can conclude that
+$\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}\lesssim C_{\zeta}(h)h^{2\min\{\alpha-d/4;1\}}$
+and therefore, $E_{2}^{2}\lesssim C_{\zeta}(h)^{2}h^{4\min\{\alpha-d/4;1\}}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$.
+On the other hand, the term $\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$can be bounded using Equation 12 and Equation 6 which yield that
+$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\eta}\lesssim\sum_{i=1}^{N_{h}}i^{-2\eta/d}\lesssim\int_{1}^{\infty}t^{-2\eta/d}=\frac{1}{2\eta/d-1}\lesssim|\log h|,$
+since $12\eta/d=1+4\varepsilon_{h}/d>$ 1 and by definition of $\varepsilon_{h}$. Hence, we can conclude that
+$E_{2}$ $\lesssim C_{\zeta}(h)|\log h|^{1/2}h^{2\min\{\alpha-d/4;1\}}\lesssim K_{2}(h)h^{2\min\{\alpha-d/4;1\}},$
+where $K_{2}(h)=C_{\zeta}(h)|\log h|^{1/2}$, i.e., $K_{2}(h)=|\log h|^{3/2}$ if $1<\alpha\leq 1+d/4+\varepsilon_{h}$, and $K_{2}(h)=|\log h|^{1/2}$ if $\alpha>1+d/4+\varepsilon_{h}$.
+Hence, by putting together the estimates for $E_{1}$ and $E_{2}$, and noting that $K_{1}(h)\leq K_{2}(h)$, we can conclude that if $\alpha>1$, then
+$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq E_{1}+E_{2}\lesssim K_{2}(h)h^{2\min\{\alpha-d/4;1\}}.$
+In particular note that for any $\alpha>1$, we can clearly upper-bound $K_{2}(h)$ by $K_{2}(h)\lesssim|\log h|^{3/2}$.
+But if moreover $\alpha\geq 1+d/2$, we have y definition of $\varepsilon_{h}$, $1+d/4+\varepsilon_{h}<1+d/4+\min\{d/2;\alpha-d/2\}/2\leq 1+d/2\leq\alpha$, and therefore $K_{2}(h)=|\log h|^{1/2}$.
+Hence, we can conclude that if $\alpha>1$, $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim K^{\prime}_{2}(h)h^{2\min\{\alpha-d/4;1\}}$ where $K^{\prime}_{2}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2 and $K^{\prime}_{2}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
+And finally gathering the estimates obtained in the two cases $\alpha\leq 1$ and $\alpha>1$, we obtain the result stated in the lemma.
+∎
+For the error between $\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}$ and $\widehat{\mathsf{Z}}_{h}$ we get the following estimate, inspired by [6, Lemma 4.4].
 Lemma 3.5.
@@
 Proof.
-Let $\mathcal{E}_{h}=\big{\|}\left(\tilde{\mathcal{Z}}_{h}\right)^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}=\big{\|}\left(\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$.
+Let
+$\mathcal{E}_{h}=\big{\|}\left(\tilde{\mathcal{Z}}_{h}\right)^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}=\big{\|}\left(\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
 We note that $\widetilde{\mathcal{W}}$ is an $S_{h}^{\ell}$-valued random variable.
@@
 First, if $d=1$, then we note that $\min\{\alpha+d/4;1\}\geq\alpha+d/4>d/2$ since $\alpha>d/4$.
-Besides, Bonito et al.
-(2024, Lemma 4.2) yields that for all $r\in(d/2,2)$, there is a constant $C>0$ such that
+Besides, [6, Lemma 4.2] yields that for all $r\in(d/2,2)$, there is a constant $C>0$ such that
 (42) $\|\mathcal{L}_{h}^{-r/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\leq C\frac{r}{r-d/2}.$
@@
 $\mathcal{E}_{h}$ $\lesssim h^{2}\frac{\min\{\alpha+d/4;1\}}{\min\{\alpha+d/4;1\}-d/2}\lesssim h^{2}.$
-If now $d=2$, since $\alpha+d/4>d/2=1$, we have $\min\{\alpha+d/4;1\}=1$. We then note that by the proof of Bonito et al. (2024, Lemma 4.2), for any $\varepsilon>0$
+If now $d=2$, since $\alpha+d/4>d/2=1$, we have $\min\{\alpha+d/4;1\}=1$. We then note that by the proof of [6, Lemma 4.2], for any $\varepsilon>0$
 $\|\mathcal{L}_{h}^{-1/2}\widetilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\lesssim\sum_{j=1}^{N_{h}}\lambda_{j}^{-1}\lesssim\lambda_{N_{h}}^{\varepsilon}\sum_{j=1}^{N_{h}}\lambda_{j}^{-1-\varepsilon}.$
@@
 Lemma 3.6.
-It holds that there is a constant $C>0$ such that
-$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq Ch^{2}.$
+Let $h\in(0,h_{0})$. It holds that there is a constant such that
+$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\lesssim h^{2}.$
 Proof.
@@
 $\left\|(\sigma^{1/2}-1)\right\|^{2}_{L^{\infty}(\mathcal{M}_{h})}\leq\left\|(\sigma-1)\right\|^{2}_{L^{\infty}(\mathcal{M}_{h})}\leq Ch^{2},$
-where Dziuk and Elliott (2013, Lemma 4.1) was applied in the final step. We conclude that
+where [11, Lemma 4.1] was applied in the final step. We conclude that
 $\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}\leq Ch^{4}\sum_{j=1}^{N_{h}}\gamma(\Lambda_{j}^{h})^{2},$
@@
 By summing the three estimates derived in Lemmas 3.6, 3.5 and 3.4, we get
-$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}+|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}.$
-Let then $\epsilon_{h}$ be the quantity defined by $\epsilon_{h}=\big{(}K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}\big{)}^{-1}\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$, which satisfies
-$\epsilon_{h}$ $\lesssim C_{\alpha}(h)K_{\alpha}(h)^{-1}+|\log(h)|^{(d-1)/2}K_{\alpha}(h)^{-1}h^{2-\min\{\alpha-d/4;2\}}.$
-If $\alpha-d/4\in(0,2)$, we have $C_{\alpha}(h)=K_{\alpha}(h)=1$ and $02-\min\{\alpha-d/4;2\}>$ 0, and therefore $\epsilon_{h}\lesssim 1+$.
-Similarly, if $\alpha-d/4=2$, it holds $C_{\alpha}(h)=K_{\alpha}(h)=|\log h|$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim 1+|\log(h)|^{(d-3)/2}\lesssim 1$.
-Finally, for $\alpha-d/4>2$, $C_{\alpha}(h)=1$, $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ and $02-\min\{\alpha-d/4;2\}=$ 0, and therefore $\epsilon_{h}\lesssim|\log h|^{-(d-1)/2}+1\lesssim 1$.
-In conclusion, we obtain for any $\alpha>d/4$, $\epsilon_{h}\lesssim 1$ which in turn yields
-$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}}.\qed$
+$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}+|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}+h^{2}.$
+Noe then that $2\min\{\alpha-d/4;1\}\leq 2$ and that $|\log(h)|^{(d-1)/2}\lesssim C_{\alpha}(h)$.
+Hence, we have the bound $|\log(h)|^{(d-1)/2}\leavevmode\nobreak\ h^{2}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}$ which allows us to conclude that
+$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
+and the proof is complete. ∎
 Having bounded the SFEM error, we are now ready to derive the additional error associated with the Chebyshev approximation which we use to compute SFEM–Galerkin approximations in practice.
-3.4. Error analysis of the Chebyshev approximation
-To bound the error between $\mathsf{Z}_{h,M}$ and $\mathsf{Z}_{h}$, we can directly apply Lang and Pereira (2023, Theorem 5.8). We obtain the following result.
-Theorem 3.7.
+3.4. Error analysis of the Galerkin–Chebyshev approximation
+The error between the field $\mathcal{Z}$ and its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M}$ described in Section 3.2 is derived by combining Theorem 3.1 with an error bound between the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ and the Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M}$ obtained using [19, Theorem 5.8].
+The latter bound is expressed in the next result.
+Lemma 3.7.
 Let $[\lambda_{\min},\lambda_{\max}]\subset(0,\infty)$ be the interval on which the Chebyshev polynomial approximation $P_{\gamma,M}$ is computed, and let $\xi=\lambda_{\min}/(\lambda_{\max}-\lambda_{\min})$.
@@
 with $\epsilon_{\xi}=\xi+\sqrt{\xi(2+\xi)}$. In particular, setting $\lambda_{\min}=V_{-}$ and $\lambda_{\max}=\Lambda_{N_{h}}^{h}$, the error is bounded by
-(39) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C_{V}^{-1}{h}^{-(d/2+1)}\exp(-C_{V}h^{-1}M)$
+(44) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq C_{V}^{-1}{h}^{-(d/2+1)}\exp(-C_{V}hM)$
 for some constant $C_{V}>0$ proportional to $\sqrt{V_{-}}$.
@@
 In particular, note that $E_{\xi}\subset H_{\pi/2}$ and for any $z\in E_{\xi}$, $\operatorname{Re}(z)\geq\lambda_{\min}/2>0$.
-Hence, since $\gamma$ is a power spectral density, by definition $\gamma$ is holomorphic and bounded inside $E_{\xi}$.
-We can then adapt the same proof as in Lang and Pereira (2023, Theorem 5.8) to obtain the stated proposition.
+Hence, since $\gamma$ is an amplitude spectral density, by definition $\gamma$ is holomorphic and bounded inside $E_{\xi}$.
+We can then adapt the same proof as in [19, Theorem 5.8] to obtain the stated proposition.
 ∎
-For a fixed mesh size $h$, the approximation error converges to 00 as the order of the polynomial approximation $M$ goes to infinity.
-Choosing $M$ as a function of $h$ that grows fast enough then allows us to ensure the convergence of the approximation error as $n$ goes to infinity (Lang and Pereira, 2023, Section 5.2).
+Hence, for a fixed mesh size $h$, the approximation error $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ converges to 00 as the order of the polynomial approximation $M$ goes to infinity.
+Choosing $M$ as a function of $h$ that grows fast enough then allows us to ensure the convergence of the approximation error as $n$ goes to infinity [19, Section 5.2].
+In particular, by taking $M=M_{h}=(d/2+3)C_{V}^{-1}h^{-1}|\log h|=(d/2+3)C_{V}^{-1}h^{-1}\log(h^{-1})$ the bound (44) becomes
+(45) $\|\mathsf{Z}_{h,M}-\mathsf{Z}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ $\leq C_{V}^{-1}\exp((d/2+1)\log(h^{-1})-(d/2+3)\log(h^{-1}))$
+$=C_{V}^{-1}h^{2}.$
+Putting together Theorem 3.1 and Equation 45, we can now state the following result which provides a bound between the field $\mathcal{Z}$ and its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M_{h}}$.
+Theorem 3.8.
+Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$, let $h\in(0,h_{0})$ and take $M_{h}=(d/2+3)C_{V}^{-1}h^{-1}|\log h|$.
+Then, The strong error between the random field $\mathcal{Z}$ by its Galerkin–Chebyshev approximation $\mathsf{Z}_{h,M_{h}}^{\ell}$ satisfies the bound
+(46) $\|\mathcal{Z}-\mathsf{Z}_{h,M_{h}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
+where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
+Remark 3.9.
+In practice, it is common to “chop” a Chebyshev polynomial approximation, i.e.
+to truncate the polynomial approximation at an early order.
+Indeed, let us assume that we have computed the coefficients $\{c_{k}\}_{0\leq k\leq M_{h}}$ of the Chebyshev series up to the order $M=M_{h}$ and let $c_{\max}=\max_{0\leq k\leq M_{h}}|c_{k}|$.
+Let $\varepsilon>0$ be fixed but arbitrary and assume that there exists $m_{\varepsilon}\in\{0,\dots,M_{h}-1\}$ such that for any $k>m_{\varepsilon}$, $|c_{k}|/c_{\max}<\varepsilon$.
+The error between the field $\mathsf{Z}_{h,M_{h}}$ and its “chopped” counterpart $\mathsf{Z}_{h,m_{\varepsilon}}$ is given by
+$\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$ $=\|\sum_{i=1}^{N_{h}}(P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h}))W_{i}^{h}E_{i}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}^{2}$
+$=\sum_{i=1}^{N_{h}}(P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h}))^{2}$
+where $P_{\gamma,M_{h}}(\lambda)=\sum_{k=0}^{M_{h}}c_{k}T_{k}(\lambda)$ is the $M_{h}$-th order Chebyshev approximation of $\gamma$ and $T_{k}$ denotes the $k$-th (shifted) Chebyshev polynomial.
+In particular, we have $|P_{\gamma,M_{h}}(\Lambda_{i}^{h})-P_{\gamma,m_{\varepsilon}}(\Lambda_{i}^{h})|=|\sum_{k=m_{\varepsilon}+1}^{M_{h}}c_{k}T_{k}(\lambda)|\leq\sum_{k=m_{\varepsilon}+1}^{M_{h}}|c_{k}|\leavevmode\nobreak\ |T_{k}(\lambda)|\leq\sum_{k=m_{\varepsilon}+1}^{M_{h}}|c_{k}|\leq(M_{h}-m_{\varepsilon})c_{\max}\varepsilon\leq M_{h}c_{\max}\varepsilon$.
+Hence, the error between $\mathsf{Z}_{h,M_{h}}$ and its “chopped” counterpart $\mathsf{Z}_{h,m_{\varepsilon}}$ satisfies
+$\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$ $\leq N_{h}^{1/2}M_{h}c_{\max}\varepsilon\lesssim h^{-(1+d/2)}|\log h|\leavevmode\nobreak\ \varepsilon$
+Hence, if we have $\varepsilon\lesssim|\log h|^{-1}h^{(3+d/2)}$ then we can conclude that $\|\mathsf{Z}_{h,M_{h}}-\mathsf{Z}_{h,m_{\varepsilon}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\lesssim h^{2}$, thus implying that once again we have $\|\mathcal{Z}-\mathsf{Z}_{h,m_{\varepsilon}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}}$.
+This means that in practice, if we notice that the coefficients of the Chebyshev polynomial approximation $\{c_{k}\}_{0\leq k\leq M_{h}}$ decay fast enough that we can find some $m_{\varepsilon}\in\{0,\dots,M_{h}-1\}$ such that for any $k>m_{\varepsilon}$, $|c_{k}|/c_{\max}<\varepsilon\lesssim|\log h|^{-1}h^{(3+d/2)}$, we can replace the field $\mathsf{Z}_{h,M_{h}}$ by its chopped counterpart while still maintaining the same strong error bound with respect to the field $\mathcal{Z}$.
+4. Numerical experiments
+In this section, we present numerical experiments confirming the strong error bound in Equation 46.
+We consider two cases: when the manifold $\mathcal{M}$ is a circle (hence $d=1$), and when the manifold $\mathcal{M}$ is a sphere (hence $d=2$).
+Indeed, in both cases we are able to easily define nested meshes of various sizes $h$, and define appropriate noise projections at the different levels.
+In each case, the field $\mathcal{Z}$ is approximated by a Galerkin–Chebyshev field $\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}$ computed at a very fine mesh size $h_{\text{fine}}$.
+More precisely, 25 samples of $\{\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}^{(i)}\}_{1\leq i\leq 25}$ of the field $\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}$ are computed.
+In particular, let us denote by $\mathsf{W}^{(i)}_{h_{\text{fine}}}$ the projected white noise used to compute the sample $\mathcal{Z}^{(i)}=P_{\gamma,M_{h_{\text{fine}}}}(\mathsf{L}_{h_{\text{fine}}})\mathsf{W}^{(i)}_{h_{\text{fine}}}$.
+For a coarse mesh size $h>h_{\text{fine}}$, each sample $\mathcal{Z}^{(i)}$ is compared to the sample of $\mathsf{Z}_{h,M_{h}}^{(i)}$ computed from the white noise $\mathsf{W}^{(i)}_{h}$ obtained by projecting the noise $\mathsf{W}^{(i)}_{h_{\text{fine}}}$ (defined in the fine mesh) down to the coarse mesh.
+Note that, following Remark 3.9, the Galerkin–Chebyshev fields are systematically computed by chopping the Chebyshev polynomial approximations at a small level $\epsilon=10^{-12}$.
+Besides, when computing the representation of a Galerkin–Chebyshev field in the nodal basis through (36), the matrix $\sqrt{\bm{C}}$ is computed as a Cholesky factorization of the mass matrix $\bm{C}$ in the first experiment, and is approximate using a (diagonal) mass lumping approach in the second experiment.
+Finally, the strong error at a mesh size $h$ is then estimated from the samples using
+(47) $\|\mathcal{Z}-\mathsf{Z}_{h,M_{h}}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\approx\bigg{(}\frac{1}{25}\sum_{i=1}^{25}\|\mathsf{Z}_{h_{\text{fine}},M_{h_{\text{fine}}}}^{(i)}-\mathsf{Z}_{h,M_{h}}^{(i)}\|_{L^{2}(\mathcal{M}_{h})}^{2}\bigg{)}^{1/2}.$
+In the first experiment, we consider a random field $\mathcal{Z}$ defined on the unit circle $\mathcal{M}$ with amplitude spectral density given by the formula $\gamma(\lambda)=(V_{0})^{\alpha}(\lambda+\cos(0.9\pi)\sqrt{\lambda})^{-\alpha}$ where $V_{0}=10^{4}$ and for $\alpha\in\{0.5,1.05,1.5\}$.
+The bilinear form $\mathcal{M}$ is taken as follows.
+The diffusion matrix $\mathcal{D}$ is defined by $\mathcal{D}(x)=I-\nu(x)\nu(x)^{T}$, $x\in\mathcal{M}$, where $\nu(x)=x/\|x\|$ is the normal vector to a point $x\in\mathcal{M}$ of the unit circle.
+The function $V$ is defined as $V(x)=3V_{0}$, $x\in\mathcal{M}$, if $\theta(x)\in]\pi/2,3\pi/2[$ and $V(x)=V_{0}$ otherwise, and $\theta(x)\in[0,2\pi[$ denotes the circular coordinate of the point $x\in\mathcal{M}$.
+The fine mesh size is taken to be $h_{\text{fine}}=2^{-19}\pi$ and the coarse mesh sizes are $h=\{2^{-13}\pi,2^{-14}\pi,2^{-15}\pi,2^{-16}\pi\}$.
+An example of sample (on the fine mesh) obtained with this choice of parameters is displayed in Figure 5(a).
+The results of this first experiment are presented in Figure 5(b), where the log of the strong error (47) is plotted against the log of the mesh size $h$, for the three amplitude spectral density obtained by taking $\alpha\in\{0.5,1.05,1.5\}$.
+As shown in the figure, we retrieve the rates (respectively $0.5$,$1.6$ and $2$) predicted by Theorem 3.8.
+(a) Sample (in green) of the random field with $\alpha=1.75$ on the fine circle mesh (in black). The distance to the circle indicates the value of taken by the field.
+(b) Strong error as a function the mesh size.
+Figure 5. First numerical experiment (on the circle).
+In the second experiment, we consider a random field $\mathcal{Z}$ defined on the unit sphere $\mathcal{M}$ with amplitude spectral density given by the formula $\gamma(\lambda)=C_{0}\lambda^{-\alpha}$ where $C_{0}=500$ and for $\alpha\in\{0.75,1.25,2.25\}$.
+The bilinear form $\mathcal{M}$ is taken as follows.
+The diffusion matrix $\mathcal{D}$ is defined $\mathcal{D}(x)=\nabla_{\mathcal{M}}f(x)(\nabla_{\mathcal{M}}f(x))^{T}+\rho(x)\nabla_{\mathcal{M}}^{\perp}f(x)(\nabla_{\mathcal{M}}^{\perp}f(x))^{T}$, where $f(x)=2\cos(\theta(x))\cos(\phi(x))\sin^{2}(\theta(x))$ and $\rho(x)=0.1+0.6/(1+e^{-4\cos(\theta(x))})$.
+Here $x\in\mathcal{M}$ and $(\theta(x),\phi(x))$ denotes spherical coordinates.
+The function $V$ is defined as $V(x)=500(1+5\cos^{2}(\pi\theta(x)))$.
+The fine mesh size is taken to be $h_{\text{fine}}=2.16\times 10^{-3}$ and the coarse mesh sizes are $h=\{3.45\times 10^{-2},1.73\times 10^{-2},8.63\times 10^{-3},4.32\times 10^{-3}\}$.
+An example of sample (on the fine mesh) obtained with this choice of parameters is displayed in Figure 6(a).
+The results of this second experiment are presented in Figure 6(b), where once again the log of the strong error (47) is plotted against the log of the mesh size $h$, for the three amplitude spectral density obtained by taking $\alpha\in\{0.75,1.25,2.25\}$.
+As seen in Figure 6(b), we retrieve the rates (respectively $0.5$,$1.5$ and $2$) predicted by Theorem 3.8.
+(a) Sample of the random field with $\alpha=1.75$ on the fine mesh. The colors indicate the value of taken by the field.
+(b) Strong error as a function the mesh size.
+Figure 6. Second numerical experiment (on the sphere).
 Appendix A Deterministic proofs
@@
 A standard result in the spectral theory of elliptic operators on compact Riemannian manifolds (see e.g.
-(Shubin, 2001, Section 8)) ensures that there exists a set of eigenpairs $\{(\lambda_{i},f_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ such that $0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{f_{i}\}_{i\in\amsmathbb{N}}$ is an orthonormal basis of $L^{2}(\mathcal{M})$ composed of possibly complex-valued functions.
+[23, Section 8]) ensures that there exists a set of eigenpairs $\{(\lambda_{i},f_{i})\}_{i\in\amsmathbb{N}}$ of $\mathcal{L}$ such that $0\leq\lambda_{1}\leq\lambda_{2}\leq\cdots$ with $\lambda_{i}\rightarrow+\infty$ as $i\rightarrow+\infty$, and $\{f_{i}\}_{i\in\amsmathbb{N}}$ is an orthonormal basis of $L^{2}(\mathcal{M})$ composed of possibly complex-valued functions.
 Hence, let us prove that $\lambda_{1}\geq\delta$ and that we can build an orthonormal basis $\{e_{i}\}_{i\in\amsmathbb{N}}$ of $L^{2}(\mathcal{M})$ such that each $e_{i}$ is real-valued and an eigenfunction of $\mathcal{L}$ with eigenvalue $\lambda_{i}$.
@@
 On the other hand, let $\lambda>0$ be one of the eigenvalues of $\mathcal{L}$, and $E_{\lambda}\subset L^{2}(\mathcal{M})$ the associated eigenspace.
-Following again the results from (Shubin, 2001, Section 8), we get that $E_{\lambda}\subset C^{\infty}(\mathcal{M})$, $\dim E_{\lambda}<\infty$ and that if $\lambda^{\prime}\neq\lambda$ is another eigenvalue of $\mathcal{L}$, then $E_{\lambda}$ and $E_{\lambda^{\prime}}$ are orthogonal.
+Following again the results from [23, Section 8], we get that $E_{\lambda}\subset C^{\infty}(\mathcal{M})$, $\dim E_{\lambda}<\infty$ and that if $\lambda^{\prime}\neq\lambda$ is another eigenvalue of $\mathcal{L}$, then $E_{\lambda}$ and $E_{\lambda^{\prime}}$ are orthogonal.
 Besides, $E_{\lambda}$ is in fact generated by the set $\{f_{j}\}_{j\in J_{\lambda}}$, where $J_{\lambda}=\{i\in\amsmathbb{N}:\lambda_{i}=\lambda\}$ is finite (since $\dim E_{\lambda}<\infty$).
@@
 ∎
-A.2. Proof of Lemma 2.4
+A.2. A useful inequality
+We end this section by introducing a lemma which will be used to obtain practical error bounds.
+Lemma A.1.
+For any $a,b\in\amsmathbb{R}$ and any $\varepsilon>0$, $\min\{a-\varepsilon;b\}+\varepsilon\geq\min\{a;b\}$.
 Proof.
-Let us start by proving the following claims: for any $z\in\Gamma$, $f\in L^{2}(\mathcal{M})$,
-(40) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq Ch^{2},$
-(41) $\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq\frac{C}{|z|}.$
-We omit the proof of Equation 41 and refer the reader to Fujita and Suzuki (1991, Theorem 5.1), where the proof may be adapted verbatim.
-To prove Equation 40, we can largely proceed as in Fujita and Suzuki (1991, Theorem 7.1), substituting the Euclidean elliptic regularity estimates and interpolation bounds with their surface counterparts, see for instance Taylor (2011, Chapter 11), Dziuk and Elliott (2013, Lemma 4.3) and Bonito et al.
-(2024, Equation (4.8)).
-We must in essence only adapt Fujita and Suzuki (1991, Lemma 7.1) to also hold for $z\in\Gamma_{0}$, as their estimate only holds for $z\in\partial G_{\theta}$.
-In other words, we must show that for any $z=\delta_{0}e^{it}$, where $t\in[-\theta,\theta]$, there is a constant $\tilde{\delta}>0$ such that for all $v\in H^{1}(\mathcal{M})$,
-(42) $|z|\|v\|_{L^{2}(\mathcal{M})}^{2}+\|v\|_{H^{1}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|z\|v\|_{L^{2}(\mathcal{M})}^{2}-\mathsf{A}_{\mathcal{M}}(v,v)\right|.$
-To prove this inequality, note that by coercivity of $\mathsf{A}_{\mathcal{M}}$,
-$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq\delta\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}.$
-Let then $\tilde{\delta}=\frac{\delta_{0}}{1+\delta_{0}}$. Since the only requirement on $\delta_{0}$ was that $\delta_{0}<\delta/2$, we can then write
-$\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\geq 2\delta_{0}\|v\|_{H^{1}(\mathcal{M})}^{2}-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-\tilde{\delta}(1+\delta_{0})\|v\|_{L^{2}(\mathcal{M})}^{2}$
-$=2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}+\left(\delta_{0}\tilde{\delta}-(2\delta_{0}+1)\tilde{\delta}\right)\|v\|_{L^{2}(\mathcal{M})}^{2}$
-$\geq 2\tilde{\delta}(1+\delta_{0})\|v\|_{H^{1}(\mathcal{M})}^{2}-(2\delta_{0}+1)\tilde{\delta}\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\tilde{\delta}\|v\|_{L^{2}(\mathcal{M})}^{2}$
-$=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right)=\tilde{\delta}\left(\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\right),$
-since $|z|=\delta_{0}$ if $z\in\Gamma_{0}$. Therefore,
-$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left(\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right).$
-This last inequality implies in particular that $\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\geq 0$ and therefore that
-$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-\delta_{0}\|v\|_{L^{2}(\mathcal{M})}^{2}\right|.$
-Using once again that $|z|=\delta_{0}$ and that $\mathsf{A}_{\mathcal{M}}(v,v)\geq 0$, we deduce that
-$\|v\|_{H^{1}(\mathcal{M})}^{2}+|z|\|v\|_{L^{2}(\mathcal{M})}^{2}$ $\leq\tilde{\delta}^{-1}\left||\mathsf{A}_{\mathcal{M}}(v,v)|-|z|\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|\leq\tilde{\delta}^{-1}\left|\mathsf{A}_{\mathcal{M}}(v,v)-z\,\|v\|_{L^{2}(\mathcal{M})}^{2}\right|$
-by the reverse triangle inequality.
-To conclude, we have proven that the relation (42) holds and equipped with this estimate, we can proceed as in the proof of Fujita and Suzuki (1991, Theorem 7.1) and obtain Equation 40.
-Finally, by interpolating between the bounds obtained in Equations 40 and 41, we obtain that, for any $\beta\in[0,1]$,
-$\|(z-\mathcal{L}_{h})^{-1}P_{h}f-(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}\leq C\frac{h^{2\beta}}{|z|^{1-\beta}}.\qed$
-Appendix B Error between functions of discrete operators
-The aim of this section is to prove Equation 24.
-To do so, we rely on a geometric consistency estimate, and on extensions of the results of Bonito et al.
-(2024, Lemma A.1), which we start by stating and proving.
+Indeed, if $a<b$, we have $\min\{a-\varepsilon;b\}+\varepsilon=a-\varepsilon+\varepsilon=a=\min\{a;b\}$.
+If $b\leq a<b+\varepsilon$, $\min\{a-\varepsilon;b\}+\varepsilon=a-\varepsilon+\varepsilon=a\geq b=\min\{a;b\}$.
+And finally if $a\geq b+\varepsilon$, $\min\{a-\varepsilon;b\}+\varepsilon=b+\varepsilon>b=\min\{a;b\}$.
+∎
+Appendix B Error estimates
 B.1. Geometric consistency estimate
 The following geometric consistency estimate quantifies the error between the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$.
-Its proof is a straightforward adaptation of the proof of Dziuk and Elliott (2013, Lemma 4.7) to account for the diffusion matrix $\mathcal{D}$.
+Its proof is a straightforward adaptation of the proof of [11, Lemma 4.7] to account for the diffusion matrix $\mathcal{D}$.
 Lemma B.1.
@@
 $\mathcal{M}_{h}$), and let $\mathcal{H}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ be the extended Weingarten map of $\mathcal{M}$ (cf.
-Dziuk and Elliott (2013, Definition 2.5)).
+[11, Definition 2.5]).
 Recall in particular that $\mathcal{H}(x)\nu(x)=0$ for any $x\in\mathcal{M}$, meaning in particular that $\mathcal{H}\Pi=\Pi\mathcal{H}=\mathcal{H}$.
@@
 where $d_{s}$ is the oriented distance function restricted to $\mathcal{M}_{h}$ and introduced in Section 2.2. On the one hand, note that for any $u_{h},v_{h}\in S_{h}$,
-$(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot\nabla_{\mathcal{M}_{h}}\overline{v}_{h}$ $=\big{(}\mathcal{D}^{-\ell}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot\big{(}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell}\big{)}$
-$=\sigma\big{(}\mathcal{Q}_{h}^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell},$
+$(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}{u_{h}}){\cdot}\nabla_{\mathcal{M}_{h}}{\overline{v}_{h}}$
+$\qquad=\big{(}\mathcal{D}^{-\ell}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}{\cdot}\big{(}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell}\big{)}$
+$\qquad=\sigma\big{(}\mathcal{Q}_{h}^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell},$
 which gives, after integrating both sides over $\mathcal{M}_{h}$ and using Equation 9,
@@
 (50) $\begin{split}&\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\int_{\mathcal{M}}\big{(}\mathcal{Q}_{h}\nabla_{\mathcal{M}}u_{h}^{\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A+\int_{\mathcal{M}}\big{(}\sigma^{\ell}\big{)}^{-1}Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\,\mathrm{d}A.\end{split}$
-Note that following Equations 44 and 9, $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}$ satisfies for any $u_{h},v_{h}\in S_{h}$ the equality $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})$.
-Therefore, for any $u_{h},v_{h}\in S_{h}$, we bound
+Note that following Equations 49 and 9, $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}$ satisfies for any $u_{h},v_{h}\in S_{h}$ the equality
+$\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})$. Therefore, for any $u_{h},v_{h}\in S_{h}$, we bound
 (51) $\begin{split}\big{|}\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})&-\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})\big{|}=\big{|}\mathsf{A}_{\mathcal{M}}(u_{h}^{\ell},v_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})\big{|}\\ &\leq\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\right|+\left|\int_{\mathcal{M}}(1-\big{(}\sigma^{\ell}\big{)}^{-1})Vu_{h}^{\ell}\overline{v}_{h}^{\ell}\leavevmode\nobreak\ \,\mathrm{d}A\right|.\end{split}$
-We now bound these two terms. Recall that Dziuk and Elliott (2013, Lemma 4.1) shows
+We now bound these two terms. Recall that [11, Lemma 4.1] shows
 (52) $\|\sigma\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim 1,\quad\|\sigma^{-1}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim 1,\quad\|\sigma-1\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2},\quad\|\sigma^{-1}-1\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}.$
@@
 $\leq\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}+\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}d_{s}^{\ell}\mathcal{H}\Pi\big{\|}+\big{\|}d_{s}^{\ell}\mathcal{H}\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi\big{\|}.$
-Since $\|d_{s}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}$ by Dziuk and Elliott (2013, Lemma 4.1) and $\mathcal{H}$ is defined independently of $h$, we conclude that
+Since $\|d_{s}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}$ by [11, Lemma 4.1] and $\mathcal{H}$ is defined independently of $h$, we conclude that
 (55) $\big{\|}\Pi(I-d_{s}^{\ell}\mathcal{H})\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}(I-d_{s}^{\ell}\mathcal{H})\Pi-\mathcal{D}\big{\|}$ $\lesssim\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}+h^{2}.$
@@
 $\leq\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\big{\|}\mathcal{D}\Pi\Pi_{h}^{\ell}\Pi\big{\|}+\big{\|}\Pi\mathcal{D}\big{\|}\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}.$
-Using that $\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\lesssim h^{2}$ by the proof of Dziuk and Elliott (2013, Lemma 4.1), we deduce that $\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}\lesssim h^{2}$.
+Using that $\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\lesssim h^{2}$ by the proof of [11, Lemma 4.1], we deduce that $\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}\lesssim h^{2}$.
 Injecting this inequality into Equation 55, and the resulting inequality into Equation 54, we conclude that
@@
 This allows us to write
-$\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})$ $\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A\big{|}\leq\int_{\mathcal{M}}\|(\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A$
-$\lesssim\int_{\mathcal{M}}h^{2}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A\leq h^{2}\int_{\mathcal{M}}(\mu_{\min})^{-1}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A,$
+$\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A\right|\leq\int_{\mathcal{M}}\|(\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A$
+$\qquad\lesssim\int_{\mathcal{M}}h^{2}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A\leq h^{2}\int_{\mathcal{M}}(\mu_{\min})^{-1}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A,$
 where $\mu_{\min}:\mathcal{M}\rightarrow\amsmathbb{R}_{+}$ maps any $x\in\mathcal{M}$ to the smallest eigenvalue of $\mathcal{D}(x)$ associated with an eigenvector in $\nu^{\perp}$.
@@
 Then, using the Cauchy–Schwartz inequality yields
-(51) $\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.$
+(56) $\begin{split}&\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\right|\\ &\qquad\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.\end{split}$
 Inserting the derived bounds Equation 53 and Equation 56 into Equation 51, we derive
@@
 and the result follows. ∎
-B.2. Norm and error estimates of discrete operators
+B.2. Norm estimates
+We start by introducing a few norm estimates involving the resolvents of $\mathcal{L}$ and $\mathcal{L}_{h}$.
+These results are straightforward extensions of the results introduced in [7, Lemmma 6.3], but adapted to the contour $\Gamma$ used to define functions of operators in this paper.
+We also recall a finite element error estimate which derives from a result in [7, Lemma 6.1].
+Lemma B.2.
+For any $z\in\Gamma$, $s\in[0,1]$, $v\in L^{2}(\mathcal{M})$, and $v_{h}\in S_{h}^{\ell}$, it holds
+(57) $\|\mathcal{L}^{s}(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})},$
+(58) $\|\mathcal{L}_{h}^{s}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-s)}\|v_{h}\|_{L^{2}(\mathcal{M})}.$
+Besides, for any $\beta\in[0,1]$ and for any $\varphi\in L^{2}(\mathcal{M})$,
+(59) $\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}$
+Proof.
+We start with the proof of Equation 57.
+To this end, let $z\in\Gamma$ and $s\in[0,1]$.
+Let then $v\in L^{2}(\mathcal{M})$ and $w=(z-\mathcal{L})^{-1}v\in\mathcal{D}(\mathcal{L})$, where $\mathcal{D}(\cdot)$ denotes the domain of an operator.
+Since $\mathcal{L}$ is self-adjoint, we have $\mathcal{D}(\mathcal{L}^{s})=[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}$ with isometry, where $[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}$ the intermediate space between $L^{2}(\mathcal{M})$ and $\mathcal{D}(\mathcal{L})$ obtained by the complex interpolation method [28, Theorem 16.1].
+In particular, we have
+(60) $\|\mathcal{L}^{s}w\|_{L^{2}(\mathcal{M})}=\|w\|_{[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}}\leq\|\mathcal{L}w\|_{L^{2}(\mathcal{M})}^{s}\|w\|_{L^{2}(\mathcal{M})}^{1-s}$
+where the inequality derives from a classical result on interpolation spaces (see e.g., [28, Section 5.1]). On the one hand, we obtain by Equation 20 that
+$\|w\|_{L^{2}(\mathcal{M})}=\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-1}\|v\|_{L^{2}(\mathcal{M})}.$
+On the other hand, adding and subtracting $z(z-\mathcal{L})^{-1}v$ yields
+$\|\mathcal{L}w\|_{L^{2}(\mathcal{M})}$ $=\|z(z{-}\mathcal{L})^{-1}v-v\|_{L^{2}(\mathcal{M})}\leq|z|\|(z{-}\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}+\|v\|_{L^{2}(\mathcal{M})}\lesssim 2\|v\|_{L^{2}(\mathcal{M})},$
+where we once again use Equation 20 to derive the last inequality. Combining these two estimates with Equation 60, we get
+$\|\mathcal{L}^{s}(z{-}\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}{=}\|\mathcal{L}^{s}w\|_{L^{2}(\mathcal{M})}\lesssim\|v\|^{s}_{L^{2}(\mathcal{M})}|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})}^{1-s}=|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})},$
+hence proving Equation 57.
+As for Equation 58, it is proven using the same steps as Equation 57 but substituting $\mathcal{L}$ by $\mathcal{L}_{h}$, Equation 20 by Equation 21, and $\mathcal{D}(\mathcal{L})$ by $S_{h}^{\ell}$.
+Finally, let us prove Equation 59. Let $\beta\in(0,1]$ and $\varphi\in L^{2}(\mathcal{M})$. Then, we have
+$\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$ $\lesssim\|\mathcal{L}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$
+$\lesssim\|\mathcal{L}^{(1-\beta)/2}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$
+where we use [7, Lemma 5.2] to derive the first inequality, and [7, Lemma 5.1] to derive the second inequality.
+Finally, noting that $\mathcal{L}$ satisfies elliptic regularity for indices $\alpha\in(0,1]$ (cf.
+[7, Assumption 1]), we can apply [7, Lemma 6.1] with $\alpha=\beta$ and $s=(1-\beta)/2$ to conclude that $\|\mathcal{L}^{(1-\beta)/2}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}$ and therefore
+$\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$ $\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}.$
+This concludes the proof of Lemma B.2. ∎
 We now prove some estimates for the norm of shifted inverses of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$, and for the error between inverses of these two operators.
-These results can be seen as extensions of the ones stated in Bonito et al.
-(2024, Lemma A.1).
+These results can be seen as extensions of the ones stated in [6, Lemma A.1].
 Lemma B.3.
-Let $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$ be arbitrary. Then, for all $z\in\Gamma$, for any $q\in[-1,1]$, and any $r\in(0,2)$,
-(52) $\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},$
+Let $v\in L^{2}(\mathcal{M})$, $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$ be arbitrary. Then, for all $z\in\Gamma$, for any $q\in[-1,1]$, and any $r\in(0,2)$, and any $s\in[0,1]$,
 (61) $\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}$ $\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},$
@@
 Proof.
-We start with the proof of Equation 52. To this end, let $z\in\Gamma$ and $q\in[-1,1]$. For any $s\in[0,1]$ and $v_{h}\in S_{h}^{\ell}$, we expand
-$\mathcal{L}_{h}^{s}v_{h}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{s}(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}e_{i}^{h},$
-implying that
-$\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}=\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}=\sum_{i=1}^{N_{h}}\left(|\lambda_{i}^{h}||(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|\right)^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2(1-s)}.$
-Hölder’s inequality yields
-(56) $\begin{split}\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}&\leq\left(\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{s}\left(\sum_{i=1}^{N_{h}}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{1-s}\\ &=\|\mathcal{L}_{h}v_{h}\|_{L^{2}(\mathcal{M})}^{2s}\|v_{h}\|_{L^{2}(\mathcal{M})}^{2(1-s)}.\end{split}$
-Let $w_{h}=(z-\mathcal{L}_{h})^{-1}v_{h}$, then we obtain by Equation 20 that
-$\|w_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-1}\|v_{h}\|_{L^{2}(\mathcal{M})}.$
-Hence, adding and subtracting $z(z-\mathcal{L}_{h})^{-1}$ yields
-$\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}$ $=\|z(z-\mathcal{L}_{h})^{-1}v_{h}-v_{h}\|_{L^{2}(\mathcal{M})}\leq|z|\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}+\|v_{h}\|_{L^{2}(\mathcal{M})}$
-$\lesssim 2\|v_{h}\|_{L^{2}(\mathcal{M})}.$
-Combining this estimate with Equation 56, we get for any $s\in[0,1]$
-(57) $\|\mathcal{L}_{h}^{s}w_{h}\|_{L^{2}(\mathcal{M})}\leq\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}^{s}\|w_{h}\|_{L^{2}(\mathcal{M})}^{(1-s)}\leq|z|^{-(1-s)}\|v_{h}\|^{s}_{L^{2}(\mathcal{M})}\|v_{h}\|_{L^{2}(\mathcal{M})}^{1-s}=|z|^{s-1}\|v_{h}\|_{L^{2}(\mathcal{M})},$
-yielding for any $q\in[-1,1]$ that
-(58) $\begin{split}\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}&=\|\mathcal{L}_{h}^{1-(1+q)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})}\\ &\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},\end{split}$
-where we applied Equation 57 with $s=1-(1+q)/2$ in the last step. Hence, we retrieve Equation 52.
-Next, to prove Equation 53, we first observe that the estimates in Equation 57 and Equation 58 carry over to the case when $\mathsf{L}_{h}$ is used instead of $\mathcal{L}_{h}$, as the proof in essence is a standard manipulation of a finite eigenexpansion.
-In particular, we obtain (by taking $q=1$ and $s=0$)
-(59) $\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}.$
-Now, Equation 59 combined with the equivalence of norms in Equation 19, yields that
-$\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},$
-which shows Equation 53.
+To prove Equation 61, we first observe that the estimate in Equation 58 carry over to the case when $\mathsf{L}_{h}$ is used instead of $\mathcal{L}_{h}$, and $L^{2}(\mathcal{M})$ (resp.
+$S_{h}^{\ell}$) is replaced by its counterpart $L^{2}(\mathcal{M}_{h})$ (resp.
+$S_{h}$) on the polyhedral surface.
+Hence, for any $z\in\Gamma$, $V_{h}\in L^{2}(\mathcal{M}_{h})$, we have
+$\|\mathsf{L}_{h}^{s}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-(1-s)}\|V_{h}\|_{L^{2}(\mathcal{M}_{h})}$
+In particular, we retrieve (by taking $s=1/2$)
+(64) $\begin{split}\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}&=\|\mathsf{L}_{h}^{1/2}(z-\mathsf{L}_{h})^{-1}\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\\ &\lesssim|z|^{-1/2}\|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},\end{split}$
+where we used the equivalence of norms (19) in the last inequality.
 To bound Equation 62, we apply Equation 58 with $s=r/2$ to obtain
-$\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}=\|\mathcal{L}_{h}^{r/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{r/2-1}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
+$\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}=\|\mathcal{L}_{h}^{r/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-(1-r/2)}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
 Finally, to prove the bound in Equation 63, we rely on the geometric consistency estimate of Lemma B.1.
@@
 for all $w_{h}\in S_{h}^{\ell}$. Likewise, for $\mathsf{L}_{h}^{-1}$ we obtain
-$(\mathsf{L}_{h}U_{h},W_{h})_{L^{2}(\mathcal{M}_{h})}=\mathsf{A}_{\mathcal{M}_{h}}(U_{h},W_{h})=(\mathsf{P}_{h}(\sigma v_{h}^{-\ell}),W_{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma v_{h}^{-\ell},W_{h})_{L^{2}(\mathcal{M}_{h})}=(v_{h},W_{h}^{\ell})_{L^{2}(\mathcal{M})},$
+$(\mathsf{L}_{h}U_{h},W_{h})_{L^{2}(\mathcal{M}_{h})}$ $=\mathsf{A}_{\mathcal{M}_{h}}(U_{h},W_{h})=(\mathsf{P}_{h}(\sigma v_{h}^{-\ell}),W_{h})_{L^{2}(\mathcal{M}_{h})}$
+$=(\sigma v_{h}^{-\ell},W_{h})_{L^{2}(\mathcal{M}_{h})}=(v_{h},W_{h}^{\ell})_{L^{2}(\mathcal{M})},$
 for all $W_{h}\in S_{h}$, where we used the definition of $\sigma$ in the last step. Let us now select a fixed, but arbitrary, $\Xi_{h}\in S_{h}$. Then, by combining the last two equations,
-$|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|$ $=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}_{h}}(U_{h},\Xi_{h})|=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-(v_{h},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}|$
-$=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}}(u_{h},\Xi_{h}^{\ell})|,$
+$|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|$ $=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}_{h}}(U_{h},\Xi_{h})|$
+$=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-(v_{h},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}|=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}}(u_{h},\Xi_{h}^{\ell})|,$
 meaning that an application of Lemma B.1 results in the bound
@@
 Then, applying the triangle inequality to the (last) right-hand side of Equation 66 gives
-$\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})+\big{|}\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})\big{|}\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2},$
+$\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})+\big{|}\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})\big{|}$
+$\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2},$
 where we used Equation 67 and Lemma B.1 to derive the second inequality.
@@
 where the equivalence of norms (19) together with the definition of $u_{h}$ are used in the final step. This concludes the proof of Equation 63. ∎
-B.3. Proof of Equation 24
+Finally, we recall the Bramble-–Hilbert lemma (cf. [6, Equation (4.8)]), which is used in several proofs in this paper.
+Lemma B.4 (Bramble-–Hilbert lemma).
+For any $t\in[0,2]$ and any $\varphi\in H^{t}(\mathcal{M})$,
+(69) $\|(I-P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{t}\|\varphi\|_{H^{t}(\mathcal{M})}\lesssim h^{t}\|\mathcal{L}^{t/2}\varphi\|_{L^{2}(\mathcal{M})},$
+where we used the equivalence of Sobolev norms dot-spaces norms in the last inequality.
+B.3. Proof of Lemma 2.5
+Proof.
+Let $h\in(0,h_{0})$, $z\in\Gamma$.
+Let $f\in L^{2}(\mathcal{M})$ and let $p\in[0,1]$ such that $\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}<\infty$.
+Finally, let $\beta\in[0,1]$ such that $p\in[0,(1+\beta)/2]$.
+We then define $\mathcal{F}_{h}(z)$ by $\mathcal{F}_{h}(z)=(z-\mathcal{L}_{h})^{-1}P_{h}-P_{h}(z-\mathcal{L})^{-1}$.
+Note then
+$\mathcal{F}_{h}(z)$ $=(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(\mathcal{L}_{h}^{-1}P_{h}(z-\mathcal{L})\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}(z-\mathcal{L}_{h})P_{h}\mathcal{L}^{-1})\mathcal{L}(z-\mathcal{L})^{-1}$
+$=(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}$
+Hence,
+$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $=\|(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
+$=\|\mathcal{L}_{h}^{1-(1-\beta)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{(1-\beta)/2}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
+$\lesssim|z|^{-(1-\beta)/2}\|\mathcal{L}_{h}^{(1-\beta)/2}(P_{h}\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
+where we used Equation 58 with $s=(1-\beta)/2\in[0,1/2]$ to derive the inequality. Note then that, using Equation 59, we have
+$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-\beta)/2}h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\mathcal{L}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
+$=h^{2\beta}|z|^{-(1-\beta)/2}\|\mathcal{L}^{(1+\beta)/2}(z-\mathcal{L})^{-1}f\|_{L^{2}(\mathcal{M})}$
+$=h^{2\beta}|z|^{-(1-\beta)/2}\|\mathcal{L}^{(1+\beta)/2-p}(z-\mathcal{L})^{-1}\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
+Using then Equation 57 with $s=(1+\beta)/2-p\in[0,1]$, we retrieve
+$\|\mathcal{F}_{h}(z)f\|_{L^{2}(\mathcal{M})}$ $\lesssim h^{2\beta}|z|^{-(1-\beta)/2}|z|^{-(1-(1+\beta)/2+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}$
+$=h^{2\beta}|z|^{-(1-\beta+p)}\|\mathcal{L}^{p}f\|_{L^{2}(\mathcal{M})}.$
+This concludes the proof. ∎
+B.4. Proof of Equation 26
 Based on the results in the previous subsections, we can now move on to the proof of Equation 26.
@@
 Note that by the integral representations of the operators (22)
-$\mathcal{E}_{h}=\big{\|}\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}(z)\,\mathrm{d}z\big{\|}_{L^{2}(\mathcal{M}_{h})},$
+$\mathcal{E}_{h}=\big{\|}\frac{1}{2\pi i}\int_{\Gamma}\gamma(z)\mathcal{F}(z)\tilde{f}\,\mathrm{d}z\big{\|}_{L^{2}(\mathcal{M}_{h})},$
 where we take for any $z\in\Gamma$, $\mathcal{F}(z)=\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}-(z-\mathsf{L}_{h})^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}$.
 Similarly, as in the proof of Proposition 2.4, we use the splitting (23) of $\Gamma$ and the triangle inequality to deduce that
-(64) $\mathcal{E}_{h}$ $\leq\frac{1}{2\pi}\int_{\Omega_{+}}|\gamma(g_{+}(t))|\left\|\mathcal{F}(g_{+}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{\Omega_{0}}|\gamma(g_{0}(t))|\|\mathcal{F}(g_{0}(t)\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t$
-$\qquad+\frac{1}{2\pi}\int_{\Omega_{-}}|\gamma(g_{-}(t))|\left\|\mathcal{F}(g_{-}(t))\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
+(70) $\mathcal{E}_{h}$ $\leq\frac{1}{2\pi}\int_{\Omega_{+}}|\gamma(g_{+}(t))|\left\|\mathcal{F}(g_{+}(t))\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t+\frac{\delta_{0}}{2\pi}\int_{\Omega_{0}}|\gamma(g_{0}(t))|\|\mathcal{F}(g_{0}(t))\tilde{f}\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t$
+$\qquad+\frac{1}{2\pi}\int_{\Omega_{-}}|\gamma(g_{-}(t))|\left\|\mathcal{F}(g_{-}(t))\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\,\mathrm{d}t,$
 […]
@@
 (71) $\mathcal{E}_{h}^{*}$ $\lesssim\int_{\Gamma_{t}}|\gamma(g_{*}(t))|\big{(}\left\|\mathcal{F}_{1}(g_{*}(t))\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}+\big{\|}\mathcal{F}_{2}(g_{*}(t))\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}\right)\,\mathrm{d}t.$
-We first bound $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$. Using successively Equation 23 and the geometric estimates in Bonito et al. (2018, Corollary 2.2) results in
+We first bound $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$. Using successively Equation 25 and the geometric estimates in [5, Corollary 2.2] results in
 $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$ $=|z|\left\|(z-\mathsf{L}_{h})^{-1}(I-\mathsf{P}_{h}\sigma)\mathcal{P}_{\ell}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$
@@
 (72) $\left\|\mathcal{F}_{1}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}|z|^{-(1-p/2)}\left\|\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}^{2}.$
-To bound $\big{\|}\mathcal{F}_{2}(z)\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}$, we apply Equations 53 and 55 to obtain
+To bound $\big{\|}\mathcal{F}_{2}(z)\tilde{f}\big{\|}_{L^{2}(\mathcal{M}_{h})}$, we apply Equations 61 and 63 to obtain
 $\left\|\mathcal{F}_{2}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}$ $\lesssim|z|^{-1/2}\left\|\left(\mathsf{L}_{h}^{-1}\mathsf{P}_{h}\sigma\mathcal{P}_{\ell}-\mathcal{P}_{\ell}\mathcal{L}_{h}^{-1}\right)\mathcal{L}_{h}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{H^{1}(\mathcal{M}_{h})}$
 $\lesssim|z|^{-1/2}h^{2}\left\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}\tilde{f}\right\|_{L^{2}(\mathcal{M})}$
-and with Equation 52 for any $p\in[-1,1]$ ($p=-q$)
+$=|z|^{-1/2}h^{2}\left\|\mathcal{L}_{h}^{(1+p)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}$
+and with Equation 58 (applied with $s=(1+p)/2$)
 (73) $\left\|\mathcal{F}_{2}(z)\tilde{f}\right\|_{L^{2}(\mathcal{M}_{h})}\lesssim h^{2}|z|^{-(1-p/2)}\left\|\mathcal{L}_{h}^{-p/2}\tilde{f}\right\|_{L^{2}(\mathcal{M})}.$
-Using Equations 67 and 66 with $p=\min\{\alpha+d/4;1\}$ together with Equation 65 gives
+Using Equations 73 and 72 with $p=\min\{\alpha+d/4;1\}$ together with Equation 71 gives
 $\mathcal{E}_{h}^{*}$ $\lesssim\int_{\Omega_{*}}|\gamma(g_{*}(t))|h^{2}|g_{*}(t)|^{-(1-p/2)}\|\mathcal{L}_{h}^{-p/2}\tilde{f}\|_{L^{2}(\mathcal{M})}\,\mathrm{d}t,$
-which yields in turn (since $\gamma$ is an $\alpha$-power spectral density)
+which yields in turn (since $\gamma$ is an $\alpha$-amplitude spectral density)
 $\mathcal{E}_{h}^{*}\lesssim h^{2}\|\mathcal{L}_{h}^{-p/2}\tilde{f}\|_{L^{2}(\mathcal{M})}\int_{\Omega_{*}}|g_{*}(t)|^{-(1+\alpha-p/2)}\,\mathrm{d}t\lesssim h^{2}\|\mathcal{L}_{h}^{-\min\{\alpha+d/4;1\}/2}\tilde{f}\|_{L^{2}(\mathcal{M})},$